### v1

# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

Abstract
We study the quantum dynamics of a homogeneous ideal Fermi gas coupled to an impurity particle on a three-dimensional box with periodic boundary condition. For large Fermi momentum $k_{\textnormal{F}}$, we prove that the effective dynamics is generated by a Fröhlich-type polaron Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. Moreover, we prove that the effective dynamics can be approximated by an explicit coupled coherent state. Our method is applicable to two relevant settings: first, an interaction coupling $\lambda=1$ and masses of order 1 for time scales of order $k_{\textnormal{F}}^{-1}$; second to the case of $\lambda=k_{\textnormal{F}}^{-1}$ and a heavy Fermi gas with masses of order $k_{\textnormal{F}}$ for time scales of order 1.
1 Introduction
The study of impurities in quantum gases has garnered considerable attention due to its relevance in various physical contexts, ranging from solid-state physics to cold atom experiments. In this context quasi-particles such as polarons stand as intriguing entities emerging from the interaction of a single impurity particle with a surrounding medium. The concept of a polaron, originally introduced by Lev Landau to study the motion of an electron in a dielectric crystal [1965], most famously emerges from the celebrated Fröhlich Hamiltonian in second quantization formalism describing electron-phonon interactions [Froehlich1954]. Subsequently, the polaron concept was extended to all kind of surrounding media including Bose and Fermi gases. The formation conditions and properties of polarons are believed to play a central role to understand the transport properties and the effective mass of impurities within the host material.
In this article, we study with mathematical rigor the dynamics of an impurity particle immersed in a dense gas of fermions as surrounding medium. Interactions between fermions are neglected and we assume that the initial state $\psi$ of the system is a product state between the impurity state and a filled Fermi ball. This mathematical framework finds resonance with recent experimental and theoretical advancements in the study of ultracold atoms [SWSZ09, KSN+12, CJL+15]. We show that the effective dynamics of the system is governed by a Fröhlich-type Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. More specifically, the excitations relative to the filled Fermi ball are up to a constant described by the Hamiltonian
$\mathbb{H}^{\text{F}}=(-\Delta_{y})\otimes 1+1\otimes\mathbb{D}_{\text{B}}+\Phi(h_{y})$ (1.1)
with $(-\Delta_{y})$ describing the kinetic energy of the impurity particle, $\mathbb{D}_{\text{B}}$ describing the kinetic energy of the excitation field and $\Phi(h_{y})\coloneqq c^{\ast}(h_{y})+c(h_{y})$ the linear coupling between impurity particle and excitations. The operators $c^{\ast}$ and $c$ describing this excitation field coincide with those introduced in a series of pioneering studies on the correlation energy of interacting fermions [BNP+19, BNP+21a, BNP+21b]. We note that an effective Hamiltonian of a similar type to (1.1) has recently been derived in another microscopic setting involving a tracer particle interacting with excitations of a Bose Einstein condensate [LP22, MS20].
Subsequently, we show that the effective time evolved state can be up to a phase factor approximated by a time-dependent coupled coherent state $W(\eta_{t})\phi\otimes\Omega$ where $W$ is the Weyl operator which is simply parameterized by a function $\eta_{t}$. An explicit expression for $\eta_{t}$ is derived which allows for determining the number of collective excitations over time. We believe that such quantities are in particularly helpful to gain deeper insights into the formation process of quasi-particles as studied in experiments such as [CJL+16]. Eventually, we show that the linear coupling term $\Phi(h_{y})$ cannot be omitted in the effective description but adds a leading order effect to the effective dynamics in our setting.
Our results hold for a variety of time scales and couplings, describing different interaction strengths and mass ratio, which will be specified in the subsequent section. We note that the same microscopic model has been studied in [JMPP17, JMP18, MP21] but with very specific choices of couplings different from ours, leading to an effective decoupling of impurity and gas.
1.1 The microscopic model
We consider an impurity particle interacting with $N$ spinless fermions on a 3-dimensional box with periodic boundaries described by $\Lambda\coloneqq\mathbb{T}^{3}\coloneqq\mathbb{R}^{3}/(2\pi\mathbb{Z}^{3})$. The system is described by a state in the Hilbert space $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\mathcal{H}_{N}^{-}=L^{2}(\Lambda)^{\wedge N}$ where $y$ is the coordinate of the impurity particle and $\{x_{i}\}_{i=1,\ldots,N}$ are the coordinates of the fermions. The Hamiltonian of the system is given by
$H_{N}=-\beta\Delta_{y}+\kappa\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda\sum_{i=1}^{N}V(x_{i}-y)$ (1.2)
and parameterized by the coupling constants $\beta,\kappa,\lambda>0$. Note that the different parts of the Hamiltonian on different tensor components of our Hilbert space writing, i.e. we used the short-hand notation writing, for example, $-\Delta_{y}\equiv-\Delta_{y}\otimes 1$ for the Laplacian acting on the impurity particle. The interaction $V$ is assumed to have a Fourier transform $\hat{V}$ with compact support satisfying $\hat{V}(k)=\hat{V}(-k)$ for all $k\in\mathbb{Z}^{3}$. It is well-known that under this assumption the Hamiltonian (1.2) defines a self-adjoint operator which generates by Stone’s theorem the unitary time evolution $e^{-H_{N}t}$.
We are interested in the dynamics of the system governed by the time-dependent Schrödinger equation of the form
$i\frac{\textnormal{d}}{\textnormal{d}t}\psi_{t}=H_{N}\psi_{t},\qquad\psi_{0}\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}.$ (1.3)
Note that the ground state of the non-interacting Fermion system is well-known, non-degenerate and explicitly given by
$\Omega_{0}\coloneqq\bigwedge_{k\in B_{\textnormal{F}}}f_{k},\qquad f_{k}(x)\coloneqq\frac{\exp(ikx)}{(2\pi)^{3/2}}\in L^{2}(\Lambda).$ (1.4)
We choose the initial state to be of product form
$\psi_{0}\big{(}y;x_{1},\ldots x_{N}\big{)}\coloneqq\phi(y)\otimes\Omega_{0}(x_{1},\ldots x_{N}),$ (1.5)
with a general state for the impurity particle i.e. the system is initially prepared in a state describing the ground state of the ideal Fermi gas which does not interact with the impurity particle.
Furthermore, we choose the Fermi momentum $k_{F}$ to be our parameter of the system in the sense that the particle number is defined as
$N\equiv N(k_{\textnormal{F}})\coloneqq|B_{\textnormal{F}}|,\qquad B_{\textnormal{F}}\coloneqq\{k\in\mathbb{Z}^{3}:|k|\leq k_{\textnormal{F}}\},$ (1.6)
i.e. the particle number $N$ of the Fermi gas is chosen such that the Fermi ball is completely filled. Note that the average density is in this case proportional to the number $N$ of gas particles due to the following relation
$k_{\textnormal{F}}=\left(\frac{3}{4\pi}\right)^{1/3}N^{\frac{1}{3}}+\mathcal{O}(1)$ (1.7)
which is a consequence of Gauss’ counting algorithm.
1.2 Scaling regimes
In the following, we present the choices of $\beta,\kappa$ and $\lambda$ we are aiming for, and discuss the physical meaning of the couplings.
  • The coupling $\lambda$ models the coupling strength between the Fermi gas and the impurity. For small $\lambda$ we expect a decoupling between the gas and the impurity in the sense that the time evolution given by (1.3) does not entangle an initial state of product form $\phi\otimes\Omega_{0}$. Such a result was shown in [MP21] with $\beta,\kappa=1$ and $\lambda=k_{\textnormal{F}}^{-1/2}\sim N^{-\frac{1}{6}}$ in three dimensions and [JMPP17, JMP18] with $\beta=\kappa=\lambda=1$ in two dimensions.
  • The couplings $\beta$ and $\kappa$ determine the order of the kinetic energy of the impurity and Fermi gas respectively.
  • In addition, it is important to discuss on which time scales results hold. Define $\tau$ to be the relevant time scale for our statements, i.e. our statements $A(t)$ shall be considered for $t\in[0,T]$ with $T\in\mathcal{O}(\tau)$.
Our results hold for the following set of couplings
$S=\left\{(\beta,\kappa,\lambda,\tau):\beta\tau\in o(1),\ \kappa=\lambda,\ \lambda k_{\textnormal{F}}\tau\in\mathcal{O}(1)\right\}\subset\mathbb{R}_{\geq 0}^{4}.$ (1.8)
The physical interpretation can be better understood in the following specific cases:
Fermi time dynamics $\kappa=\lambda=1$, $\beta=k_{\textnormal{F}}^{-b}$ with $b>-1$, $\tau=k_{\textnormal{F}}^{-1}$
In this setting the kinetic energy of the Fermi gas and the interaction term are chosen independent of the the Fermi momentum $k_{\textnormal{F}}$ and, in particular of the particle number $N$. In return the considered time scale is $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$ and therefore short for large $k_{\textnormal{F}}$. Planck’s constant $\hbar$ is set to be $1$, the Fermion mass is of order 1. For convenience we set $m=1/2$ in this case. The mass $m_{\text{imp}}=\frac{1}{2}k_{\textnormal{F}}^{b}$ of the impurity particle might depend on $k_{\textnormal{F}}$ and represent for $b\in(-1,0)$ and large $k_{\textnormal{F}}$ a light, for $b>0$ a heavy or for $b=0$ an equally heavy impurity particle in comparison to the fermions. We remark that the results in [MP21] can be transferred to this setting of $\kappa=\lambda=1$, however, allowing only for shorter timescales of $\tau=k_{\textnormal{F}}^{-1-\delta}$ for a $\delta>0$. This seems compatible since we are not interested in a free time evolution of the impurity particle. The times $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$ are on the time scale of the fermions near the Fermi surface, which typically have a momentum of $k_{\textnormal{F}}$.
Heavy fermion regime $\kappa=\lambda=k_{\textnormal{F}}^{-1}$, $\beta=k_{\textnormal{F}}^{-b}$ with $b>0$, $\tau=1$
In this setting, we identify the couplings $\kappa$ and $\beta$ with the masses of the fermions and the impurity particle, respectively. More concretely, the fermion mass is $m=\frac{1}{2}k_{\textnormal{F}}$ and the mass of the impurity particle is $m_{\text{imp}}=\frac{1}{2}k_{\textnormal{F}}^{b}$ where the factor $\frac{1}{2}$ is chosen for convenience. Since we are interested in large $k_{\textnormal{F}}$, this case corresponds to a heavy Fermi gas setting where the impurity particle has a relative mass ratio of $\kappa/\beta$ depending on $b>0$. The coupling strength $\lambda$ between both impurity particle and gas is weak but much stronger than in the mean-field setting where one would introduce an averaging factor of $\lambda=N^{-1}\sim k_{\textnormal{F}}^{-3}$. Planck’s constant $\hbar$ is set to be $1$ in this case and is particularly independent of the number of gas particles.
Semiclassical regime $\kappa=\lambda=k_{\textnormal{F}}^{-2}$, $\beta=k_{\textnormal{F}}^{-b}$ with $b>1$, $\tau=1$
We can also identify the couplings with Planck’s constant $\hbar$ instead of setting it to $1$. In this case, we consider the time evolution governed by the Schrödinger equation
$i\hbar\partial_{t}\psi_{t}=\left(\hbar^{2}\beta^{\prime}(-\Delta_{y})+\hbar^{2}\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda\sum_{i=1}^{N}V(x_{i}-y)\right)\psi_{t}$ (1.9)
with $\hbar\coloneqq\kappa^{1/2}=\lambda^{1/2}=k_{\textnormal{F}}^{-1}$. The fermion mass is given by $m=1/2$ and mass of the impurity particle $m_{\text{imp}}=\beta^{\prime-1}=\frac{1}{2}k_{\textnormal{F}}^{b^{\prime}}$ with $b^{\prime}>-1$. The time evolution is therefore given by the operator $U_{t}=e^{-iH_{N}t/\hbar}$. Since the time scale $\tau=1$ is absorbed by the factor $t/\hbar$ is of order $k_{\textnormal{F}}$. [^1] Note that the effective Hamiltonian is expected to be of order $\hbar$ here such that the exponent of the time evolution operator is of order $1$. Since we are interested in $k_{\textnormal{F}}\gg 1$, the interpretation is that the system is considered in the semiclassical regime with $\hbar\ll 1$. Such a semi-classical regime with $\hbar\equiv k_{\textnormal{F}}^{-1}$ has been widely studied in the analysis of Fermi gases, as can be seen for example in [Ben22, Saf23].
From a technical point of view all settings are very similar and our results will apply to all described cases. We will focus our presentation on the setting with $\hbar\equiv 1$ and give some remarks how the results can be translated to the semiclassical setting.
2 Preliminaries
2.1 Second quantization
It is convenient to consider $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ as $N$-particle sector of $L^{2}(\Lambda)\otimes\mathcal{F}$ with the fermionic Fock space $\mathcal{F}$ constructed over $L^{2}(\Lambda)$. This way, we have access to the powerful formalism of second quantization with the fermionic creation operator $a_{p}^{\ast}$ creating a particle with momentum $p\in\mathbb{Z}^{3}$ and the annihilation operator $a_{p}$ annihilating a particle with momentum $p\in\mathbb{Z}^{3}$. Those operators satisfy the canonical anticommutation relations (CAR)
$\forall p,q\in\mathbb{Z}^{3}:\{a_{p},a_{q}^{\ast}\}=\delta_{p,q},\qquad\{a_{p},a_{q}\}=0=\{a_{p}^{\ast},a_{q}^{\ast}\}.$ (2.1)
Furthermore we introduce the fermionic number operator $\mathcal{N}\coloneqq\sum_{p\in\mathbb{Z}^{3}}a_{p}^{\ast}a_{p}$ and the vacuum $\Omega$ satisfying $a_{p}\Omega=0$ for all $p\in\mathbb{Z}^{3}$.
We lift our $N$-particle Hamiltonian $H_{N}$ to Fock space as
$\mathbb{H}=\underbrace{-\beta\Delta_{y}}_{\eqqcolon h_{0}}+\underbrace{\kappa\sum_{k\in\mathbb{Z}^{3}}|k|^{2}a_{k}^{\ast}a_{k}}_{\eqqcolon\mathbb{H}^{\textnormal{kin}}}+\underbrace{\lambda\sum_{k,p\in\mathbb{Z}^{3}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}}_{\eqqcolon\mathbb{V}}$ (2.2)
[…]
We will mostly use the abuse of notation $\mathbb{A}\equiv 1\otimes\mathbb{A}$ as operator on $L^{2}(\Lambda)\otimes\mathcal{F}$ where $\mathbb{A}$ acts as an operator on the Fock space part.
2.2 Particle-hole transformation
In our analysis, a primary objective is to focus on excitations relative to the non-interacting Fermi ball. In particular, we want to use a description of our fermionic system in which the non-interacting Fermi ball $\Omega_{0}=\prod_{k\in B_{\textnormal{F}}}a^{\ast}_{k}\Omega$ is mapped to the vacuum. To achieve this, we employ the particle-hole transformation, which is a specific type of fermionic Bogoliubov transformation as creation operators are mapped to linear combinations of creation and annihilation operators while preserving the CAR. The particle-hole transformation is defined as the map $R:\mathcal{F}\to\mathcal{F}$ satisfying
$R^{\ast}a_{k}^{\ast}R\coloneqq\begin{cases}a_{k}^{\ast}&,k\in B_{\textnormal{F}}^{c}\\ a_{k}&,k\in B_{\textnormal{F}}\end{cases},\qquad R\Omega\coloneqq\Omega_{0}\ .$ (2.3)
It is easy to check that the map is well-defined, unitary and satisfies $R^{-1}=R^{\ast}=R$.
With this, we can re-write the initial state (1.5) representing a non-interacting impurity particle and a Fermi gas as
$\psi_{0}=\phi\otimes\Omega_{0}=(1\otimes R)\left(\phi\otimes\Omega\right)\eqqcolon R\psi.$ (2.4)
Later on, we will mostly use the product state $\psi=\phi\otimes\Omega$ of the impurity and the vacuum instead of $\psi_{0}$.
Furthermore, we define
$E_{N}^{\text{pw}}\coloneqq\kappa\sum_{k\in B_{\textnormal{F}}}|k|^{2}=\langle R\Omega,H_{N},R\Omega\rangle$ (2.5)
to be the energy of the non-interacting Fermi ball.
Of greatest interest is of course the action of the particle-hole transformation on the microscopic Hamiltonian as generator of the dynamics. The conjugation with $R$ of $\mathbb{H}=-\beta\Delta_{y}+\mathbb{H}^{\textnormal{kin}}+\mathbb{V}$ yields
$\mathbb{H}_{0}$ $\coloneqq R^{\ast}\mathbb{H}^{\textnormal{kin}}R-E_{N}^{\text{pw}}=\kappa\sum_{k\in\mathbb{Z}^{3}}|k|^{2}R^{\ast}a_{k}^{\ast}RR^{\ast}a_{k}R-E_{N}^{\text{pw}}$ (2.6)
$=\kappa\sum_{k\in B_{\textnormal{F}}}|k|^{2}a_{k}a_{k}^{\ast}+\kappa\sum_{k\in B_{\textnormal{F}}^{c}}|k|^{2}a_{k}^{\ast}a_{k}-E_{N}^{\text{pw}}$ (2.7)
$=\sum_{k\in\mathbb{Z}^{3}}e(k)a_{k}^{\ast}a_{k}\qquad\text{with }e(k)\coloneqq\begin{cases}\kappa|k|^{2}&,k\in B_{\textnormal{F}}^{c}\\ -\kappa|k|^{2}&,k\in B_{\textnormal{F}}.\end{cases}$ (2.8)
Similarly, we can see that
$R^{\ast}\mathbb{V}R$ $=\lambda\sum_{k,p\in\mathbb{Z}^{3}}\hat{V}(k)e^{iky}R^{\ast}a_{p}^{\ast}RR^{\ast}a_{p-k}R$
$=\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p-k\in B_{\textnormal{F}}^{c},\\ p\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}a_{p-k}+\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}^{\ast}$ (2.9a)
$\quad+\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}},\\ p-k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}a_{p-k}^{\ast}+\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}^{c}\end{subarray}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}.$ (2.9b)
For later purposes we shall introduce for $\varphi\in l^{2}(\mathbb{Z}^{3})$ the short-notation
$b(\varphi)$ $\coloneqq\sum_{k\in\mathbb{Z}^{3}}\overline{\varphi(k)}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}a_{p-k}a_{p},$ (2.10)
$b^{\ast}(\varphi)$ $=\sum_{k\in\mathbb{Z}^{3}}\varphi(k)\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}a_{p}^{\ast}a_{p-k}^{\ast}.$ (2.11)
We can then write
$R^{\ast}\mathbb{H}R=-\beta\Delta_{y}+\mathbb{H}_{0}+b^{\ast}(\tilde{h})+b(\tilde{h})+\mathcal{E}$ (2.12)
with $\tilde{h}(k)\coloneqq\tilde{h}_{y}(k)\coloneqq\lambda\hat{V}(k)e^{iky}$ and $\mathcal{E}$ is given by the terms of (2.9b) since
$\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p-k\in B_{\textnormal{F}}^{c},\\ p\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}a_{p-k}=$ $\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}\tilde{p}\in B_{\textnormal{F}}^{c},\\ \tilde{p}+k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{\tilde{p}+k}a_{\tilde{p}}$
$=$ $\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{-iky}a_{p-k}a_{p}$ (2.13)
where we used that $\hat{V}(-k)=\hat{V}(k)$.
2.3 Almost-bosonic operators and patch decomposition
Our effective description of the microscopic system described by (1.2) will involve the emergence of almost-bosonic particles describing pair excitations of the Fermi ball. Those pair excitations will be delocalized over the Fermi surface in the sense that they correspond to a linear combination of two fermionic creation operators. As mentioned before the almost-bosonic pair operators which occur in this article coincide with the ones introduced in the series of seminal works [BNP+19, BNP+21a, BNP+21b, BPSS23] on the correlation energy of a weakly interacting Fermi gas. We give a brief introduction to the construction of those operators with the most relevant properties in the this subsection and in Section A.
A key ingredient for the approximation of the microscopic fermionic system by almost-bosonic excitations is the decomposition of the Fermi surface into patches. This will allow to approximate the fermionic kinetic energy term by a term quadratic in the almost-bosonic pair operators.
Introduce the bisecting subset of $\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}$
$\Gamma^{\textnormal{nor}}\coloneqq\left\{(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}:k_{3}>0\vee k_{3}=0,k_{2}>0\vee k_{2}=k_{3}=0,k_{1}>0\right\}$ (2.14)
allowing the decomposition $\Gamma^{\textnormal{nor}}\sqcup(-\Gamma^{\textnormal{nor}})=\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}$.
The construction works as follows:
  1. (i).
    Choose the number $M$ of patches satisfying
    $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta},\qquad\delta\in(0,\frac{1}{6}).$ (2.15)
    Equivalently, we can also write the condition as $M=N^{c\delta}$ with $\delta\in(0,1/6)$ and $c\in(2,\frac{2}{3\delta}-2$). The lower bound on $M$ is needed to control the number of momenta inside each patch whereas the upper bound is needed to suppress Pauli’s principle. The choice of $\delta$ and $c$ will be taken later.
  2. (ii).
    Define equal-area disjoint patches $p_{\alpha}$ centered around $\omega_{\alpha}\in S^{2}$ as follows
    • $p_{1}$ is spherical cap of area $4\pi/M$,
    • decompose remaining semi-sphere into $\sqrt{M}/2$ collars,
    • leave corridors of width $2R\coloneqq 2\ \text{supp}\hat{V}$ between adjacent patches,
    • define patches of southern semi-sphere by reflection $k\mapsto-k$.
  3. (iii).
    For given $k\in\Gamma^{\textnormal{nor}}$ define the index set $\mathcal{I}_{k}\coloneqq\mathcal{I}_{k}^{+}\cup\mathcal{I}_{k}^{-}$ consisting of north and south patch indices
    $\mathcal{I}_{k}^{+}$$\coloneqq\{\alpha\in\{1,\ldots,M\}\mid k\cdot\hat{\omega}_{\alpha}\geq N^{-\delta}\},$ (2.16)
    $\mathcal{I}_{k}^{-}$$\coloneqq\{\alpha\in\{1,\ldots,M\}\mid k\cdot\hat{\omega}_{\alpha}\leq-N^{-\delta}\}.$ (2.17)
    A thin strip around the equator has to be excluded since the number of momenta per equator patch is too small. Note that $\delta>0$ coincides with the parameter in step 1.
  4. (iv).
    Define the collective almost-bosonic creation operator and its normalization factor as
    $c_{\alpha}^{\ast}(k)\coloneqq b_{\alpha}^{\ast}(\pm k),\ \ n_{\alpha}(k)\coloneqq m_{\alpha}(\pm k)\iff\alpha\in\mathcal{I}_{k}^{\pm}$ (2.18)
    with
    $b_{\alpha}^{\ast}(k)\coloneqq\frac{1}{m_{\alpha}(k)}\sum_{p\in B_{F}^{c}\cap B_{\alpha},p-k\in B_{F}\cap B_{\alpha}}a_{p}^{\ast}a_{p-k}^{\ast}$ (2.19)
    being sensitive to being on the north or south hemisphere. The creation operator can be seen as collective in the sense that it involves a superposition of all possible fermion pairs with relative momentum $k$.
Similarly to (2.11) introduced in the previous subsection, we define
$c^{\ast}(\eta)\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}c_{\alpha}^{\ast}(k)$ (2.20)
with inner product $\langle\eta,\varphi\rangle\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}\varphi_{\alpha}(k)$ for all $\eta,\varphi\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$.
The following statements hold as a consequence of the above construction.
  • The surface area of a patch satisfies $\sigma(p_{\alpha})\in\mathcal{O}(1/M)$.
  • The Canonical Anticommutation Relations (CAR) are satisfied up to an error term (see [BNP+19, Lemma 4.1]): It holds for all $k^{\prime},k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$
    $[c_{\alpha}(k),c_{\beta}(k^{\prime})]$ $=[c_{\alpha}^{\ast}(k),c_{\beta}^{\ast}(k^{\prime})]=0,$ (2.21)
    ${}[c_{\alpha}(k),c_{\beta}^{\ast}(k^{\prime})]$ $=\delta_{\alpha,\beta}\big{(}\delta_{k,k^{\prime}}+\mathcal{E}_{\alpha}(k,k^{\prime})\big{)}$ (2.22)
    satisfying $\mathcal{E}_{\alpha}(k,k)\leq 0$, $\mathcal{E}_{\alpha}(k,l)=\mathcal{E}_{\alpha}(l,k)^{\ast}$ and
    $\forall\psi\in\mathcal{F}:\qquad\|\mathcal{E}_{\alpha}(k,k^{\prime})\psi\|\leq\frac{2}{n_{\alpha}(k)n_{\alpha}(k^{\prime})}\|\mathcal{N}\psi\|.$ (2.23)
  • The almost-bosonic operators change the number operator by two (see [BNP+19, Lemma 2.3]) in the following sense
    $c_{\alpha}(k)\mathcal{N}$ $=(\mathcal{N}+2)c_{\alpha}(k).$ (2.24)
  • The normalization constant satisfies (see [BNP+19, Proposition 3.1])
    $n_{\alpha}(k)^{2}=\frac{4\pi k_{\textnormal{F}}^{2}}{M}|k\cdot\hat{\omega}_{\alpha}|\big{(}1+o(1)\big{)}.$ (2.25)
Also note that the summation in the definition of the almost-bosonic operators $c_{\alpha}^{\ast}(k)$ and $c_{\alpha}(k)$ involves only finite sets. Unlike in the exactly bosonic case our almost-bosonic operators therefore inherit boundedness from the fermionic constituents which satisfy $\|a_{k}^{\ast}\|=\|a_{k}\|=1$. Subtle questions about the self-adjointness and domain of the almost-bosonic operators remain trivial in our case.
2.4 Almost-bosonic coherent state
Since $B\coloneqq c^{\ast}(\eta)-c(\eta)$ defines for all $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ a bounded operator and satisfies $B=-B^{\ast}$, the exponential operator $e^{B}$ is well-defined and is unitary.
Definition 2.1.
Define for $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ the Weyl operator
$W(\eta)\coloneqq e^{B}\coloneqq e^{c^{\ast}(\eta)-c(\eta)}.$ (2.26)
If $\eta\equiv\eta^{y}$ is additionally a bounded multiplication operator for each $y\in\mathbb{R}^{3}$, we call $W(\eta)\phi\otimes\Omega$ a coupled coherent state with $\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$.
Remark 2.2.
If $c^{\ast},c$ would satisfy the CCR without error, one could use the Baker-Campbell-Hausdorff formula to formally write
$W(\eta)\phi\otimes\Omega=$ $e^{-\|\eta\|^{2}/2}e^{c^{\ast}(\eta)}\phi\otimes\bigg{\{}1,0,\ldots,0,\ldots\bigg{\}}$
$=$ $e^{-\|\eta\|^{2}/2}e^{c^{\ast}(\eta)}\bigg{\{}\phi,0,\ldots,0,\ldots\bigg{\}}$
$=$ $e^{-\|\eta_{s}\|^{2}/2}\bigg{\{}\phi,\eta\phi,\frac{\eta^{\otimes 2}\phi}{\sqrt{2!}},\ldots,\frac{\eta^{\otimes n}\phi}{\sqrt{n!}},\ldots\bigg{\}}$ (2.27)
i.e. the coupled coherent state corresponds to a superposition of different particle number. Later on , the terms $c^{\ast}(\eta)$ and $c(\varphi)$ with the function $(\eta)_{\alpha}(k)\equiv(\eta^{y})_{\alpha}(k)\equiv e^{iky}\varphi_{\alpha}(k)$ will take the role of interaction term with an multiplication operator acting on $\phi$. Thus the occurrence of $\phi$ in each component of the vector representation stresses the coupling between the $L^{2}$ function $\phi$ and the bosonic $n$-particle state via the interaction $\eta$.
In the following we show that well-known properties of the Weyl operator (cf. for example [BPS16, Chapter 3] or [FZ17, Appendix A]) hold up to certain error terms.
Lemma 2.3 (Approximate shift property).
Let $\eta,\xi\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ and $W_{\lambda}(\eta)\coloneqq e^{\lambda B}\coloneqq e^{\lambda c^{\ast}(\eta)-\lambda c(\eta)}$ for all $\lambda\in[0,1]$, then it holds
$W_{\lambda}(\eta)^{\ast}c(\xi)W_{\lambda}(\eta)$ $=c(\xi)+\lambda\langle\xi,\eta\rangle+\langle\xi,\mathcal{R}^{\lambda}\rangle_{k,\alpha},$
$W_{\lambda}(\eta)^{\ast}c^{\ast}(\xi)W_{\lambda}(\eta)$ $=c^{\ast}(\xi)+\lambda\langle\eta,\xi\rangle+\langle\mathcal{R}^{\lambda},\xi\rangle_{k,\alpha},$
with the short-hand notation
$\langle\xi,\mathcal{R}^{\lambda}\rangle_{k,\alpha}\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\xi_{\alpha}(k)^{\ast}\mathcal{R}^{\lambda}_{\alpha}(k)$ (2.28)
and a $\lambda$-dependent error term $\mathcal{R}^{\lambda}_{\alpha}(k)\coloneqq\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}$ with $\mathcal{R}\coloneqq\mathcal{R}^{1}$.
Remark 2.4.
We will later give an estimate for $\mathcal{R}$ to show that this term corresponds indeed to a small error. Note that it holds $\mathcal{E}_{\alpha}(k,l)=\mathcal{E}_{\alpha}(l,k)^{\ast}$ for all $l,k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}$ from Lemma A.3 and therefore
$\mathcal{R}^{\lambda}_{\gamma}(l)^{\ast}\xi_{\gamma}(l)$ $=\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma^{\textnormal{nor}}}\mathcal{E}_{\gamma}(l,k)\overline{\eta_{\gamma}(k)}\xi_{\gamma}(l)\right)e^{\tau B},$ (2.29)
$\overline{\xi_{\gamma}(l)}\mathcal{R}^{\lambda}_{\gamma}(l)$ $=\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma^{\textnormal{nor}}}\overline{\xi_{\gamma}(l)}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}.$ (2.30)
For $\xi=\eta$ the above equations coincide, i.e.
$\langle\eta,\mathcal{R}^{\lambda}\rangle_{k,\alpha}=\langle\mathcal{R}^{\lambda},\eta\rangle_{k,\alpha}$ (2.31)
from which it follows immediately that $\big{(}c^{\ast}(\eta)-c(\eta)\big{)}W_{\lambda}(\eta)=W_{\lambda}(\eta)\big{(}c^{\ast}(\eta)-c(\eta)\big{)}$, i.e. $[B,W_{\lambda}(\eta)]=0$.
Proof.
The desired statement follows from Duhamel’s formula with the CCR as stated in (2.22)
$[c_{\gamma}(l),e^{\lambda B}]$ $=e^{\lambda B}\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}[c_{\gamma}(l),B]e^{\tau B}$
$=e^{\lambda B}\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\eta_{\gamma}(l)+\sum_{k\in\Gamma^{\textnormal{nor}}}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}$
$\equiv\lambda\eta_{\gamma}(l)e^{\lambda B}+e^{\lambda B}\mathcal{R}^{\lambda}_{\gamma}(l)$ (2.32)
Since $c(\xi)\equiv\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\xi_{\alpha}(k)}c_{\alpha}(k)$ is linear the result follows from the above identity. ∎
The following statement shows that $W(\eta)\phi\otimes\Omega$ corresponds to a distribution to the Fock space with expectation approximately being $2\|\eta\|^{2}$.
Proposition 2.5 (Expectation of the number operator).
Let $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$, then it holds for all $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$
$\langle W(\eta)\psi,\mathcal{N}W(\eta)\psi\rangle=2\|\eta\|^{2}+4\int_{0}^{1}\textnormal{d}\lambda\langle\zeta,\langle\eta,\mathcal{R}^{\lambda}\rangle_{k,\alpha}\zeta\rangle.$
Proof.
Using $W^{\ast}W=\textnormal{id}$ yields for all $\zeta\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\|\zeta\|=1$
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=\langle W(\eta)\zeta,[\mathcal{N},W(\eta)]\zeta\rangle+\langle\zeta,\mathcal{N}\zeta\rangle.$ (2.33)
We use Duhamel’s formula to calculate
$[\mathcal{N},e^{B}]$ $=e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}[\mathcal{N},B]e^{\tau B}$
$=2e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(c^{\ast}(\eta)+c(\eta)\right)e^{\tau B}$
$=2e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(B+2c(\eta)\right)e^{\tau B}$
$=2e^{B}B+4e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}c(\eta)e^{\tau B}$ (2.34)
where we used $[\mathcal{N},c_{\alpha}^{\ast}(k)]=2c_{\alpha}^{\ast}(k)$. Therefore
$\langle W(\eta)\zeta,[\mathcal{N},W(\eta)]\zeta\rangle$ $=2\langle\zeta,B\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\ \langle e^{\tau B}\zeta,c(\eta)e^{\tau B}\zeta\rangle$
$=2\langle\zeta,B\zeta\rangle+2\|\eta\|^{2}+4\langle\zeta,c(\eta)\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{k,\alpha}\zeta\rangle$ (2.35)
where we used the shift property Lemma 2.3 and that $e^{\tau B}$ is unitary
$\langle e^{\tau B}\zeta,c(\eta)e^{\tau B}\zeta\rangle$ $=\langle e^{\tau B}\zeta,[c(\eta),e^{\tau B}]\zeta\rangle+\langle\zeta,c(\eta)\zeta\rangle$
$=\tau\|\eta\|^{2}+\langle\zeta,\langle\eta,\mathcal{R}\rangle_{k,\alpha}\zeta\rangle+\langle\zeta,c(\eta)\zeta\rangle.$ (2.36)
Inserting (2.35) and (2.34) into (2.33) we obtain
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=2\|\eta\|^{2}+2\langle\zeta,\left(c^{\ast}(\eta)+c(\eta)\right)\zeta\rangle+\langle\zeta,\mathcal{N}\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{k,\alpha}\zeta\rangle.$ (2.37)
The desired result holds for $\zeta\equiv\phi\otimes\Omega$ since $c(\eta)\phi\otimes\Omega=0$.
For later purposes, we can bound the expectation of the number operator in the following way:
Proposition 2.6 (Stability of the number operator).
Let $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$, there exists a constant $C>0$ such that it holds for all $\tau\in[-1,1],n\in\mathbb{N}$ and $\zeta\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
$\langle e^{\tau B}\zeta,(\mathcal{N}+1)^{n}e^{\tau B}\zeta\rangle\leq e^{C\|\eta\|n|\tau|}\langle\zeta,(\mathcal{N}+3)^{n}\zeta\rangle.$
Proof.
The proof works analogously to the proof of Proposition 4.1 with a Grönwall argument and $B$ instead of $\mathbb{H}^{\textnormal{eff}}$ which is given later. Note that
$[\mathcal{N},B]=[\mathcal{N},c^{\ast}(\eta)-c(\eta)]=2\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\eta_{\alpha}(k)\big{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(k)\big{)}$ (2.38)
where we again used (2.24). The result is than obtained by using the same estimates with $\|\eta\|$ taking the role of $\|h_{y}\|$. ∎
We will later consider time-dependent functions $\eta_{t}$ which are differentiable in $t$ with $\eta_{t},\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ for each $t\in\mathbb{R}$. In this context, we are interested in the time derivative of the almost-bosonic Weyl operator $W(\eta_{t})$ which can be calculated by the following statement:
Lemma 2.7.
Let $\eta_{t}\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ be differentiable in $t$ with derivative $\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ for all $t\in\mathbb{R}$. Then it holds for all $t\in\mathbb{R}$
$\partial_{t}W(\eta_{t})=\left(c^{\ast}(\dot{\eta}_{t})-c(\dot{\eta}_{t})+i\textnormal{Im}\langle\dot{\eta}_{t},\eta_{t}\rangle\right)W(\eta_{t})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{(1-\tau)}(\eta_{t})\textnormal{Im}\langle\dot{\eta}_{t},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{t})$
with the shorthand notation $\textnormal{Im}\langle A,B\rangle_{k,\alpha}\coloneqq-\frac{i}{2}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}(A_{\alpha}^{\ast}(k)B_{\alpha}(k)-B_{\alpha}^{\ast}(k)A_{\alpha}(k))$
Proof.
For arbitrary $s\in\mathbb{R}$ it holds
$W(\eta_{s})^{\ast}\partial_{s}W(\eta_{s})$ $=\left.e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}\right|_{\tau=0}^{\tau=1}=\int_{0}^{1}\textnormal{d}\tau\ \partial_{\tau}\left(e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}\right)$
$=\int_{0}^{1}\textnormal{d}\tau\left\{-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\partial_{\tau}e^{\tau B_{s}}\right\}$ (2.39)
$=\int_{0}^{1}\textnormal{d}\tau\left\{-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\left(B_{s}e^{\tau B_{s}}\right)\right\}$ (2.40)
$=\int_{0}^{1}\textnormal{d}\tau\left\{-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}+e^{-\tau B_{s}}B_{s}\partial_{s}e^{\tau B_{s}}\right\}$ (2.41)
$=\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}.$ (2.42)
Thus
$\partial_{s}W(\eta_{s})$ $=\int_{0}^{1}\textnormal{d}\tau\ e^{(1-\tau)B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}$
$=\int_{0}^{1}\textnormal{d}\tau\ \left(\partial_{s}B_{s}\right)e^{(1-\tau)B_{s}}e^{\tau B_{s}}+\int_{0}^{1}\textnormal{d}\tau\ \left[e^{(1-\tau)B_{s}},\partial_{s}B_{s}\right]e^{\tau B_{s}}$ (2.43)
$=\left(\partial_{s}B_{s}\right)e^{B_{s}}+\int_{0}^{1}\textnormal{d}\tau\ \left[e^{(1-\tau)B_{s}},\partial_{s}B_{s}\right]e^{\tau B_{s}}.$ (2.44)
With
$\partial_{s}B_{s}=\partial_{s}\left\{c^{\ast}(\eta_{s})-c(\eta_{s})\right\}=c^{\ast}(\dot{\eta}_{s})-c(\dot{\eta}_{s})$ (2.45)
from the linearity of $c(\eta_{s})$ it follows
$\left[e^{(1-\tau)B_{s}},\partial_{s}B_{s}\right]e^{\tau B_{s}}$
=$=\left\{[W_{(1-\tau)}(\eta_{s}),c^{\ast}(\dot{\eta}_{s})]-[W_{(1-\tau)}(\eta_{s}),c(\dot{\eta}_{s})]\right\}W_{\tau}(\eta_{s})$ (2.46)
$=W_{(1-\tau)}(\eta_{s})\bigg{\{}(1-\tau)\langle\dot{\eta}_{s},\eta_{s}\rangle-(1-\tau)\langle\eta_{s},\dot{\eta}_{s}\rangle$ (2.47)
[…]
$=W(\eta_{s})(1-\tau)2i\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle+2iW_{(1-\tau)}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}).$ (2.49)
Inserting the above identity yields
$\partial_{s}W(\eta_{s})=\left(c^{\ast}(\dot{\eta}_{s})-c(\dot{\eta}_{s})+i\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{(1-\tau)}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}).$ (2.50)
3 Main results
3.1 Effective time evolution
We are now focusing on the effective time evolution of the initial state $\psi_{0}=R\psi\equiv\phi\otimes R\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$, i.e. a uncorrelated product state with the non-interacting Fermi gas prepared as its non-degenerate ground state with no initial excitations. This set-up corresponds to a system where the impurity does not interact with the cold Fermi gas at time $t=0$. Over time, we expect that the influence of the impurity particle creates and annihilates excitations of the Fermi ball. Therefore we will use the particle-hole transformation as defined in (2.3) to connect the microscopic description to the following effective description: Let
$\mathbb{H}^{\text{eff}}\coloneqq$ $-\beta\Delta_{y}+\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)+E_{N}^{\text{pw}}$
$+\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\hat{V}(k)e^{iky}n_{\alpha}(k)\bigg{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(-k)\bigg{)}$ (3.1)
be our effective Hamiltonian with $\epsilon_{\alpha}(k)=2\kappa k_{F}|k\cdot\omega_{\alpha}|$ and $E_{N}^{\text{pw}}$ as defined in (2.5). We introduce for all $k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k}$
$(h_{y})_{\alpha}(k)\coloneqq\lambda\hat{V}(k)e^{iky}n_{\alpha}(k).$ (3.2)
Note that the effective Hamiltonian acts on the components of the Hilbert space $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ in the sense that we can write
$\mathbb{H}^{\text{eff}}=(-\beta\Delta_{y})\otimes 1+1\otimes\mathbb{D}_{\text{B}}+\Phi(h_{y})+E_{N}^{\text{pw}}$ (3.3)
with $\Phi(h_{y})\coloneqq c^{\ast}(h_{y})+c(h_{y})$ and
$c^{\ast}(h_{y})$ $\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\ c_{\alpha}^{\ast}(k),$ (3.4)
$c(h_{y})$ $\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{(h_{y})_{\alpha}(k)}\ c_{\alpha}(k),$ (3.5)
$\mathbb{D}_{\text{B}}$ $\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)\quad\text{with }\epsilon_{\alpha}(k)=2\kappa k_{F}|k\cdot\omega_{\alpha}|$ (3.6)
describing the kinetic energy of the almost-bosonic pair excitations with linear dispersion relation. Note that by the Kato-Rellich theorem the effective Hamiltonian (3.1) is self-adjoint in its natural domain and generates a unitary time evolution.
To state an effective description of the time evolution of those excitations we compare particle-hole transformed microscopic dynamic of $R^{\ast}e^{-i\mathbb{H}t}R\psi$ with the effective time evolution $e^{-i\mathbb{H}^{\text{eff}}t}\psi$ in Hilbert space norm:
Theorem 3.1 (Effective dynamics of the system).
Let $(\beta,\kappa,\lambda,\tau)\in S$ as defined in (1.8) and $\epsilon>0$ small. Then it holds for the number of patches $M=N^{2\delta-\tilde{\varepsilon}}$ with $\delta=\frac{1}{6}-\tilde{\varepsilon}$ for a sufficiently small $\tilde{\varepsilon}>0$ and the initial state $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ that there is a $C_{V}>0$ such that for all $N\in\mathbb{N}$ and $t\geq 0$
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\leq C_{V}\left(e^{C_{V}\lambda k_{\textnormal{F}}t}-1\right)k_{\textnormal{F}}^{-\frac{1}{4}+\epsilon}.$
Remark 3.2.
The above result holds also in the semiclassical regime, i.e. for $\beta=k_{\textnormal{F}}^{-b}$ with $b\in\mathbb{R}$ and $\lambda=\kappa=k_{\textnormal{F}}^{-2}\in\mathcal{O}(\hbar^{2})$, in the form of
$\|R^{\ast}e^{-i\mathbb{H}t/\hbar}R\psi-e^{-i\mathbb{H}^{\text{eff}}t/\hbar}\psi\|\leq C\left(e^{Ct}-1\right)k_{\textnormal{F}}^{-\frac{1}{4}+\epsilon}.$ (3.7)
That the same estimate is obtained is based on the fact that $\lambda k_{\textnormal{F}}/\hbar$ is of order 1 in this case.
Remark 3.3.
In fact the result does hold even for a larger set then $S$.
  • One can also choose $\kappa\not=\lambda$ as long as $\kappa/\lambda+(\kappa/\lambda)^{2}\in o(k_{\textnormal{F}}^{1/4-\tilde{\varepsilon}})$. This includes in particular the choice of $\kappa\equiv 1$ and $\lambda\gtrsim k_{\textnormal{F}}^{-1/8}$ and $\tau\lesssim k_{\textnormal{F}}^{-7/8}$.
  • The coupling $\beta$ can be chosen as $\beta=k_{\textnormal{F}}^{-b}$ with any $b\in\mathbb{R}$ without changing the bound.
3.2 Effective coherent state
Our second result shows how the dynamics can be approximated on the level of states. Consider now the following state for all times $t\in\mathbb{R}$
$\psi_{t}$ $\coloneqq e^{iP(t)}W(\eta_{t})\phi\otimes\Omega$ (3.8)
with $\text{with }P(t)$ $=2\text{Im}(\nu_{t})-E_{N}^{\text{pw}}t-\text{Im}\int_{0}^{t}\textnormal{d}s\langle\dot{\eta}_{s},\eta_{s}\rangle$ (3.9)
with the choices of
$\big{(}\eta_{s}\big{)}_{\alpha}(k)\coloneqq$ $\big{(}\eta_{s}^{y}\big{)}_{\alpha}(k)\coloneqq\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}(h_{y})_{\alpha}(k)=\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}\lambda\hat{V}(k)n_{\alpha}(k)e^{iky},$ (3.10)
$\big{(}\nu_{s}\big{)}_{\alpha}(k)\coloneqq$ $\frac{e^{-is\epsilon_{\alpha}(k)}+is\epsilon_{\alpha}(k)-1}{\epsilon_{\alpha}(k)^{2}}|(h_{y})_{\alpha}(k)|^{2}$ (3.11)
for all $k\in\Gamma^{\textnormal{nor}},\alpha\in\mathcal{I}_{k}$. Due to $\big{(}\eta_{s}\big{)}_{\alpha}(k)=-ie^{-is\epsilon_{\alpha}(k)/2}\frac{\text{sin}\left(\epsilon_{\alpha}(k)s/2\right)}{\epsilon_{\alpha}(k)/2}(h_{y})_{\alpha}(k)$ and Lemma A.2 the norm is bounded for all $s\in\mathbb{R}$
$\|\eta_{s}\|$ $\equiv\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\eta_{s}\big{)}_{\alpha}(k)|^{2}\bigg{)}^{1/2}\leq\min\left\{\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}\kappa k_{\textnormal{F}}s,\sqrt{2\pi}\|\hat{V}\|_{2}\log(4\kappa k_{\textnormal{F}}s+2)\right\}$ (3.12)
as shown later in Lemma 5.1 and similarly
$\|\nu_{s}\|\equiv\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\nu_{s}\big{)}_{\alpha}(k)|^{2}\bigg{)}^{1/2}.$ (3.13)
In particular, it holds $(\eta_{0},\nu_{0})=(0,0)$ and $\lim_{\epsilon_{\alpha}(k)\to 0}(\eta_{s},\nu_{s})=(-is\varphi,0)$ and therefore $\psi_{0}=\phi\otimes\Omega$. Thus, as mentioned before, $\eta\equiv\eta^{y}$ as defined in (3.10) corresponds to an interaction term with a bounded multiplication operator acting on $L^{2}(\Lambda,\textnormal{d}y)$. The state $\psi_{t}=e^{iP(t)}W(\eta_{t})\phi\otimes\Omega$ can therefore be seen as a time-dependent coupled coherent state.
The following theorem states that $\psi_{t}$ approximately corresponds to the effective time evolution generated by $\mathbb{H}^{\textnormal{eff}}$.
Theorem 3.4 (Effective coherent dynamics).
Let $(\beta,\kappa,\lambda,\tau)\in S$ as defined in (1.8) and $\epsilon>0$ small. Consider the initial state $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\sum_{i=1}^{3}\big{(}\|\partial_{y_{i}}\phi\|+\|\partial_{y_{i}}^{2}\phi\|\big{)}\leq c<\infty$ for a constant $c>0$. Then it holds for the number of patches $M=N^{2\delta-\tilde{\varepsilon}}$ with $\delta=\frac{1}{6}-\tilde{\varepsilon}$ for a sufficiently small $\tilde{\varepsilon}>0$ that there exists a constant $C_{V}>0$ and a function $Q_{V}:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ monotonically increasing with $Q_{V}(0)=0$ such that for all $t\geq 0$
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|\leq C_{V}Q_{V}(\kappa k_{\textnormal{F}}t)\max\{\kappa k_{\textnormal{F}}^{\epsilon},\beta\}t$
with $P(t)$ given by (3.9) and $\eta_{t}$ given by (5.16).
Remark 3.5.
Note that above upper bound is indeed meaningful in the sense that the bound is small for $(\beta,\kappa,\lambda,\tau)\in S$, i.e. for $\beta t,\kappa k_{\textnormal{F}}t\in o(1)$. In particular, this includes the semiclassical regime for $\beta=k_{\textnormal{F}}^{-b}$ with $b>1$ and $\kappa=\lambda=k_{\textnormal{F}}^{-2}$
$\|e^{-i\mathbb{H}^{\text{eff}}t/\hbar}\psi-e^{iP(t/\hbar)}W(\eta_{t/\hbar})\psi\|\leq C_{V}Q_{V}(\kappa k_{\textnormal{F}}t/\hbar)t/\hbar\max\{\kappa k_{\textnormal{F}}^{\epsilon},\beta\}$
Note that also in this case $\|\eta_{t/\hbar}\|$ is of order $1$ due to $\lambda k_{\textnormal{F}}/\hbar\in\mathcal{O}(1)$. Also note that the function $Q_{V}$ is explicitly constructed in the proof.
Remark 3.6.
Note that $\psi_{t}$, as coupled coherent state, does not include the kinetic energy term $\beta(-\Delta_{y})$ of the impurity particle. This is because $\beta(-\Delta_{y})$ generates non-trivial correlations making the coherent state form generally inapplicable. However, if $\beta$ is sufficiently small, the approximation is expected to be valid. In Lemma 5.4 it is shown that $\|\Delta_{y}W(\eta_{t})\phi\otimes\Omega\|$ is of order 1 and therefore $\beta$ can be chosen to be of much larger order than $\kappa$.
Remark 3.7.
We emphasize that due to the explicit formulation of the coupled coherent state provided in (3.10), we can quantify the number of collective excitations over time. More concretely, using Proposition 2.5 the term $\|\eta_{t}\|^{2}$, which is calculated in Lemma 5.1, represents the expected number of excitations generated by the interaction with the impurity. Assuming that the impurity, together with its excitations, can be interpreted as polaron-like quasi-particle, the quantity $\|\eta_{t}\|^{2}$ offers insight into the quasi-particle formation process. As displayed in Figure 1 on page 1 the graph shows a parabolic growth followed by a logarithmic increase as qualitative feature.
Figure 1: Plot of $\|\eta_{t}\|^{2}$ using Lemma 5.1 with constant $\hat{V}$ and $\lambda=\kappa=k_{\textnormal{F}}^{-1}$.
Remark 3.8.
The proof is based on the following observation: It holds by virtue of Duhamel’s formula
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|$
$=\|\psi-e^{i\mathbb{H}^{\text{eff}}t}e^{-iE_{N}^{\text{pw}}t}e^{i2\text{Im}(\nu_{t})}e^{-i\text{Im}\int_{0}^{t}\textnormal{d}s\langle\dot{\eta}_{s},\eta_{s}\rangle}W(\eta_{s})\psi\|$ (3.14)
$=\|\int_{0}^{t}\textnormal{d}s\ e^{i\mathbb{H}^{\text{eff}}s}e^{iP(s)}\left\{\big{(}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\right\}\|$ (3.15)
$\leq\int_{0}^{t}\textnormal{d}s\|\big{(}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\|.$ (3.16)
If the collective operators $c_{\alpha}^{\ast}(k)$ and $c_{\alpha}(k)$ would be exactly bosonic, i.e. the CCR holds without error, we can use the shift property Lemma 2.3 of the Weyl operator to commute $c_{\alpha}(k)$ to the vacuum $\Omega$ with the cost of some inner product terms. In this case we observe with the short-hand notation $c^{\ast}c(\epsilon)\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)$ and by applying Lemma 2.7 that
$\big{[}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-i\left\{c^{\ast}(\dot{\eta}_{s})-c(\dot{\eta}_{s})\right\}\big{]}W(\eta_{s})\psi$
$=$ $\big{[}h_{0}+c^{\ast}c(\epsilon)+c^{\ast}(h_{y})+c(h_{y})+2\text{Im}(\dot{\nu}_{s})+c^{\ast}(-i\dot{\eta}_{s})-c(i\dot{\eta}_{s})\big{]}W(\eta_{s})\phi\otimes\Omega$
$=$ $\big{[}h_{0}+c^{\ast}(\epsilon\eta_{s})+c^{\ast}(h_{y})+\langle h_{y},\eta_{s}\rangle+2\text{Im}(\dot{\nu}_{s})+c^{\ast}(-i\dot{\eta}_{s})-\langle i\dot{\eta}_{s},\eta_{s}\rangle\big{]}W(\eta_{s})\phi\otimes\Omega$
$=$ $h_{0}W(\eta_{s})\phi\otimes\Omega$ (3.17)
is exactly vanishing up to the kinetic energy term $h_{0}$ of the impurity particle since $\eta_{t}$ as defined in (3.10) solves the ODE $\epsilon\eta_{t}+h_{y}=i\dot{\eta}_{t}$ and $\nu_{t}$ as defined in (3.11) absorbs all scalar terms.
Now by virtue of the previous statement we are able to show that the linear coupling term of the effective Hamiltonian is essential for the effective description and cannot be neglected on an approximate level.
Let
$\widetilde{\mathbb{H}^{\text{eff}}}\coloneqq\mathbb{H}^{\textnormal{eff}}-c^{\ast}(h_{y})-c(h_{y})\equiv\beta(-\Delta_{y})+\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)+E_{N}^{\text{pw}}$ (3.18)
be the effective Hamiltonian without the annihilation term from the linear coupling.
Corollary 3.9.
Let $(\beta,\kappa,\lambda,\tau)\in S$ as defined in (1.8) and $T\in\mathcal{O}(\tau)$. Under the assumptions of Theorem 3.4, there exists a monotonically increasing function $C:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ such that for all $t\in[0,T]$
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|\geq C_{\kappa,\lambda}(t)-\mathcal{O}\left(\max\{k_{\textnormal{F}}^{-\frac{1}{4}+\varepsilon},\lambda k_{\textnormal{F}}^{\varepsilon},\beta\}\right).$
In particular it holds $C_{\kappa,\lambda}(t)\in\mathcal{O}(1)$ for all $t\in[0,T]$ with $T\in\mathcal{O}(\tau)$.
Remark 3.10.
The result also holds for the semiclassical regime with $\kappa=\lambda=k_{\textnormal{F}}^{-2}$ and $t\mapsto t/\hbar$. The lower bound $C_{\kappa,\lambda}(t/\hbar)$ in this case is still of order $1$.
Remark 3.11.
A priori, it is not clear why the linear coupling term cannot be neglected. Consider the almost-bosonic annihilation operator $c(h_{y})$, for example, which consists of pairs of fermionic annihilation operators. Such an operator only gives a non-vanishing contribution when acting on a state which also contains at least one pair excitation. The occurrence of those pair excitations is unclear when considering the microscopic Hamiltonian (1.2). In particular, when comparing the non-bosonizable terms (2.9b) of the microscopic interaction term one might think that those terms should be generally the bosonizable terms (2.9a). However, the key observation is that the time evolved state is close to a coupled coherent state as shown in Theorem 3.4 which can be seen as superposition of product states of pair operators.
4 Proof of Theorem 3.1
In all of our estimates, we need to control the number operator $\mathcal{N}$ acting on the time evolved state $\psi_{t}=e^{-i\mathbb{H}^{\textnormal{eff}}t}\phi\otimes\Omega$. The following statement shows that the effective time evolution preserves the order of magnitude the number of excitations:
Proposition 4.1 (Effective time evolution of the number operator).
There exists a constant $C(\lambda,k_{\textnormal{F}})\equiv C\lambda k_{\textnormal{F}}$ such that it holds for all $t\in\mathbb{R},n\in\mathbb{N}$ and $\psi\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
$\langle e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi,(\mathcal{N}+1)^{n}e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi\rangle\leq e^{nC(\lambda,k_{\textnormal{F}})t}\langle\psi,(\mathcal{N}+3)^{n}\psi\rangle.$
Remark 4.2.
In the semiclassical regime it holds for $\lambda\equiv k_{\textnormal{F}}^{-2}\sim N^{-\frac{3}{2}}$ that $C(\lambda,k_{\textnormal{F}})\sim\hbar$ and therefore the exponential bound in
$\langle e^{-i\mathbb{H}^{\text{eff}}t/\hbar}\psi,(\mathcal{N}+1)^{n}e^{-i\mathbb{H}^{\text{eff}}t/\hbar}\psi\rangle\leq e^{nC(\lambda,k_{\textnormal{F}})t/\hbar}\langle\psi,(\mathcal{N}+3)^{n}\psi\rangle$
is of order 1 for times $t\in\mathcal{O}(1)$.
Proof.
We want to use Grönwall’s lemma and therefore estimate the derivative
$\bigg{|}i\partial_{t}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\bigg{|}$
$\leq$ $\bigg{|}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,\sum_{j=0}^{n-1}(\mathcal{N}+3)^{j}[\mathcal{N},\mathbb{H}^{\text{eff}}](\mathcal{N}+3)^{n-j-1}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\bigg{|}$
$=$ $\bigg{|}2\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\sum_{j=0}^{n-1}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{j}\big{(}c_{\alpha}^{\ast}(k)-c_{\alpha}(k)\big{)}(\mathcal{N}+3)^{n-j-1}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\bigg{|}$ (4.1)
We split the difference of $\big{(}c_{\alpha}^{\ast}(k)-c_{\alpha}(k)\big{)}$ and consider the term with $c_{\alpha}^{\ast}(k)$ first. Insert here $\textnormal{id}=(\mathcal{N}+1)^{\frac{n}{2}-1-j}(\mathcal{N}+1)^{j+1-\frac{n}{2}}$ between $(\mathcal{N}+3)^{j}$ and $c_{\alpha}^{\ast}(k)$ and use the commutation $\mathcal{N}c_{\alpha}^{\ast}(k)=c_{\alpha}^{\ast}(k)(\mathcal{N}+2)$ to obtain
$(\mathcal{N}+3)^{j}c_{\alpha}^{\ast}(k)(\mathcal{N}+3)^{n-j-1}=(\mathcal{N}+3)^{j}(\mathcal{N}+1)^{\frac{n}{2}-1-j}c_{\alpha}^{\ast}(k)(\mathcal{N}+3)^{\frac{n}{2}}.$ (4.2)
We introduce the notation $\xi_{j}\coloneqq(\mathcal{N}+1)^{\frac{n}{2}-1-j}(\mathcal{N}+3)^{j}e^{-i\mathbb{H}^{\text{eff}}t}\psi$ and $\tilde{\xi}\coloneqq(\mathcal{N}+3)^{\frac{n}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi$ to estimate
$\bigg{|}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\sum_{j=0}^{n-1}\langle\xi_{j},c_{\alpha}^{\ast}(k)\tilde{\xi}\rangle\bigg{|}$
$\leq$ $\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|(h_{y})_{\alpha}(k)|\sum_{j=0}^{n-1}\|c_{\alpha}(k)\xi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $C\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}\bigg{)}^{1/2}\sum_{j=0}^{n-1}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\xi_{j}\|^{2}\bigg{)}^{1/2}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\sum_{j=0}^{n-1}\|\mathcal{N}^{1/2}\xi_{j}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\sum_{j=0}^{n-1}\|(\mathcal{N}+1)\xi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $Cn\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle$ (4.3)
where we used Lemma A.4 and Lemma A.1 in the fourth line and the operator inequality $\mathcal{N}^{1/2}\leq(\mathcal{N}+1)$ in the fifth line. Note that $|k|<C$ for all $k\in\Gamma^{\textnormal{nor}}$.
The second term with $c_{\alpha}(k)$ can be treated by inserting $\textnormal{id}=(\mathcal{N}+1)^{\frac{n}{2}-j}(\mathcal{N}+1)^{j-\frac{n}{2}}$ and using the commutation $c_{\alpha}(k)\mathcal{N}=(\mathcal{N}+2)c_{\alpha}(k)$.
$(\mathcal{N}+3)^{j}c_{\alpha}(k)(\mathcal{N}+3)^{n-j-1}=(\mathcal{N}+3)^{\frac{n}{2}}c_{\alpha}(k)(\mathcal{N}+1)^{j-\frac{n}{2}}(\mathcal{N}+3)^{n-j-1}.$ (4.4)
We introduce the notation $\chi_{j}\coloneqq(\mathcal{N}+1)^{j-\frac{n}{2}}(\mathcal{N}+3)^{n-j-1}e^{-i\mathbb{H}^{\text{eff}}t}\psi$
$\bigg{|}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\sum_{j=0}^{n-1}\langle\tilde{\xi},c_{\alpha}(k)\chi_{j}\rangle\bigg{|}$
$\leq$ $\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|(h_{y})_{\alpha}(k)|\sum_{j=0}^{n-1}\|c_{\alpha}(k)\chi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $C\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}\bigg{)}^{1/2}\sum_{j=0}^{n-1}\bigg{(}\sum_{\alpha}\|c_{\alpha}(k)\chi_{j}\|^{2}\bigg{)}^{1/2}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\sum_{j=0}^{n-1}\|\mathcal{N}^{1/2}\chi_{j}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\sum_{j=0}^{n-1}\|(\mathcal{N}+1)\chi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $Cn\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle.$ (4.5)
Altogether it follows with the Grönwall’s lemma that
$\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\leq\exp\big{(}Cn\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}t\big{)}\langle\psi,(\mathcal{N}+3)^{n}\psi\rangle$ (4.6)
which is the desired result with since $\|\hat{V}\|_{1}<C$. ∎
The following statement shows that the fermionic kinetic energy term (2.8) can be approximated by the almost-bosonic kinetic energy term.
Proposition 4.3 (Approximation of the kinetic energy).
There exists a constant $C>0$ such that it holds for all $t\in\mathbb{R}$ and $\psi\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
$\left|\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi\|-\|(\mathbb{H}_{0}-\mathbb{D}_{B})\psi\|\right|$
$\leq C(1+\kappa\lambda^{-1})\kappa N^{\frac{1}{3}}\left(e^{C(\lambda,k_{\textnormal{F}})t}-1\right)\left(M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)$ (4.7)
with $C(\lambda,k_{\textnormal{F}})$ from Proposition 4.1.
Remark 4.4.
Note that in the case of $\psi\equiv\phi\otimes\Omega$ it holds $(\mathbb{H}_{0}-\mathbb{D}_{B})\psi=0$ and $\|(\mathcal{N}+3)^{2}\psi\|=9\leq C$. Also note that $\left(e^{C(\lambda,k_{\textnormal{F}})t}-1\right)=\lambda k_{\textnormal{F}}t+\mathcal{O}((\lambda k_{\textnormal{F}}t)^{2})$.
Remark 4.5.
Also here the above inequality holds in the semiclassical regime with $\lambda\equiv\kappa\equiv k_{\textnormal{F}}^{-2}$ in the form of
$\left|\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\textnormal{eff}}t/\hbar}\psi\|-\|(\mathbb{H}_{0}-\mathbb{D}_{B})\psi\|\right|$
$\leq C(1+\kappa\lambda^{-1})\kappa N^{\frac{1}{3}}\left(e^{C(\lambda,k_{\textnormal{F}})t/\hbar}-1\right)\left(M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)$
with $C(\lambda,k_{\textnormal{F}})t/\hbar\sim 1$ for times $t$ of order $1$.
Proof.
It holds
$\left|\partial_{t}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\right|$
$=\left|\partial_{t}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$=\left|\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,[(\mathbb{H}_{0}-\mathbb{D}_{B})^{2},\mathbb{H}^{\text{eff}}]e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$=\left|\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,\left((\mathbb{H}_{0}-\mathbb{D}_{B})[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]+[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}](\mathbb{H}_{0}-\mathbb{D}_{B})\right)e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$\leq 2\left|\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$\leq 2\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\ \|[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.8)
and make use of
$[(\mathbb{H}_{0}-\mathbb{D}_{B}),c_{\alpha}^{\ast}(k)]$ $\eqqcolon\mathfrak{E}_{\alpha}^{\text{lin}}(k)^{\ast}-\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)^{\ast}\eqqcolon\mathfrak{E}_{\alpha}(k)^{\ast},$ (4.9)
${}[(\mathbb{H}_{0}-\mathbb{D}_{B}),c_{\alpha}(k)]$ $=-\mathfrak{E}_{\alpha}(k)$ (4.10)
to arrive at
$[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]$ $=[(\mathbb{H}_{0}-\mathbb{D}_{B}),c^{\ast}c(\epsilon)+c^{\ast}(h_{y})+c(h_{y})]$ (4.11)
$=\langle\mathfrak{E},c(\epsilon)\rangle_{k,\alpha}-\langle c(\epsilon),\mathfrak{E}\rangle_{k,\alpha}+\langle\mathfrak{E},h_{y}\rangle_{k,\alpha}-\langle h_{y},\mathfrak{E}\rangle_{k,\alpha}$ (4.12)
with the short notation $\langle A,B\rangle_{k,\alpha}\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}A_{\alpha}^{\ast}(k)B_{\alpha}(k)$.
Bounds for the error terms are readily provided in Lemma A.5 and Lemma A.6 of the appendix
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|^{2}$ $\leq C\left(\kappa N^{\frac{1}{3}}M^{-\frac{1}{2}}\right)^{2}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|^{2},$ (4.13)
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$ $\leq C\left(\kappa k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2}.$ (4.14)
Furthermore use the bounds $\sum_{\alpha\in\mathcal{I}_{k}}c_{\alpha}^{\ast}(k)c_{\alpha}(k)\leq\mathcal{N}$, $\epsilon_{\alpha}(k)\leq C\kappa k_{\textnormal{F}}$ and that $\mathfrak{E}^{\textnormal{B}}_{\alpha}(k),\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k),c_{\alpha}(k)$ all annihilate exactly two fermions to estimate
$\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})\langle c(\epsilon),\mathfrak{E}^{\textnormal{lin}}-\mathfrak{E}^{\textnormal{B}}\rangle_{k,\alpha}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle$
$\leq$ $\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle\epsilon_{\alpha}(k)c_{\alpha}(k)(\mathcal{N}+1)^{-1/2}(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$+\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle\epsilon_{\alpha}(k)c_{\alpha}(k)(\mathcal{N}+1)^{-1/2}(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$ (4.15)
$\leq$ $\sum_{k\in\Gamma^{\textnormal{nor}}}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\epsilon_{\alpha}(k)c_{\alpha}(k)(\mathcal{N}+1)^{-1/2}(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}$
×$\quad\times\left\{\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}+\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}\right\}$ (4.16)
$\leq$ $C\kappa k_{\textnormal{F}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\left\{\kappa N^{\frac{1}{3}}M^{-\frac{1}{2}}\|(\mathcal{N}+1)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+\kappa k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+1)^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right\}$ (4.17)
$\leq$ $C\kappa^{2}k_{\textnormal{F}}N^{\frac{1}{3}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\left(M^{-\frac{1}{2}}\|(\mathcal{N}+1)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+1)^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right).$ (4.18)
Analogously, the term $\langle c(\epsilon),\mathfrak{E}\rangle_{k,\alpha}$ can be treated. The other terms can be estimated similarly using $|(h_{y})_{\alpha}(k)|\leq C\lambda|\hat{V}(k)||n_{\alpha}(k)|$ and Cauchy-Schwarz inequality with (A.1)
$\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})\langle\mathfrak{E}^{\textnormal{lin}}-\mathfrak{E}^{\textnormal{B}},h_{y}\rangle_{k,\alpha}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle$
$\leq$ $C\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle n_{\alpha}(k)(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$+C\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle n_{\alpha}(k)(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$ (4.19)
$\leq$ $C\lambda k_{\textnormal{F}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\Bigg{\{}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}$ (4.20)
[…]
$\leq$ $C\lambda\|\hat{V}\|_{1}k_{\textnormal{F}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\bigg{\{}\kappa M^{-\frac{1}{2}}N^{\frac{1}{3}}\|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.22)
[…]
$\leq$ $C\kappa\lambda k_{\textnormal{F}}M^{-\frac{1}{2}}N^{\frac{1}{3}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\bigg{\{}\|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.23)
[…]
Thus we derive
$\partial_{t}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$
$\leq$ $C\kappa\lambda k_{\textnormal{F}}M^{-\frac{1}{2}}N^{\frac{1}{3}}\left(\|(\mathcal{N}+1)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+1)^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right)$
$+C\lambda k_{\textnormal{F}}\kappa N^{\frac{1}{3}}\left(\|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right)$ (4.25)
$\leq$ $C\kappa^{2}k_{\textnormal{F}}N^{\frac{1}{3}}\left(e^{C(\lambda,k_{\textnormal{F}})t}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}e^{2C(\lambda,k_{\textnormal{F}})t}\|(\mathcal{N}+3)^{2}\psi\|\right)$
$+C\kappa\lambda k_{\textnormal{F}}N^{\frac{1}{3}}\left(e^{\frac{1}{2}C(\lambda,k_{\textnormal{F}})t}\|(\mathcal{N}+3)^{\frac{1}{2}}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}e^{\frac{3}{2}C(\lambda,k_{\textnormal{F}})t}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|\right)$ (4.26)
which corresponds to the integrated version
$\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|-\|(\mathbb{H}_{0}-\mathbb{D}_{B})\psi\|$
$\leq$ $C\kappa^{2}\lambda^{-1}M^{-\frac{1}{2}}N^{\frac{1}{3}}\left(e^{C(\lambda,k_{\textnormal{F}})t}-1\right)\left(\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)$
$+C\kappa M^{-\frac{1}{2}}N^{\frac{1}{3}}\left(e^{C(\lambda,k_{\textnormal{F}})t}-1\right)\left(\|(\mathcal{N}+3)^{\frac{1}{2}}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|\right)$ (4.27)
which yields the desired result. ∎
We are now ready to give the proof of the main theorem.
Proof of of Theorem 3.1.
We employ Duhamel’s formula
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ $=\|R\psi-e^{i\mathbb{H}t}Re^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.28)
$=\|\int_{0}^{t}\textnormal{d}s\ e^{i\mathbb{H}s}\big{(}\mathbb{H}R-R\mathbb{H}^{\text{eff}}\big{)}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$ (4.29)
$\leq\int_{0}^{t}\textnormal{d}s\ \|(R^{\ast}\mathbb{H}R-\mathbb{H}^{\text{eff}})e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$ (4.30)
where we used that $R$ is unitary.
From our considerations in (2.12) it follows that
$R^{\ast}\mathbb{H}R-\mathbb{H}^{\textnormal{eff}}$ $=-\beta\Delta_{y}+\mathbb{H}_{0}+b^{\ast}(\tilde{h}_{y})+b(\tilde{h}_{y})+E_{N}^{\text{pw}}+\mathcal{E}-\mathbb{H}^{\textnormal{eff}}$
$=\mathbb{H}_{0}-\mathbb{D}_{B}+b^{\ast}(\tilde{h}_{y})-c^{\ast}(h_{y})+b(\tilde{h}_{y})-c(h_{y})+\mathcal{E}.$ (4.31)
The interaction terms can be approximated in the sense of Lemma A.7:
$\|\left(b^{\ast}(\tilde{h}_{y})-c^{\ast}(h_{y})\right)\psi\|$ $\leq\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\ \|\big{(}b(k)-\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)c_{\alpha}(k)\big{)}\psi\|$
$\leq C\lambda N^{\frac{1}{3}}\|\hat{V}\|_{1}(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|.$
Therefore by combining the bounds from Proposition 4.3 with $(\mathbb{H}_{0}-\mathbb{D}_{B})\psi=0$, Lemma A.8 with $\|\hat{V}\|_{1}<C$ and Proposition 4.1 we obtain
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$
$\leq\int_{0}^{t}\textnormal{d}s\ \|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}s}\psi\|+\int_{0}^{t}\textnormal{d}s\ \|\mathcal{E}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\quad+\int_{0}^{t}\textnormal{d}s\ \|\left(b^{\ast}(\tilde{h}_{y})-c^{\ast}(h_{y})\right)e^{-i\mathbb{H}^{\text{eff}}s}\psi\|+\int_{0}^{t}\textnormal{d}s\ \|\left(b(\tilde{h}_{y})-c(h_{y})\right)e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\leq C(1+\kappa\lambda^{-1})\kappa N^{\frac{1}{3}}\left(M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)\int_{0}^{t}\textnormal{d}s\left(e^{C(\lambda,k_{\textnormal{F}})t}-1\right)$
$\quad+C\lambda\int_{0}^{t}\textnormal{d}s\ \|\mathcal{N}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\quad+C\lambda N^{\frac{1}{3}}(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\int_{0}^{t}\textnormal{d}s\ \|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\leq C(1+\kappa\lambda^{-1})\kappa\lambda^{-1}N^{\frac{1}{3}}\left(k_{\textnormal{F}}^{-1}M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)\left(e^{C\lambda k_{\textnormal{F}}t}-\lambda k_{\textnormal{F}}t-1\right)$
$\quad+Ck_{\textnormal{F}}^{-1}\|(\mathcal{N}+3)\psi\|\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)$
$\quad+Ck_{\textnormal{F}}^{-1}N^{\frac{1}{3}}(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\|(\mathcal{N}+3)^{\frac{1}{2}}\psi\|\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)$
$\leq C\tilde{C}\|(\mathcal{N}+3)^{2}\psi\|\left(e^{C\lambda k_{\textnormal{F}}t}-1\right).$ (4.32)
The prefactor is given by
$\tilde{C}$ $=\max\left\{(1+\kappa\lambda^{-1})\kappa\lambda^{-1}k_{\textnormal{F}}^{-1}N^{\frac{1}{3}}\left(M^{-\frac{1}{2}}+k_{\textnormal{F}}MN^{-1+\delta}\right),k_{\textnormal{F}}^{-1},k_{\textnormal{F}}^{-1}N^{\frac{1}{3}}(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\right\}$ (4.33)
$=N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}+\frac{a}{4}\delta}$
where we used $\kappa\equiv\lambda$ and $M=N^{a\delta}$with $22<a<\frac{2}{3\delta}-$ 2 and $\delta\in(0,\frac{1}{6})$. It holds $\tilde{C}\leq N^{-\frac{1}{12}+\varepsilon}$ for $\delta=\frac{1}{6}-\tilde{\varepsilon},a=2+\tilde{\varepsilon}$ with a sufficiently small $\tilde{\varepsilon}>0$. ∎
5 Proof of Theorem 3.4
We first collect some useful observations on the function $\eta_{s}$ in the following statement:
Lemma 5.1.
Let $\eta_{s}$ be defined as in (3.10) for $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$, then it holds for all $s\in\mathbb{R}$
$\|\eta_{s}\|^{2}$ $=\frac{\pi\lambda^{2}}{\kappa^{2}}\sum_{k\in\Gamma^{\textnormal{nor}}}\frac{\hat{V}(k)^{2}}{|k|}\left\{\log(2\kappa k_{\textnormal{F}}|k|s)-\textnormal{Ci}(2\kappa k_{\textnormal{F}}|k|s)+\gamma\right\}$
×$\times\left\{1+g(\kappa k_{\textnormal{F}}|k|s)\ \mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$
where $g:\mathbb{R}\to\mathbb{R}_{\geq 0}$ is a function independent of $k_{\textnormal{F}}$ and monotonically increasing.
Furthermore, define $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ by $(y,x)\mapsto f_{y}(x)\coloneqq\min\{e^{byx},(4x+2)^{ay}\}$ with $a\coloneqq\sqrt{2\pi}\|\hat{V}\|_{2},b\coloneqq\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}$. Then there exists a $C>0$ independent of $k_{\textnormal{F}}$ such that for $\lambda=\kappa,c_{0}>0$ and $k_{\textnormal{F}}$ sufficiently large
$\|\eta_{s}\|$ $\leq\log\left(f_{1}(\kappa k_{\textnormal{F}}s)\right),$
$e^{c_{0}\|\eta_{s}\|}$ $\leq f_{c_{0}}(\kappa k_{\textnormal{F}}s).$
Remark 5.2.
Note that $f_{y}$ is for all $y\geq 0$ monotonically increasing with $f_{y}(0)=1$. The same bounds hold also for $\eta_{s/\hbar}$ in the semiclassical regime with $\lambda=k_{\textnormal{F}}^{-2}$.
Proof.
Recall that by definition it holds
$\|\eta_{s}\|^{2}=\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\eta_{s}\big{)}_{\alpha}(k)|^{2}=\lambda^{2}\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|^{2}\underbrace{\sum_{\alpha\in\mathcal{I}_{k}}\bigg{|}n_{\alpha}(k)\frac{\sin\big{(}\epsilon_{\alpha}(k)s/2\big{)}}{\epsilon_{\alpha}(k)/2}\bigg{|}^{2}}_{\eqqcolon S_{k}}$ (5.1)
with $\epsilon_{\alpha}(k)=2\kappa k_{F}|k\cdot\omega_{\alpha}|$. We will first approximate $S_{k}$ and then give a upper bound.
Firstly, we approximate the $\alpha$-sum $S_{k}$ by an integral by identifying
$n_{\alpha}(k)^{2}=k_{\textnormal{F}}^{2}|k|\sigma(p_{\alpha})u_{\alpha}(k)^{2}\left(1+\mathcal{O}(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta})\right)$ (5.2)
with $\cos\theta_{\alpha}\coloneqq|\hat{k}\cdot\hat{\omega}_{\alpha}|\equiv u_{\alpha}(k)^{2}$ analogously to Lemma A.1. Thus, we calculate the half-sphere integral as approximation for $S_{k}$
$S_{k}$ $=\frac{1}{\kappa^{2}|k|}\int_{0}^{2\pi}\textnormal{d}\varphi\int_{0}^{\pi/2}\textnormal{d}\theta\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\sin\theta+\mathcal{E}$
$=\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(\kappa k_{\textnormal{F}}|k|su)^{2}}{u}+\mathcal{E}$
$=\frac{\pi}{\kappa^{2}|k|}\left\{\log(2\kappa k_{\textnormal{F}}|k|s)-\text{Ci}(2\kappa k_{\textnormal{F}}|k|s)+\gamma\right\}+\mathcal{E}$ (5.3)
with total error $\mathcal{E}=\mathcal{E}_{1}+\mathcal{E}_{2}+\mathcal{E}_{3}$ consisting three terms: the error $\mathcal{E}_{1}$ from (5.2), the error $\mathcal{E}_{2}$ from approximating the discrete variables $\theta_{\alpha}$ by continuous $\theta$ given by
$\mathcal{E}_{2}$ $=\left|\int_{p_{\alpha}}\textnormal{d}\sigma\ \frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}-\sigma(p_{\alpha})\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta_{\alpha})^{2}}{\cos\theta_{\alpha}}\right|$
$\leq\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\textnormal{d}}{\textnormal{d}\theta}\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\sup_{(\theta,\varphi)\in p_{\alpha}}\left|\theta-\theta_{\alpha}\right|\sigma(p_{\alpha})$
$\leq C(\kappa k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}$ (5.4)
and the error $\mathcal{E}_{3}$ from the patch construction which is given by
$\mathcal{E}_{3}=\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\left|\sigma\Big{(}\bigcup_{\alpha\in\mathcal{I}_{k}}p_{\alpha}\Big{)}-\sigma(S)\right|\leq C\kappa k_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right).$ (5.5)
Thus with the triangle inequality
$\left|S_{k}-\frac{\pi}{\kappa^{2}|k|}\left\{\log(2\kappa k_{\textnormal{F}}|k|s)-\text{Ci}(2\kappa k_{\textnormal{F}}|k|s)+\gamma\right\}\right|$
$\leq\frac{C}{\kappa^{2}|k|}\left\{M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+(\kappa k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}+\kappa k_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)\right\}.$ (5.6)
We will now give an estimate for $S_{k}$ by approximating
$\frac{\sin(\alpha x)^{2}}{x}=\frac{1}{2}\frac{1-\cos(2\alpha x)}{x}\leq\begin{cases}\alpha^{2}x&,2\alpha x\in[0,\pi/2]\\ \frac{1}{x}&,2\alpha x>\pi/2\end{cases}$
where we used $1-x^{2}/2\leq\cos(x)$ for all $x\in[0,\pi/2]$. Therefore
$\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(\kappa k_{\textnormal{F}}|k|su)^{2}}{u}$
$\leq\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(\kappa k_{\textnormal{F}}|k|s)^{2}u\ \chi(2\kappa k_{\textnormal{F}}|k|su\leq\pi/2)+\frac{1}{u}\chi(2\kappa k_{\textnormal{F}}|k|su>\pi/2)\right\}$
$\leq\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(\kappa k_{\textnormal{F}}|k|s)^{2}u\ \chi\Big{(}u\leq\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}+\frac{1}{u}\ \chi\Big{(}u>\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}\right\}$
$=\frac{2\pi}{\kappa^{2}|k|}\left\{\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2}\Big{(}\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}^{2}-\log\Big{(}\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}\right\}\ \chi\Big{(}1>\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}$
$\qquad+\frac{2\pi}{\kappa^{2}|k|}\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}1\leq\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}$
$=\frac{2\pi}{\kappa^{2}|k|}\left\{\frac{\pi^{2}}{32}-\log(\pi)+\log(4\kappa k_{\textnormal{F}}|k|s)\right\}\ \chi\Big{(}\kappa k_{\textnormal{F}}|k|s>\frac{\pi}{4}\Big{)}$
$\qquad+\frac{2\pi}{\kappa^{2}|k|}\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}\kappa k_{\textnormal{F}}|k|s\leq\frac{\pi}{4}\Big{)}$
$\leq\frac{2\pi}{\kappa^{2}|k|}\min\left\{\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2},\log(4\kappa k_{\textnormal{F}}|k|s+1)\right\}.$ (5.7)
Note that $\log(4\kappa k_{\textnormal{F}}|k|s+1)\leq\log(4\kappa k_{\textnormal{F}}s+1)+\log|k|$ using $\frac{1}{2}x^{2}\leq\log(4x+1)$ for all $x\leq 2$ and $|k|\geq 1$ for all $k\in\mathbb{Z}_{\ast}^{3}$. Thus, for $\lambda=\kappa$
$\|\eta_{s}\|^{2}$ $\leq 2\pi\sum_{k\in\mathbb{Z}_{\ast}^{3}}|\hat{V}(k)|^{2}\min\left\{\frac{(\kappa k_{\textnormal{F}}s)^{2}}{2}|k|,\frac{\log(4\kappa k_{\textnormal{F}}|k|s+1)}{|k|}\right\}$
$\leq 2\pi\min\left\{\|\hat{V}(\cdot)^{1/2}\|_{2}^{2}(\lambda k_{\textnormal{F}}s)^{2}/2,\|\hat{V}\|_{2}^{2}\left(\log(4\kappa k_{\textnormal{F}}s+1)+1\right)\right\}.$ (5.8)
Thus by using $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ and $\sqrt{\log(x+1)}\leq\log(x+2)$ we obtain the desired
$\|\eta_{s}\|$ $\leq\min\left\{b\kappa k_{\textnormal{F}}s,\log(4\kappa k_{\textnormal{F}}s+2)^{a}+a\right\},$ (5.9)
$e^{c_{0}\|\eta_{s}\|}$ $\leq\min\left\{e^{c_{0}b\kappa k_{\textnormal{F}}s},e^{c_{0}a}(4\kappa k_{\textnormal{F}}s+2)^{c_{0}a}\right\}$ (5.10)
with $a\coloneqq\sqrt{2\pi}\|\hat{V}\|_{2},b\coloneqq\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}.$
Lemma 5.3.
It holds for all $s\in\mathbb{R}$, $n\in\mathbb{N}$, $\psi\in\mathcal{F}$
  1. (i).
    $\sum_{k\in\Gamma^{\textnormal{nor}}}\||k|^{n}\eta_{s}(k)\|_{l^{2}}\leq C\|\eta_{s}\|$,
  2. (ii).
    $\langle\eta_{s},|k|^{n}\eta_{s}\rangle\leq C\|\eta_{s}\|^{2}$,
  3. (iii).
    $\|c^{\ast}(|k|^{n}\eta_{s})\psi\|\leq C\|\eta_{s}\|\ \|(\mathcal{N}+1)^{1/2}\psi\|$.
Proof.
For the inequalities, we observe that for $n\in\mathbb{N}_{0}$ it holds
$\sum_{k\in\Gamma^{\textnormal{nor}}}\||k|^{n}\eta_{s}(k)\|_{l^{2}}$ $\equiv\sum_{k\in\Gamma^{\textnormal{nor}}}|k|^{n}\Big{(}\sum_{\alpha\in\mathcal{I}_{k}}|(\eta_{s})_{\alpha}(k)|^{2}\Big{)}^{1/2}=\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|k|^{n}|\hat{V}(k)|\left(S_{k}\right)^{1/2}.$ (5.11)
Since by assumption $\|(\cdot)^{n}\hat{V}\|_{1}$ is bounded for each $n\in\mathbb{N}$, the first statement follows. The second statement simply follows from the same calculation and recalling that $\langle\eta_{s},|k|^{n}\eta_{s}\rangle\equiv\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}|\eta_{\alpha}(k)|^{2}$. The third statement follows from Lemma A.4 and
$\|c^{\ast}(|k|^{n}\eta_{s})\psi\|\leq\sum_{k\in\Gamma^{\textnormal{nor}}}\|\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}\eta_{\alpha}(k)c_{\alpha}^{\ast}(k)\psi\|\leq\sum_{k\in\Gamma^{\textnormal{nor}}}\||k|^{n}\eta_{s}(k)\|_{l^{2}}\|(\mathcal{N}+1)^{1/2}\psi\|.$ (5.12)
We will now give the proof of the second main theorem.
Proof of Theorem 3.4.
We use the approach as sketched in 3.8. Since the bosonic property only holds only with an error, the equality (3.17) holds only approximately:
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|\leq\int_{0}^{t}\textnormal{d}s\|h_{0}W(\eta_{s})\phi\otimes\Omega\|+\text{Error}_{1}+\text{Error}_{2}.$ (5.13)
We will first estimate the error terms and then subsequently treat the $h_{0}$ term in a separate lemma. We give an explicit expression for the first error term using Lemma 2.3 on the approximate shift property applied to $c_{\alpha}(k)$ in the $c^{\ast}c(\epsilon),c(h_{y})$ and $c(i\dot{\eta}_{s})$ terms. Thus, the error of (3.17) is given by
$\text{Error}_{1}\coloneqq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\left\{\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right\}W(\eta_{s})\mathcal{R}_{\alpha}(k)\psi.$ (5.14)
The second error term is given by Lemma 2.7 on the time derivative of the almost-bosonic Weyl operator and therefore
$\text{Error}_{2}\coloneqq-$ $2i\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})\psi$
$\equiv$ $2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau e^{(1-\tau)B}\bigg{\{}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}e^{is\epsilon_{\alpha}(k)}\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)$
$+\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{-is\epsilon_{\alpha}(k)}(h_{y})_{\alpha}(k)\bigg{\}}e^{\tau B}\psi.$ (5.15)
Firstly, we show that the term $\mathcal{R}^{\lambda}_{\alpha}(k)\psi=\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi$ as defined in Lemma 2.3 constitutes indeed a small error. We estimate
$\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{\gamma\in\mathcal{I}_{l}}\|\mathcal{R}_{\gamma}(l)\psi\|^{2}$
$\leq\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{\gamma\in\mathcal{I}_{l}\cap\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}|\eta_{\gamma}(k)|\int_{0}^{1}\textnormal{d}\tau\|\mathcal{E}_{\gamma}(k,l)e^{\tau B}\psi\|\bigg{)}^{2}$
$\leq C\left(MN^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)\psi\|\right)^{2}$ (5.16)
where we used $\sum_{\alpha\in\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(k)|\bigg{)}^{2}\leq\sum_{k^{\prime},k\in\Gamma^{\textnormal{nor}}}\|\eta(k)\|_{l^{2}}\|\eta(k^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.1 and
$\int_{0}^{1}\textnormal{d}\tau\|\mathcal{E}_{\gamma}(k,l)e^{\tau B}\psi\|$ (5.17)
$\leq\int_{0}^{1}\textnormal{d}\tau\langle e^{\tau B}\psi,|\mathcal{E}_{\gamma}(k,l)|^{2}e^{\tau B}\psi\rangle^{1/2}\leq CMN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\langle e^{\tau B}\psi,\mathcal{N}^{2}e^{\tau B}\psi\rangle^{1/2}$ (5.18)
$\leq CMN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau e^{C\|\eta_{s}\|\tau}\|(\mathcal{N}+3)\psi\|\leq C\|\eta_{s}\|^{-1}(e^{C\|\eta_{s}\|}-1)MN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)\psi\|$ (5.19)
which follows from $e^{\tau B}$ is unitary in the first inequality, Lemma A.3 in the second inequality and Proposition 2.6 in the third inequality.
Secondly, we estimate
$\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\leq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\|c_{\alpha}^{\ast}(k)e^{(1-\tau)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|c_{\alpha}^{\ast}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
$\quad+\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|[c_{\alpha}^{\ast}(k),e^{(1-\tau)B}]\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\bigg{(}\|(\mathcal{N}+1)^{\frac{3}{2}}e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}+C\|\eta_{s}\|^{2}MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ (1-\tau)\|\mathcal{N}e^{\tau B}\psi\|$
$\quad+\sum_{k\in\Gamma^{\textnormal{nor}}}\|\eta_{s}\|\int_{0}^{1}\textnormal{d}\tau\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|\mathcal{R}_{\alpha}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}$
$\leq CMN^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{\frac{3}{2}}\psi\|+CMN^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-C\|\eta_{s}\|-1)\|(\mathcal{N}+3)\psi\|$
$\quad+CM^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)^{2}\|(\mathcal{N}+3)^{2}\psi\|$
$\leq CM^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{2}\psi\|.$ (5.20)
where we used $[c_{\alpha}^{\ast}(k),e^{\lambda B}]=-\lambda\eta_{\alpha}(k)e^{\lambda B}-e^{\lambda B}\mathcal{R}_{\alpha}(k)$ from Lemma 2.3, the Cauchy-Schwarz inequality for the $\alpha$-summation, Lemma A.3 in the third inequality and in the fourth inequality we used Proposition 2.6 and (5.16).
In total by combining (5.16) and (5.20) we end up with the following estimate
$\|\text{Error}_{1}\|$ $\leq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\left\{\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right\}e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\leq C\kappa k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\quad+C\lambda\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\|n_{\alpha}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\leq C\kappa k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\quad+C\lambda k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\|\hat{V}\|_{2}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}_{\alpha}(k)\psi\|^{2}\bigg{)}^{1/2}$
$\leq C\kappa k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\ (e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{2}\psi\|$
$\quad+C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\ (e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)\psi\|$
$\leq C\ (f_{C}(\kappa k_{\textnormal{F}}t)-1)(\lambda+\kappa)k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)^{2}\psi\|.$ (5.21)
where we used (A.2) and $e^{B}$ unitary in the third inequality and Lemma 5.1 in the last line. Using the choices of $\kappa=\lambda$ and $\psi=\phi\otimes\Omega$ we obtain the desired bound.
Similarly, we obtain an estimate for the second error term using Cauchy-Schwarz, (5.16) and Proposition 2.6
$\|\text{Error}_{2}\|$
$\leq$ $2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\bigg{\{}\|\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)e^{\tau B}\psi\|\bigg{\}}$
$\leq$ $2\lambda\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|\hat{V}(k)n_{\alpha}(k)|\bigg{\{}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|\bigg{\}}$
$\leq$ $C\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ \lambda k_{\textnormal{F}}\|\hat{V}\|_{2}\bigg{\{}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}+\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}\bigg{\}}$
$\leq$ $C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|(1-\tau)}-1)\|(\mathcal{N}+3)e^{\tau B}\psi\|$
$\leq$ $C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+5)\psi\|$
$\leq$ $C(f_{C}(\kappa k_{\textnormal{F}}t)-1)\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+5)\psi\|.$ (5.22)
Again, using the choice of $\kappa=\lambda$ and $\psi=\phi\otimes\Omega$ we obtain the desired bound.
With the subsequent Lemma 5.4, we can conclude with a bound on $h_{0}$ of the form
$\int_{0}^{t}\textnormal{d}s\|h_{0}W(\eta_{s})\phi\otimes\Omega\|\leq C\beta\int_{0}^{t}\textnormal{d}s\left\{(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)\right\}$
$\leq C\beta t\left\{\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]^{4}\right\}\left\{f_{C}(\kappa k_{\textnormal{F}}t)+1\right\}.$ (5.23)
[…]
Lemma 5.4.
Under the assumptions of Theorem 3.4, it holds that for all $t\geq 0$
$\|\Delta_{y}W(\eta_{t})\phi\otimes\Omega\|\leq C(\|\eta_{t}\|+\|\eta_{t}\|^{4})(e^{C\|\eta_{t}\|}+1)$
for $C>0$ independent of $k_{\textnormal{F}}$.
Proof.
We explicitly calculate the action of the Laplacian on the coupled coherent state $W(\eta_{s})\psi=e^{B}\psi$ i.e.
$-\Delta_{y}W(\eta_{s})\psi=-\big{(}\Delta_{y}W(\eta_{s})\big{)}\psi-2\beta\nabla_{y}W(\eta_{s})\cdot\nabla_{y}\psi-W(\eta_{s})\Delta_{y}\psi.$ (5.24)
In total we expect, that all terms can be bounded by assumption on the initial condition on $\phi$. We will first focus on the term $\Delta_{y}W(\eta_{s})$. It holds
$W(\eta_{s})^{\ast}\partial_{y_{i}}W(\eta_{s})$ $=\left.e^{-\tau B_{s}}\partial_{y_{i}}e^{\tau B}\right|_{\tau=0}^{\tau=1}=\int_{0}^{1}\textnormal{d}\tau\ \partial_{\tau}\left(e^{-\tau B}\partial_{y_{i}}e^{\tau B}\right)$
$=\int_{0}^{1}\textnormal{d}\tau\left\{-Be^{-\tau B}\partial_{y_{i}}e^{\tau B}+e^{-\tau B}\partial_{y_{i}}\left(B_{s}e^{\tau B}\right)\right\}$
$=\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(\partial_{y_{i}}B_{s}\right)e^{\tau B}.$ (5.25)
With
$\partial_{y_{i}}B_{s}=$ $\partial_{y_{i}}\left\{c^{\ast}(\eta_{s})-c(\eta_{s})\right\}=c^{\ast}(\partial_{y_{i}}\eta_{s})-c(\partial_{y_{i}}\eta_{s}),$ (5.26)
$\partial_{y_{i}}\eta_{s}=$ $\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}\lambda\hat{V}(k)n_{\alpha}(k)ik_{i}e^{iky}=ik_{i}\eta_{s}$ (5.27)
it follows analogously to Lemma 2.7 that
$\partial_{y_{i}}W(\eta_{s})$ $=\int_{0}^{1}\textnormal{d}\tau\ e^{(1-\tau)B_{s}}\left(\partial_{y_{i}}B_{s}\right)e^{\tau B_{s}}$
$=\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})$
$\hphantom{=}+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}).$ (5.28)
And repeating the differentiation with
$\partial_{y_{i}}W_{\tau}(\eta_{s})$ $=\tau\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\tau\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W_{\tau}(\eta_{s})$
$\hphantom{=}+2i\int_{0}^{\tau}\textnormal{d}\sigma\ W_{1-\sigma}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\tau-\sigma}\rangle_{k,\alpha}W_{\sigma}(\eta_{s})$ (5.29)
yields
$\Delta W(\eta_{s})=\sum_{i=1}^{3}\partial_{y_{i}}^{2}W(\eta_{s})$
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})$
$+\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\partial_{y_{i}}W(\eta_{s})$
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}$ (5.30)
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})$
$+\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\times$
×$\times\left\{\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})\right\}$
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}$ (5.31)
$\eqqcolon$ $I_{1}+I_{2}+I_{3}+I_{4}$
with
$I_{1}$ $\coloneqq\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})$
$\hphantom{\coloneqq}-\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s}),$ (5.32)
$I_{2}$ $\coloneqq 2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}),$ (5.33)
$I_{3}$ $\coloneqq 2\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\int_{0}^{1}\textnormal{d}\tau\ iW_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}),$ (5.34)
$I_{4}$ $\coloneqq 2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}.$ (5.35)
We show that each term can be bounded here by a constant at most of order $1$.
For $I_{1}$, we can treat all $c^{\ast}(\cdots)$ and $c(\cdots)$ terms with Lemma 5.1. Furthermore we use that $c_{\alpha}(k)\mathcal{N}=(\mathcal{N}+2)c_{\alpha}(k)$ to estimate
$\|I_{1}\psi\|$ $\leq C\|\eta_{s}\|\ \|(\mathcal{N}+1)^{1/2}W(\eta_{s})\psi\|+C(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})\|W(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|(\mathcal{N}+3)W(\eta_{s})\psi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}\|(\mathcal{N}+5)\psi\|+C(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})$ (5.36)
where we used Proposition 2.6.
For $I_{2}$, we use a similar approach to (5.16) to obtain
$\|I_{2}\psi\|$ $\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|(1-\tau)}-1)\|(\mathcal{N}+3)W_{\tau}(\eta_{s})\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|}e^{(\tilde{C}-C)\|\eta_{s}\|\tau}-e^{\tilde{C}\|\eta_{s}\|\tau})\|(\mathcal{N}+5)\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)\psi\|.$ (5.37)
For $I_{3}$, we first observe that for $n\in\mathbb{N}$
$\|\mathcal{N}^{n}\mathcal{R}^{\mu}_{\beta}(k)\psi\|$ $\leq\int_{0}^{\mu}\textnormal{d}\tau\ \|\mathcal{N}^{n}e^{-\tau B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\beta}(l)\mathcal{E}_{\beta}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\beta}(l)|\ \|\mathcal{E}_{\beta}(l,k)(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\beta}(l)|\ MN^{-\frac{2}{3}+\delta}\|\mathcal{N}(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\|\eta_{s}\|^{-1}(e^{C\|\eta_{s}\|\mu}-1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\beta}(l)|\ \|(\mathcal{N}+5)^{n+1}\psi\|.$ (5.38)
We use a similar approach to (5.20) and insert the above inequality (5.38) to obtain
$\|I_{3}\psi\|$ $\leq C\|\eta_{s}\|\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\left(\|\eta_{s}\|(\mathcal{N}+1)^{1/2}+C\|\eta_{s}\|^{2}\right)W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\mathcal{N}W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ e^{C\|\eta_{s}\|(1-\tau)}\|\mathcal{N}\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|^{2}MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|(1-\tau)}-1)\|(\mathcal{N}+5)^{2}\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)^{2}\psi\|$
$\quad+CMN^{-\frac{2}{3}+\delta}\|\eta_{s}\|^{2}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}(\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)^{2}\psi\|.$ (5.39)
For $I_{4}$, we first calculate
$\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}$
$=\int_{\tau}^{1}\textnormal{d}\sigma\ \partial_{y_{i}}e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}$
$\hphantom{=}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\partial_{y_{i}}e^{\sigma B}$ (5.40)
$\eqqcolon I_{4,1}+I_{4,2}+I_{4,3}+I_{4,4}+I_{4,5}$
with
$I_{4,1}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i(1-\sigma)\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)$
×$\times e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.41)
$I_{4,2}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\sigma-a}\rangle_{k,\alpha}e^{aB}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.42)
$I_{4,3}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.43)
$I_{4,4}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\sigma\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B},$ (5.44)
$I_{4,5}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{k,\alpha}e^{aB}.$ (5.45)
We approach each term similarly to (5.16).
For $I_{4,1}$, it holds
$\|I_{4,1}\psi\|$
$\leq C\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\|\left(\|\eta_{s}\|\ \|(\mathcal{N}+1)e^{(1-\sigma)B}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|^{2}\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\|\left(e^{C\|\eta_{s}\|(1-\sigma)}\|(\mathcal{N}+3)\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|\,\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)\|(\mathcal{N}+3)^{2}e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+5)^{2}\psi\|$ (5.46)
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\left((1-\tau)e^{C\|\eta_{s}\|(1-\tau)}+C\|\eta_{s}\|e^{C\|\eta_{s}\|}\right)\|(\mathcal{N}+5)^{2}\psi\|$ (5.47)
where we used Lemma A.4 and Lemma 5.1 in the first inequality, Lemma A.3 in the third inequality and Proposition 2.6 in the second and forth inequality. Therefore using $\int_{0}^{1}(1-\tau)e^{y(1-\tau)}\textnormal{d}\tau=y^{-2}\left((y-1)e^{y}+1\right)$ yields
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,1}\psi\|^{2}}$ (5.48)
$\leq CMN^{-\frac{2}{3}+\delta}\left((\|\eta_{s}\|^{3}+\|\eta_{s}\|-1)e^{C\|\eta_{s}\|}+1\right)\|(\mathcal{N}+5)^{2}\psi\|$ (5.49)
where we used $\sum_{\alpha\in\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\bigg{)}^{2}\leq\sum_{l^{\prime},l\in\Gamma^{\textnormal{nor}}}\|\eta(l)\|_{l^{2}}\|\eta(l^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.1.
For $I_{4,2}$, it holds
$\|I_{4,2}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ \|(\mathcal{N}+3)e^{aB}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|a}\|(\mathcal{N}+5)^{2}e^{\sigma B}\psi\|$
$\leq C\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma(e^{C\|\eta_{s}\|(1-\sigma)}-1)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+8)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|e^{C\|\eta_{s}\|(1-\tau)}\|(\mathcal{N}+8)^{2}\psi\|$ (5.50)
where we used Cauchy-Schwarz with (5.16) and Lemma 5.1 in the first inequality, Proposition 2.6 and Lemma A.3 in the second inequality. Therefore it follows
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,2}\psi\|^{2}}$
$\leq CM^{2}N^{-\frac{4}{3}+2\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+8)^{2}\psi\|$ (5.51)
The term $I_{4,3}$ is estimated similarly to (5.16) by
$\|I_{4,3}\psi\|\leq C\|\eta_{s}\|^{-1}MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|(e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+3)\psi\|$ (5.52)
and therefore
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,3}\psi\|^{2}}\leq CMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)\psi\|.$ (5.53)
Similarly for $I_{4,4}$, we estimate
$\|I_{4,4}\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|\mathcal{N}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{2})MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|(\mathcal{N}+1)^{2}e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\ \left(e^{C\|\eta_{s}\|}+(1-C\|\eta_{s}\|\tau)e^{C\|\eta_{s}\|\tau}\right)\|(\mathcal{N}+3)^{2}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\ e^{C\|\eta_{s}\|}\|(\mathcal{N}+3)^{2}\psi\|$ (5.54)
using $\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma e^{y\sigma}=y^{-2}\left((y-1)e^{y}-(\tau y-1)e^{y\tau}\right)$ and therefore
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,4}\psi\|^{2}}$
$\leq CMN^{-\frac{2}{3}+\delta}(\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}\|(\mathcal{N}+1)^{2}\psi\|.$ (5.55)
For $I_{4,5}$, we estimate
$\|I_{4,5}\psi\|$ $=\|2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{k,\alpha}e^{aB}\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ \|\mathcal{N}e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{k,\alpha}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|(\mathcal{N}+3)\mathcal{R}^{\sigma-a}_{\alpha}(k)e^{aB}\psi\|^{2}\bigg{)}^{1/2}$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}(e^{C\|\eta_{s}\|(\sigma-a)}-1)\|(\mathcal{N}+8)^{2}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|}e^{C^{\prime}\|\eta_{s}\|\sigma}e^{-C^{\prime\prime}\|\eta_{s}\|a}\|(\mathcal{N}+10)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\ (e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+10)^{2}\psi\|.$ (5.56)
where we used (5.38) in the third inequality. Therefore we obtain
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,5}\psi\|^{2}}\leq CM^{2}N^{-\frac{4}{3}+2\delta}\|(\mathcal{N}+10)^{2}\psi\|.$ (5.57)
Taking the five bounds together we finally obtain
$\|I_{4}\psi\|$ $=\|2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}\psi\|$
$\leq C\int_{0}^{1}\textnormal{d}\tau\sum_{n=1}^{5}\|\eta_{s}\|\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,n}\psi\|^{2}\bigg{)}^{1/2}$
$\leq CMN^{-\frac{2}{3}+\delta}\left((\|\eta_{s}\|^{3}+\|\eta_{s}\|-1)e^{C\|\eta_{s}\|}+1\right)\|(\mathcal{N}+8)^{2}\psi\|$ (5.58)
Combining (5.36), (5.37), (5.39) and (5.58) with $\psi\equiv\phi\otimes\Omega$ yields the desired final result of
$\|\big{(}\Delta_{y}W(\eta_{s})\big{)}\psi\|\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1).$ (5.59)
Similarly to (5.36) and (5.37), we further estimate the second term of (5.24)
$\|\nabla_{y}W(\eta_{s})\cdot\nabla_{y}\psi\|$
$=$ $\|\sum_{i=1}^{3}\Big{\{}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})$
[…]
$\leq$ $C\left((\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}+(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})\right)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|+CMN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|.$ (5.60)
Therefore in total for all $t\in\mathbb{R}$
$\|\Delta_{y}W(\eta_{t})\phi\otimes\Omega\|$ (5.61)
$\leq C\left\{(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)+\sum_{i=1}^{3}\big{(}(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)\|\partial_{y_{i}}\phi\|+\|\partial_{y_{i}}^{2}\phi\|\big{)}\right\}.$ (5.62)
6 Proof of Corollary 3.9
Proof of Corollary 3.9.
We observe that by the inverse triangle inequality it holds
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|$
$\geq\|e^{iP(t)}W(\eta_{t})\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|$ (6.1a)
$\hphantom{\geq}-\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|-\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|.$ (6.1b)
The second line (6.1b) is be bounded from below by bounds of order $o(1)$ from Theorem 3.1 and Theorem 3.4. Therefore it is sufficient to show that the first line (6.1a) is large.
The first line (6.1a) can be explicitly estimated by the same approach as in the proof of Theorem 3.4. That is use Duhamel’s formula, commute all $c_{\alpha}(k)$-operators with $W(\eta_{s})$ and collect the error terms via Lemma 2.3:
$\bigg{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})-1\bigg{)}\psi$
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\left\{\big{(}\widetilde{\mathbb{H}^{\text{eff}}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\eta_{s},\dot{\eta}_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\right\}$ (6.2)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{\{}\big{(}h_{0}-c^{\ast}(h_{y})-c(h_{y})\big{)}W(\eta_{s})\psi+\text{Error}\bigg{\}}$ (6.3)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{\{}-W(\eta_{s})\left(c^{\ast}(h_{y})+\langle\eta_{s},h_{y}\rangle+\langle h_{y},\eta_{s}\rangle\right)\psi+h_{0}W(\eta_{s})\psi+\widetilde{\text{Error}}\bigg{\}}$ (6.4)
with $\int_{0}^{t}\textnormal{d}s\ \text{Error}=\text{Error}_{1}+\text{Error}_{2}$ from (5.13) and where we also commuted $c^{\ast}(h_{y})$ and $c(h_{y})$ with $W(\eta_{s})$ such that the new error term is of the form
$\widetilde{\text{Error}}$ $\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\left(\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)-e^{is\epsilon_{\alpha}(k)}\right)\overline{(h_{y})_{\alpha}(k)}W(\eta_{s})\mathcal{R}_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}W(\eta_{s})\mathcal{R}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})\psi$ (6.5)
Note that $\widetilde{\text{Error}}$ is $s$-dependent even though we did not include the dependence explicitly in the notation.
We employ the Cauchy-Schwarz inequality, insert our previous finding with the triangle inequality and use that $e^{-i\widetilde{\mathbb{H}^{\text{eff}}}s}$ is unitary to estimate
$\|e^{iP(t)}W(\eta_{t})\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|$
$=\|\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})-1\big{)}\psi\|\geq\left|\langle\psi,\big{(}1-e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})\big{)}\psi\rangle\right|$
$=\left|\int_{0}^{t}\textnormal{d}s\langle e^{-iP(s)}e^{-i\widetilde{\mathbb{H}^{\text{eff}}}s}\psi,\big{(}\widetilde{\mathbb{H}^{\text{eff}}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\eta_{s},\dot{\eta}_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\rangle\right|$
$\geq\left|\int_{0}^{t}\textnormal{d}s\langle\psi,e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})\left(c^{\ast}(h_{y})+2\text{Re}\langle h_{y},\eta_{s}\rangle\right)\psi\rangle\right|$
$\hskip 113.81102pt-\int_{0}^{t}\textnormal{d}s\left|\langle e^{-iP(s)}e^{-i\widetilde{\mathbb{H}^{\text{eff}}}s}\psi,\widetilde{\text{Error}}+h_{0}W(\eta_{s})\psi\rangle\right|$ (6.6)
$\geq\left|\int_{0}^{t}\textnormal{d}s\langle\psi,\big{(}1+e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})-1\big{)}\left(c^{\ast}(h_{y})+2\text{Re}\langle h_{y},\eta_{s}\rangle\right)\psi\rangle\right|$
$\hskip 113.81102pt-\int_{0}^{t}\textnormal{d}s\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$ (6.7)
$\geq\left|\int_{0}^{t}\textnormal{d}s\left\{\langle\psi,c^{\ast}(h_{y})\psi\rangle+2\text{Re}\langle h_{y},\eta_{s}\rangle\right\}\right|$
$\hphantom{\geq}\qquad-\left|\int_{0}^{t}\textnormal{d}s\langle\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})-1\big{)}^{\ast}\psi,\left(c^{\ast}(h_{y})+2\text{Re}\langle h_{y},\eta_{s}\rangle\right)\psi\rangle\right|$
$\hphantom{\geq}\qquad\qquad-\int_{0}^{t}\textnormal{d}s\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$ (6.8)
$\geq 2\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\langle h_{y},\eta_{s}\rangle\right|-\big{(}\|c^{\ast}(h_{y})\psi\|+2\sup_{s\in[0,t]}|\text{Re}\langle h_{y},\eta_{s}\rangle|\big{)}\int_{0}^{t}\textnormal{d}s\|\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})-1\big{)}\psi\|$
$\hphantom{\geq}\qquad\qquad-t\sup_{s\in[0,t]}\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$ (6.9)
where we used $\langle\psi,c^{\ast}(h_{y})\psi\rangle=0$ since $c(h_{y})\phi\otimes\Omega=0$.
The three parts of the above inequality (6.9) can be bounded in the following:
Firstly, the error term can be bounded with (5.23) and analogously to (5.21) and (5.22):
$\sup_{s\in[0,t]}\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$
$\leq C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}(f_{C}(\kappa k_{\textnormal{F}}t)-1)+C\beta\left(\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]^{4}\right)\left(f_{C}(\kappa k_{\textnormal{F}}t)+1\right)$
$\leq CQ_{V}(\kappa k_{\textnormal{F}}t)\max\{\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta},\beta\}\eqqcolon d\in\mathcal{O}(\max\{\lambda k_{\textnormal{F}}^{\epsilon},\beta\})$ (6.10)
using that for all $t\in[0,T]$ with $T\in\mathcal{O}(\tau)$ it holds $\kappa k_{\textnormal{F}}t\in\mathcal{O}(1)$.
Secondly, the prefactor of the integral term is bounded by Lemma A.4
$\|c^{\ast}(h_{y})\psi\|+2\sup_{s\in[0,t]}|\text{Re}\langle h_{y},\eta_{s}\rangle|$
$\leq\|h_{y}\|\|(\mathcal{N}+1)^{1/2}\psi\|+2\sup_{s\in[0,t]}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\frac{1-\cos(s\epsilon_{\alpha}(k))}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}$ (6.11)
$\leq\|h_{y}\|+\frac{2}{\kappa k_{\textnormal{F}}}\|h_{y}\|^{2}\leq C\lambda k_{\textnormal{F}}\big{(}\|\hat{V}\|+\kappa^{-1}\lambda\|\hat{V}\|^{2}\big{)}\eqqcolon\alpha.$ (6.12)
Thirdly, we want to show that the remaining term $|\int_{0}^{t}\textnormal{d}s\,\text{Re}\langle h_{y},\eta_{s}\rangle|$ is large:
$\left|\int_{0}^{t}\textnormal{d}s\,\text{Re}\langle h_{y},\eta_{s}\rangle\right|$ $=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}\right|$
$=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\lambda^{2}\hat{V}(k)^{2}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}\right|$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\frac{e^{-i2\kappa k_{\textnormal{F}}|k|s}-1}{2\kappa k_{\textnormal{F}}|k|s}\right|\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$ (6.13)
where we used that the sum over $\alpha$ is approximated similarly to (A.2) by integrating over the half sphere
$\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}$ $=2\pi\int_{0}^{\pi/2}\textnormal{d}\theta\ \frac{e^{-is2\kappa k_{\textnormal{F}}|k|\cos\theta}-1}{2\kappa k_{\textnormal{F}}|k|\cos\theta}\cos\theta\sin\theta\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$=i\frac{2\pi k_{\textnormal{F}}^{2}}{2\kappa k_{\textnormal{F}}}\frac{e^{-i2\kappa k_{\textnormal{F}}|k|s}-1}{2\kappa k_{\textnormal{F}}|k|s}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$ (6.14)
Therefore we obtain
$\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\frac{e^{-i2\kappa k_{\textnormal{F}}|k|s}-1}{2\kappa k_{\textnormal{F}}|k|s}\right|=\left|\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\int_{0}^{t}\textnormal{d}s\frac{1-\cos(2\kappa k_{\textnormal{F}}|k|s)}{2\kappa k_{\textnormal{F}}|k|s}\right|.$ (6.15)
Note that it is well-known that for all $x>0$ it holds $\cos(x)<1-4x^{2}/\pi^{2}$ and thus $1-\cos(x)\geq 4x^{2}/\pi^{2}$ for all $x\in(0,\pi/2)$ Also it holds for all $x>\pi/2$ that
$\int_{\frac{\pi}{2}}^{x}\frac{1-\cos y}{y}\textnormal{d}y=\ln(x)-\ln(\frac{\pi}{2})+\text{Ci}(\frac{\pi}{2})-\text{Ci}(x)\geq\ln(x)-\ln(\frac{\pi}{2})$
since $\text{Ci}(\pi/2)>0$ and $\text{Ci}(\pi/2)\geq\text{Ci}(x)$ for all $x\geq\pi/2$. Thus we obtain
$\int_{0}^{x}\frac{1-\cos y}{y}\textnormal{d}y$ $\geq\chi(x>\frac{\pi}{2})\left\{\int_{0}^{\frac{\pi}{2}}\frac{4y}{\pi^{2}}\textnormal{d}y+\int_{\frac{\pi}{2}}^{x}\frac{1-\cos y}{y}\textnormal{d}y\right\}+\chi(x\leq\frac{\pi}{2})\int_{0}^{x}\frac{4y}{\pi^{2}}\textnormal{d}y$ (6.16)
$=\chi(x>\frac{\pi}{2})\left\{\frac{1}{2}+\ln(x)-\ln(\frac{\pi}{2})\right\}+\chi(x\leq\frac{\pi}{2})\frac{2x^{2}}{\pi^{2}}$ (6.17)
which is a differentiable lower bound.
Therefore we find
$2\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\langle h_{y},\eta_{s}\rangle\right|$
$\geq\frac{2\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\sum_{k\in\Gamma^{\textnormal{nor}}}\frac{\hat{V}(k)^{2}}{2\kappa k_{\textnormal{F}}|k|}\bigg{(}\chi(2\kappa k_{\textnormal{F}}|k|t>\frac{\pi}{2})\left\{\frac{1}{2}+\ln(2\kappa k_{\textnormal{F}}|k|t)-\ln(\frac{\pi}{2})\right\}$
$+\chi(2\kappa k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{2(2\kappa k_{\textnormal{F}}|k|t)^{2}}{\pi^{2}}\bigg{)}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$\eqqcolon b_{t}+td$ (6.18)
with $d\in\mathcal{O}(MN^{-\frac{2}{3}+\delta})$ defined in (6.10) and
$\dot{b}_{t}$ $=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\left(\chi(2\kappa k_{\textnormal{F}}|k|t>\frac{\pi}{2})\frac{1}{2\kappa k_{\textnormal{F}}|k|t}+\chi(2\kappa k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{8\kappa k_{\textnormal{F}}|k|t}{\pi^{2}}\right)-d>-d.$ (6.19)
In total, we can re-write the inequality (6.9) for $g_{t}\coloneqq\|\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})-1\big{)}\psi\|$ as the following integral inequality
$g_{t}\geq b_{t}-\alpha\int_{0}^{t}\textnormal{d}sg_{s}$ (6.20)
with $b_{t}$ and $\alpha$ defined in (6.18) and (6.12), respectively.
Claim.
We can bound $g_{t}$ from below for all $t\geq 0$ by a differentiable map $h_{t}$ which obeys the initial value problem
$h_{t}$ $=b_{t}-\alpha\int_{0}^{t}\textnormal{d}sh_{s}\qquad\text{with }h_{0}=0.$ (6.21)
Proof of claim.
Assume for the proof by contradiction $\exists t_{0}>0:h_{t_{0}}>g_{t_{0}}$ and set $I\subseteq\mathbb{R}_{>0}$ as largest open interval satisfying $t_{0}\in I$ and $h_{t}>g_{t}$ for all $t\in I$. Note that since $h_{t}$ is differentiable such an open interval exists. It holds $t_{1}\coloneqq\inf I>0$ since $h_{0}=0=g_{0}$ and for all $t\in I$
$\dot{g}_{t}\geq\dot{b}_{t}-\alpha g_{t}>\dot{b}_{t}-\alpha h_{t}=\dot{h}_{t}$
and therefore
$\int_{t_{1}}^{t_{0}}\dot{g}_{s}\textnormal{d}s>\int_{t_{1}}^{t_{0}}\dot{h}_{s}\textnormal{d}s\implies h_{t_{1}}-g_{t_{1}}>h_{t_{0}}-g_{t_{0}}>0.$
Thus we obtain the desired contradiction to $t_{1}$ being the inf of the largest set satisfying $h_{t}>g_{t}$. ∎
The solution of the initial value problem (6.21) is uniquely given by
$h_{t}=e^{-\alpha t}\int_{0}^{t}\dot{b}_{s}e^{\alpha s}\textnormal{d}s.$ (6.22)
Consequently it holds by inserting (6.19)
$h_{t}$ $=e^{-\alpha t}\int_{0}^{t}\dot{b}_{s}e^{\alpha s}\textnormal{d}s$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}e^{-\alpha t}\int_{0}^{t}\bigg{(}\chi(2\kappa k_{\textnormal{F}}|k|s>\frac{\pi}{2})e^{\alpha s}\frac{1}{2\kappa k_{\textnormal{F}}|k|s}$
$+\chi(2\kappa k_{\textnormal{F}}|k|s\leq\frac{\pi}{2})e^{\alpha s}\frac{8\kappa k_{\textnormal{F}}|k|s}{\pi^{2}}\bigg{)}\textnormal{d}s+d(e^{-\alpha t}-1)$ (6.23)
$\geq\frac{\pi\lambda^{2}k_{\textnormal{F}}}{\kappa}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\bigg{\{}\chi(|k|t\leq\frac{\pi}{4\kappa k_{\textnormal{F}}})\frac{8\kappa k_{\textnormal{F}}|k|}{\alpha^{2}\pi^{2}}\left(e^{-\alpha t}+\alpha t-1\right)+$
$+\chi(|k|t>\frac{\pi}{4\kappa k_{\textnormal{F}}})\frac{8\kappa k_{\textnormal{F}}|k|}{\alpha^{2}\pi^{2}}\big{(}e^{-\alpha\pi/(4\kappa k_{\textnormal{F}}|k|)}+\frac{\alpha\pi}{4\kappa k_{\textnormal{F}}|k|}-1\big{)}\bigg{\}}+d(e^{-\alpha t}-1)$ (6.24)
$\geq\frac{\lambda^{2}k_{\textnormal{F}}^{2}}{\alpha^{2}\pi}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}|k|\min\bigg{\{}f(t),f\big{(}\frac{\pi}{4\kappa k_{\textnormal{F}}|k|}\big{)}\bigg{\}}+d(e^{-\alpha t}-1)$ (6.25)
with $f(t)\coloneqq\left(e^{-\alpha t}+\alpha t-1\right)$ defining a non-negative monotonically increasing function. Since $d\in\mathcal{O}(MN^{-\frac{2}{3}+\delta})$ we obtain the desired result that (5.20) has a lower bound of order 1. ∎
Appendix A Estimates within the bosonization framework
In this section we collect all relevant estimates of the bosonization framework which was first developed in [BNP+19, BNP+21a, BNP+21b, BPSS23] in the semiclassical regime. In addition to the references we provide brief proof sketches in order to stress the dependence of the coupling constants.
Lemma A.1 (Approximation of $n_{\alpha}(k)$, [BNP+21b, Lemma 5.1]).
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma^{\textnormal{nor}},\alpha\in\mathcal{I}_{k}$ it holds
$n_{\alpha}(k)^{2}=\frac{4\pi k_{\textnormal{F}}^{2}}{M}|k\cdot\hat{\omega}_{\alpha}|\big{(}1+o(1)\big{)}.$
Note that $|k\cdot\hat{\omega}_{\alpha}|>N^{-\delta}$ by construction of $\mathcal{I}_{k}$.
Lemma A.2 (Approximation of $\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}$).
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma^{\textnormal{nor}}$ it holds
$\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}=k_{\textnormal{F}}^{2}|k|\pi\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$
Proof.
It holds from [BNP+19, Proposition 3.1]
$n_{\alpha}(k)^{2}=k_{\textnormal{F}}^{2}|k|\sigma(p_{\alpha})u_{\alpha}(k)^{2}\left(1+\mathcal{O}(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta})\right)$
with $\cos\theta_{\alpha}\coloneqq|\hat{k}\cdot\hat{\omega}_{\alpha}|\equiv u_{\alpha}(k)^{2}$. Choose $\varphi_{\alpha}$ for the azimuth angle of $\omega_{\alpha}$. We estimate the $\alpha$-sum by an appropriate surface integral over the patch $p_{\alpha}$
$\left|\int_{p_{\alpha}}\textnormal{d}\sigma\ \cos\theta-\sigma(p_{\alpha})\cos\theta_{\alpha}\right|$ $\leq\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\textnormal{d}}{\textnormal{d}\theta}\cos\theta\right|\sup_{(\theta,\varphi)\in p_{\alpha}}\left|\theta-\theta_{\alpha}\right|\sigma(p_{\alpha})$
$\leq CM^{-\frac{3}{2}}$
where we used $\left|\theta-\theta_{\alpha}\right|\leq CM^{-\frac{1}{2}}$ and $\sigma(p_{\alpha})\leq CM^{-1}$ by the patch construction. Note that the integral over $\tilde{S}\coloneqq\bigcup_{\alpha\in\mathcal{I}_{k}}p_{\alpha}$ which excludes a collar of width $N^{-\delta}$ can be approximated by an integral over the half-sphere $S$
$\left|\int_{\tilde{S}}\textnormal{d}\sigma\ \cos\theta-\int_{S}\textnormal{d}\sigma\ \cos\theta\right|<C\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)$
which can be calculated explicitly
$\int_{S}\textnormal{d}\sigma\ \cos\theta=2\pi\int_{0}^{\pi/2}\textnormal{d}\theta\ \cos\theta\sin\theta=\pi.$
Thus in total we obtain with the triangle inequality and $|k|<C$
$\left|\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}-k_{\textnormal{F}}^{2}|k|\pi\right|$ $=\left|\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}-k_{\textnormal{F}}^{2}|k|\int_{S}\textnormal{d}\sigma\ \cos\theta\right|$
$\leq Ck_{\textnormal{F}}^{2}|k|\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+M^{-\frac{1}{2}}+N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)$
$\leq Ck_{\textnormal{F}}^{2}|k|\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right).$
Lemma A.3 (Estimates on the CAR error term, [BNP+21b, Lemma 5.2]).
For $k^{\prime},k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$ the error term $\mathcal{E}_{\alpha}(k,k^{\prime})$ as defined in (2.22) satisfies $\mathcal{E}_{\alpha}(k,k^{\prime})=\mathcal{E}_{\alpha}(k^{\prime},k)^{\ast}$, commutes with $\mathcal{N}$ and for all $\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}$ it holds for all $\zeta\in\mathcal{F}$
$|\mathcal{E}_{\gamma}(k,k^{\prime})|^{2}\leq\sum_{\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}}|\mathcal{E}_{\gamma}(k,k^{\prime})|^{2}$ $\leq C\left(MN^{-\frac{2}{3}+\delta}\mathcal{N}\right)^{2},$
$\sum_{\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}}\|\mathcal{E}_{\gamma}(k,k^{\prime})\zeta\|$ $\leq CM^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}\|\mathcal{N}\zeta\|.$
Lemma A.4 (Pair operator bounds, [BNP+21b, Lemma 5.3]).
It holds for all $k\in\Gamma^{\textnormal{nor}}$ and $\psi\in\mathcal{F},f\in l^{2}(\mathcal{I}_{k})$:
  1. (i).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|^{2}\leq\|\mathcal{N}^{\frac{1}{2}}\psi\|^{2}$
  2. (ii).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|\leq M^{\frac{1}{2}}\|\mathcal{N}^{\frac{1}{2}}\psi\|$
  3. (iii).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|\leq M^{\frac{1}{2}}\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|$
  4. (iv).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|^{2}\leq\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|^{2}$
  5. (v).
    $\|\sum_{\alpha\in\mathcal{I}_{k}}f_{\alpha}c_{\alpha}^{\ast}(k)\psi\|\leq\|f\|_{l^{2}}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|$
  6. (vi).
    $\sum_{\alpha\in\mathcal{I}_{k}}c_{\alpha}^{\ast}(k)c_{\alpha}(k)\leq\mathcal{N}$
Lemma A.5 (Error of linearized of kinetic energy, [BNP+21b, Lemma 8.2]).
It holds for $k\in\Gamma^{\textnormal{nor}}$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{H}_{0},c_{\alpha}^{\ast}(k)]=2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
with
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|^{2}$ $\leq C\left(\kappa N^{\frac{1}{3}}M^{-\frac{1}{2}}\right)^{2}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|^{2}$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|$ $\leq C\kappa N^{\frac{1}{3}}\|\mathcal{N}^{\frac{1}{2}}\psi\|.$
Proof.
Observe that
$[\mathbb{H}_{0},c_{\alpha}^{\ast}(k)]$ $=\frac{1}{n_{\alpha}(k)}\sum_{p\in B_{\textnormal{F}}^{c}\cap B_{\alpha},p-k\in B_{\textnormal{F}}\cap B_{\alpha}}\sum_{l\in\mathbb{Z}^{3}}\Big{[}e(l)a_{l}^{\ast}a_{l},a_{p}^{\ast}a_{p-k}^{\ast}\Big{]}$
$=\frac{1}{n_{\alpha}(k)}\sum_{p\in B_{\textnormal{F}}^{c}\cap B_{\alpha},p-k\in B_{\textnormal{F}}\cap B_{\alpha}}\left(e(p)+e(p-k)\right)a_{p}^{\ast}a_{p-k}^{\ast}$
$=2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
One makes the identification $\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\equiv c_{\alpha}^{g}(k)$ representing a weighted operator with
$g(p,k)$ $=e(p)+e(p-k)-2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|=\kappa|p|^{2}-\kappa|p-k|^{2}-2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|$
$=\kappa\left(2k\cdot(p-k_{\textnormal{F}}\hat{\omega}_{\alpha})-|k|^{2}\right)$
where we used $e(p)$ as defined in (2.8). The bound follows from [BNP+21b, Lemma 5.4] which depends on $\|g\|_{l^{\infty}}$. This can be estimated by using $\text{diam}(B_{\alpha})\leq CN^{\frac{1}{3}}M^{-\frac{1}{2}}$ such that $|g(p,k)|\leq C\kappa N^{\frac{1}{3}}M^{-\frac{1}{2}}$. ∎
Lemma A.6 (Error of bosonized kinetic energy, [BNP+21b, eq. (8.6)]).
It holds for $k\in\Gamma^{\textnormal{nor}}$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{D}_{\textnormal{B}},c_{\alpha}^{\ast}(k)]=2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)^{\ast}$
with
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$ $\leq C\left(\kappa k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2},$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|$ $\leq C\kappa k_{\textnormal{F}}M^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|,$
Proof.
It holds $\mathfrak{E}_{\alpha}^{\text{B}}(k)=2\kappa k_{\textnormal{F}}\sum_{l}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\chi(\alpha\in\mathcal{I}_{l})$ and therefore
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$
$\leq\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}||2\kappa k_{\textnormal{F}}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}\leq C\kappa k_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}||\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq C\kappa k_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}CMN^{-\frac{2}{3}+\delta}||\mathcal{N}c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq C\kappa k_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{l^{\prime}\in\Gamma^{\textnormal{nor}}}\|c_{\alpha}(l)(\mathcal{N}-2)\psi\|\ \|c_{\alpha}(l^{\prime})(\mathcal{N}-2)\psi\|$
$\leq C\kappa k_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2}$
where we used in the second line $l\in\Gamma^{\textnormal{nor}}$ bounded, Lemma A.3 in the third line, $\mathcal{N}c_{\alpha}(l)=c_{\alpha}(l)(\mathcal{N}-2)$ in the fourth line and Cauchy-Schwarz and Lemma A.4 in the last line.
The second statement simply follows from Cauchy-Schwarz. ∎
Lemma A.7 (Approximation of patch decomposed operators).
It holds for all $k\in\Gamma^{\textnormal{nor}}$
$\|\big{(}b(k)-\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)c_{\alpha}(k)+\text{h.c.}\big{)}\psi\|\leq C(N^{\frac{1}{3}-\frac{\delta}{2}}+N^{\frac{1}{6}}M^{\frac{1}{4}})\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|.$
Proof.
We show that the inequality derived in [BNP+21b, eq. (4.10)] holds here independent of our adjustments. The proof in [BNP+21b, eq. (4.10)] makes use of a decomposition of the uncovered Fermi surface
$U\coloneqq\big{(}B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)\big{)}\ \backslash\ \bigcup_{\alpha\in\mathcal{I}_{k}}\big{(}B_{\alpha}\cap(B_{\alpha}+k)\big{)}$
into $Y=\{p\in U:e(p)+e(p-k)\leq 4\kappa N^{\frac{1}{3}-\delta}\}$. Although the decomposition depends on the kinetic energy the final bound is independent of the kinetic energy coupling $\kappa$. The condition $e(p)+e(p-k)\leq 4\kappa N^{\frac{1}{3}-\delta}$ is chosen to imply $\hat{p}\cdot\hat{k}\leq CN^{-\delta}$ due to $|p|\sim N^{1/3}$ and $|k|\sim 1$. Thus, one derives $|U\backslash Y|\in\mathcal{O}(N^{1/3}M^{1/2})$ which gives the bound on $U\backslash Y$ with Cauchy-Schwarz. For the bound on $Y$ insert the bound (see [BPSS23, eq. (4.10)])
$\sum_{p\in B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)}\frac{1}{e(p)+e(p-k)}\leq C\kappa^{-1}N^{\frac{1}{3}}$
into
$\sum_{p\in B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)}1=\sum_{p\in B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)}\frac{4\kappa N^{\frac{1}{3}-\delta}}{e(p)+e(p-k)}$
canceling the $\kappa$-dependence. In total we end up with the same inequalities. ∎
Lemma A.8 (Estimate of non-bosonizable terms, [BNP+21a, eq. (4.6)]).
It holds for $\mathcal{E}$ as defined in (2.9b) the following estimate for all $\psi\in\mathcal{F}$
$\|\mathcal{E}\psi\|\leq C\lambda\|\hat{V}\|_{1}\|\mathcal{N}\psi\|.$
Acknowledgment
We are grateful for helpful discussions with David Mitrouskas, Benjamin Schlein and Karla Schön. This work has been partially supported by SFB/Transregio TRR 352 – Project-ID 470903074.