### v2

# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

Abstract
We study the quantum dynamics of a homogeneous ideal Fermi gas coupled to an impurity particle on a three-dimensional box with periodic boundary condition. For large Fermi momentum $k_{\textnormal{F}}$, we prove that the effective dynamics is generated by a Fröhlich-type polaron Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. Moreover, we prove that the effective dynamics can be approximated by an explicit coupled coherent state. Our method is applicable to a range of interaction couplings, in particular including interaction couplings of order 1 and time scales of the order $k_{\textnormal{F}}^{-1}$.
1 Introduction
The study of impurities in quantum gases has garnered considerable attention due to its relevance in various physical contexts, ranging from solid-state physics to cold atom experiments. In this context quasi-particles such as polarons stand as intriguing entities emerging from the interaction of a single impurity particle with a surrounding medium. The concept of a polaron, originally introduced by Lev Landau to study the motion of an electron in a dielectric crystal [Lan65], most famously emerges from the celebrated Fröhlich Hamiltonian in second quantization formalism describing electron-phonon interactions [Frö54]. Subsequently, the polaron concept was extended to all kind of surrounding media including Bose and Fermi gases. The formation conditions and properties of polarons are believed to play a central role to understand the transport properties and the effective mass of impurities within the host material.
In this article, we study with mathematical rigor the dynamics of an impurity particle immersed in a dense gas of fermions as surrounding medium. Interactions between fermions are neglected and we assume that the impurity particle interacts with the Fermi gas via a finite-range potential in momentum space. The initial state $\psi$ of the system is a product state between the impurity state and a filled Fermi ball. This mathematical framework finds resonance with recent experimental and theoretical advancements in the study of ultracold atoms [SWSZ09, KSN+12, CJL+15]. We emphasize that our assumption on the interaction potential aligns with the growing interest among experimentalists and theorists in cases where the impurity is charged, and hence the interaction potential is long-ranged in position space [CCGB22, MJ24]. We show that the effective dynamics of the system is governed by a Fröhlich-type Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. More specifically, the excitations relative to the filled Fermi ball are up to a constant described by the Hamiltonian
$\mathbb{H}^{\text{F}}=(-\Delta_{y})\otimes 1+1\otimes\mathbb{D}_{\text{B}}+\Phi(h_{y})$ (1.1)
with $(-\Delta_{y})$ describing the kinetic energy of the impurity particle, $\mathbb{D}_{\text{B}}$ describing the kinetic energy of the excitation field and $\Phi(h_{y})\coloneqq c^{\ast}(h_{y})+c(h_{y})$ the linear coupling between impurity particle and excitations. The operators $c^{\ast}$ and $c$ describing this excitation field coincide with those introduced in a series of pioneering studies on the correlation energy of interacting fermions [BNP+19, BNP+21a, BNP+21b]. We note that an effective Hamiltonian of a similar type to (1.1) has recently been derived in another microscopic setting involving a tracer particle interacting with excitations of a Bose Einstein condensate [LP22, MS20].
Subsequently, we show that the effective time evolved state can be up to a phase factor approximated by a time-dependent coupled coherent state $W(\eta_{t})\phi\otimes\Omega$ where $W$ is the Weyl operator of the excitation field which is simply parameterized by a function $\eta_{t}$ and $\Omega$ is the Fock space vacuum. An explicit expression for $\eta_{t}$ is derived which allows for determining the number of collective excitations over time. We believe that such quantities are particularly helpful to gain deeper insights into the formation process of quasi-particles as studied in experiments such as [CJL+16]. Eventually, we show that the linear coupling term $\Phi(h_{y})$ cannot be omitted in the effective description but adds a leading order effect to the effective dynamics in the setting of interaction couplings of order 1.
Our results hold for a variety of time scales and couplings, describing different interaction strengths and mass ratios, which will be specified in the subsequent section. We note that the same microscopic model has been studied in [JMPP17, JMP18, MP21] but with very specific choices of couplings different from ours, leading to an effective decoupling of impurity and gas.
1.1 The microscopic model
We consider an impurity particle interacting with $N$ spinless fermions on a 3-dimensional box with periodic boundaries described by $\Lambda\coloneqq\mathbb{T}^{3}\coloneqq\mathbb{R}^{3}/(2\pi\mathbb{Z}^{3})$. The system is described by a state in the Hilbert space $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\mathcal{H}_{N}^{-}=L^{2}(\Lambda)^{\wedge N}$ where $y$ is the coordinate of the impurity particle and $\{x_{i}\}_{i=1,\ldots,N}$ are the coordinates of the fermions. The Hamiltonian for our main model of the system is given by
$H_{N}=-\beta\Delta_{y}+\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda\sum_{i=1}^{N}V(x_{i}-y)$ (1.2)
and parameterized by $\beta,\lambda>0$. Note that the different parts of the Hamiltonian on different tensor components of our Hilbert space writing, i.e. we used the short-hand notation writing, for example, $-\Delta_{y}\coloneqq-\Delta_{y}\otimes 1$ for the Laplacian acting on the impurity particle. The interaction $V$ is assumed to have a Fourier transform $\hat{V}$ with compact support satisfying $\hat{V}(k)=\hat{V}(-k)$ and $\hat{V}(k)\geq 0$ for all $k\in\mathbb{Z}^{3}$. It is well-known that under this assumption the Hamiltonian (1.2) defines a self-adjoint operator which generates by Stone’s theorem the unitary time evolution $e^{-H_{N}t}$.
We are interested in the dynamics of the system governed by the time-dependent Schrödinger equation of the form
$i\hbar\frac{\textnormal{d}}{\textnormal{d}t}\psi_{t}=H_{N}\psi_{t},\qquad\psi_{0}\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}.$ (1.3)
where $\hbar$ is the reduced Planck constant. For mathematical convenience we will always set $\hbar=1$ unless explicitly stated otherwise.
Note that the filled Fermi ball is a ground state for the non-interacting Fermion system. It is non-degenerate and explicitly given by
$\Omega_{0}\coloneqq\bigwedge_{k\in B_{\textnormal{F}}}f_{k},\qquad f_{k}(x)\coloneqq\frac{\exp(ikx)}{(2\pi)^{3/2}}\in L^{2}(\Lambda).$ (1.4)
We choose the initial state to be of product form
$\psi_{0}\big{(}y;x_{1},\ldots x_{N}\big{)}\coloneqq\phi(y)\otimes\Omega_{0}(x_{1},\ldots x_{N}),$ (1.5)
with a general state $\phi$ for the impurity particle i.e., the system is initially prepared in a state describing a ground state of the ideal Fermi gas which does not interact with the impurity particle.
Furthermore, we choose the Fermi momentum $k_{F}$ to be our parameter of the system in the sense that the particle number is defined as
$N\equiv N(k_{\textnormal{F}})\coloneqq|B_{\textnormal{F}}|,\qquad B_{\textnormal{F}}\coloneqq\{k\in\mathbb{Z}^{3}:|k|\leq k_{\textnormal{F}}\},$ (1.6)
i.e. the particle number $N$ of the Fermi gas is chosen such that the Fermi ball is completely filled. Note that the average density is in this case proportional to the number $N$ of gas particles due to the following relation
$k_{\textnormal{F}}=\left(\frac{3}{4\pi}\right)^{1/3}N^{\frac{1}{3}}+\mathcal{O}(1)$ (1.7)
which is a consequence of Gauss’ counting argument.
1.2 Relevant parameters and time scales
In the following, we present the ranges of $\beta$ and $\lambda$ we are aiming for, and discuss the physical meaning of the parameters. In addition, it is important to discuss on which time scales our results hold.
  • The parameter $\beta$ determines the mass ratio of the impurity and Fermi gas particles such that $\beta\ll 1$ corresponds to a relatively heavy, $\beta\gg 1$ to a relatively light impurity particle. The case of equal masses corresponds to $\beta=1$. Our work requires the restriction $\beta\in o(k_{\textnormal{F}})$ see Remark 3.7. Since $k_{\textnormal{F}}$ tends to infinity, this allows to include all three cases.
  • The parameter $\lambda\in[k_{\textnormal{F}}^{-1/6},1]$ models the coupling strength between the Fermi gas and the impurity. For small $\lambda$ we expect a decoupling between the gas and the impurity in the sense that the time evolution given by (1.3) does not entangle an initial state of product form $\phi\otimes\Omega_{0}$. Such a result was shown in [MP21] with $\beta=1$ and $\lambda=k_{\textnormal{F}}^{-1/2}$ in three dimensions and [JMPP17, JMP18] with $\beta=\lambda=1$ in two dimensions. Note that our interaction coupling $\lambda$ is much larger than in [MP21] and in particular, we are able to include $\lambda=1$. As will be discussed in Remark 3.10, it is necessary to include excitations of the Fermi gas into the effective description for $\lambda=1$.
  • Our approximations will apply for times $t\in\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$, so that a weaker coupling strength $\lambda$ corresponds to slightly larger times. The times $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$ are on the time scale of the fermions near the Fermi surface, which have an approximate momentum of $k_{\textnormal{F}}$ and thus travel a distance of order 1 in this time. We are therefore able to enter a time scale where the impurity particle can resolve the motion of the fermions. We remark that the results in [MP21] can be transferred to this setting of $\lambda=1$, however, allowing only for shorter time scales of $t\in o(k_{\textnormal{F}}^{-1})$.
We remark that results obtained in the previously mentioned parameter range remain valid under re-scaling by multiplying the Hamiltonian by an overall factor and absorbing it by re-scaling the time variable. Although we will stick to the above choices, it might still be helpful for the reader to see some connections to other scaling regimes by re-scaling. We present two of them briefly:
  • Time scales of order 1 are obtained by multiplying the Hamiltonian by an overall factor of $k_{\textnormal{F}}^{-1}$. In this case, the factor $k_{\textnormal{F}}^{-1}$ appears as new parameters in front of the kinetic energy of the Fermi gas corresponding to a heavy fermion regime.
  • In recent years the so-called semiclassical regime has been widely studied in the analysis of dense Fermi gases, as can be seen for example in [Ben22, Saf23]. This regime is associated with identifying $\hbar\coloneqq k_{\textnormal{F}}^{-1}$ as small parameter instead of setting $\hbar=1$. It is achieved by multiplying the Hamiltonian by an overall factor of $k_{\textnormal{F}}^{-2}$ and absorbing $k_{\textnormal{F}}^{-1}$ in the time variable. For this new time scale, our theorem makes a statetment for times of order 1. The re-scaled Schrödinger equation takes the form of
    $i\hbar\partial_{t}\psi_{t}=\left(\hbar^{2}\beta(-\Delta_{y})+\hbar^{2}\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda k_{\textnormal{F}}^{-2}\sum_{i=1}^{N}V(x_{i}-y)\right)\psi_{t}$ (1.8)
    One identifies $\lambda k_{\textnormal{F}}^{-2}\in[k_{\textnormal{F}}^{-13/6},k_{\textnormal{F}}^{-2}]$ as re-scaled interaction coupling.
2 Preliminaries
Second quantization
It is convenient to consider $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ as the $N$-particle sector of $L^{2}(\Lambda)\otimes\mathcal{F}$ with the fermionic Fock space $\mathcal{F}$ constructed over $L^{2}(\Lambda)$. This way, we have access to the powerful formalism of second quantization with the fermionic creation operator $a_{p}^{\ast}$ creating a particle with momentum $p\in\mathbb{Z}^{3}$ and the annihilation operator $a_{p}$ annihilating a particle with momentum $p\in\mathbb{Z}^{3}$. Those operators satisfy the canonical anticommutation relations (CAR)
$\forall p,q\in\mathbb{Z}^{3}:\{a_{p},a_{q}^{\ast}\}=\delta_{p,q},\qquad\{a_{p},a_{q}\}=0=\{a_{p}^{\ast},a_{q}^{\ast}\}.$ (2.1)
Furthermore we introduce the fermionic number operator $\mathcal{N}\coloneqq\sum_{p\in\mathbb{Z}^{3}}a_{p}^{\ast}a_{p}$ and the vacuum $\Omega$ satisfying $a_{p}\Omega=0$ for all $p\in\mathbb{Z}^{3}$.
We lift our $N$-particle Hamiltonian $H_{N}$ to Fock space as
$\mathbb{H}=\underbrace{-\beta\Delta_{y}}_{\eqqcolon h_{0}}+\underbrace{\sum_{k\in\mathbb{Z}^{3}}|k|^{2}a_{k}^{\ast}a_{k}}_{\eqqcolon\mathbb{H}^{\textnormal{kin}}}+\underbrace{\lambda\sum_{k,p\in\mathbb{Z}^{3}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}}_{\eqqcolon\mathbb{V}}$ (2.2)
[…]
We will mostly use the abuse of notation $\mathbb{A}\equiv 1\otimes\mathbb{A}$ as operator on $L^{2}(\Lambda)\otimes\mathcal{F}$ where $\mathbb{A}$ acts as an operator on the Fock space part.
Particle-hole transformation
In our analysis, the primary objective is to focus on excitations relative to the non-interacting Fermi ball. In particular, we want to use a description of our fermionic system in which the non-interacting Fermi ball $\Omega_{0}=\prod_{k\in B_{\textnormal{F}}}a^{\ast}_{k}\Omega$ is mapped to the vacuum. To achieve this, we employ the particle-hole transformation, which is a specific type of fermionic Bogoliubov transformation as creation operators are mapped to linear combinations of creation and annihilation operators while preserving the CAR. The particle-hole transformation is defined as the map $R:\mathcal{F}\to\mathcal{F}$ satisfying
$R^{\ast}a_{k}^{\ast}R\coloneqq\begin{cases}a_{k}^{\ast}&\text{if }k\in B_{\textnormal{F}}^{c},\\ a_{k}&\text{if }k\in B_{\textnormal{F}}\end{cases}\quad\text{and}\quad R\Omega\coloneqq\Omega_{0}\ .$ (2.3)
It is easy to check that the map is well-defined, unitary and satisfies $R^{-1}=R^{\ast}=R$.
With this, we can re-write the initial state (1.5) representing a non-interacting impurity particle and a Fermi gas as
$\psi_{0}=\phi\otimes\Omega_{0}=(1\otimes R)\left(\phi\otimes\Omega\right)\eqqcolon R\psi.$ (2.4)
Later on, we will mostly use the product state $\psi=\phi\otimes\Omega$ of the impurity and the vacuum instead of $\psi_{0}$.
Furthermore, we define
$E_{N}^{\text{pw}}\coloneqq\sum_{k\in B_{\textnormal{F}}}|k|^{2}=\langle R\Omega,H_{N},R\Omega\rangle$ (2.5)
to be the energy of the non-interacting Fermi ball.
Of greatest interest is of course the action of the particle-hole transformation on the microscopic Hamiltonian as generator of the dynamics. The conjugation with $R$ of $\mathbb{H}=-\beta\Delta_{y}+\mathbb{H}^{\textnormal{kin}}+\mathbb{V}$ yields
$\mathbb{H}_{0}$ $\coloneqq R^{\ast}\mathbb{H}^{\textnormal{kin}}R-E_{N}^{\text{pw}}=\sum_{k\in\mathbb{Z}^{3}}|k|^{2}R^{\ast}a_{k}^{\ast}RR^{\ast}a_{k}R-E_{N}^{\text{pw}}$ (2.6)
$=\sum_{k\in B_{\textnormal{F}}}|k|^{2}a_{k}a_{k}^{\ast}+\sum_{k\in B_{\textnormal{F}}^{c}}|k|^{2}a_{k}^{\ast}a_{k}-E_{N}^{\text{pw}}$ (2.7)
$=\sum_{k\in\mathbb{Z}^{3}}e(k)a_{k}^{\ast}a_{k}\qquad\text{with }e(k)\coloneqq\begin{cases}|k|^{2}&\text{if }k\in B_{\textnormal{F}}^{c},\\ -|k|^{2}&\text{if }k\in B_{\textnormal{F}}.\end{cases}$ (2.8)
Similarly, we can see that
$R^{\ast}\mathbb{V}R$ $=\lambda\sum_{k,p\in\mathbb{Z}^{3}}\hat{V}(k)e^{iky}R^{\ast}a_{p}^{\ast}RR^{\ast}a_{p-k}R$
$=\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p-k\in B_{\textnormal{F}}^{c},\\ p\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}a_{p-k}+\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}^{\ast}$ (2.9a)
$\quad+\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}},\\ p-k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}a_{p-k}^{\ast}+\lambda\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}^{c}\end{subarray}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}.$ (2.9b)
For later purposes we shall introduce for $\varphi\in l^{2}(\mathbb{Z}^{3})$ the short-notation
$b(\varphi)$ $\coloneqq\sum_{k\in\mathbb{Z}^{3}}\overline{\varphi(k)}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}a_{p-k}a_{p},$ (2.10)
$b^{\ast}(\varphi)$ $=\sum_{k\in\mathbb{Z}^{3}}\varphi(k)\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}a_{p}^{\ast}a_{p-k}^{\ast}.$ (2.11)
We can then write
$R^{\ast}\mathbb{H}R=-\beta\Delta_{y}+\mathbb{H}_{0}+b^{\ast}(\tilde{h}_{y})+b(\tilde{h}_{y})+\mathcal{E}$ (2.12)
with $\tilde{h}_{y}(k)\coloneqq\lambda\hat{V}(k)e^{iky}$ and $\mathcal{E}$ is given by the terms of (2.9b) since
$\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p-k\in B_{\textnormal{F}}^{c},\\ p\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{p}a_{p-k}=$ $\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}\tilde{p}\in B_{\textnormal{F}}^{c},\\ \tilde{p}+k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{iky}a_{\tilde{p}+k}a_{\tilde{p}}$
$=$ $\sum_{k\in\mathbb{Z}^{3}}\sum_{\begin{subarray}{c}p\in B_{\textnormal{F}}^{c},\\ p-k\in B_{\textnormal{F}}\end{subarray}}\hat{V}(k)e^{-iky}a_{p-k}a_{p}$ (2.13)
where we used that $\hat{V}(-k)=\hat{V}(k)$.
Almost-bosonic operators and patch decomposition
Our effective description of the microscopic system described by (1.2) will involve the emergence of almost-bosonic particles describing pair excitations of the Fermi ball. Those pair excitations will be delocalized over the Fermi surface in the sense that they correspond to a linear combination of pairs of fermionic operators. As mentioned before the almost-bosonic pair operators which occur in this article coincide with the ones introduced in the series of seminal works [BNP+19, BNP+21a, BNP+21b, BPSS23] on the correlation energy of a weakly interacting Fermi gas. We give a brief introduction to the construction of those operators with the most relevant properties in this subsection and in Section A.
A key ingredient for the approximation of the microscopic fermionic system by almost-bosonic excitations is the decomposition of the Fermi surface into patches. This will allow to approximate the fermionic kinetic energy term by a term quadratic in the almost-bosonic pair operators.
Introduce the bisecting subset of $\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}$
$\Gamma\coloneqq\left\{(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}:(k_{3}>0\vee k_{3}=0),(k_{2}>0\vee k_{2}=k_{3}=0),k_{1}>0\right\}$ (2.14)
allowing the decomposition $\Gamma\cup(-\Gamma)=\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}$.
The construction works as follows:
  1. (i).
    Choose the number $M$ of patches satisfying
    $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta},\qquad\delta\in(0,\frac{1}{6}).$ (2.15)
    The lower bound on $M$ is needed to control the number of momenta inside each patch whereas the upper bound is needed to suppress Pauli’s principle. The choice of $\delta$ and $c$ will be taken later.
  2. (ii).
    Define equal-area disjoint patches $p_{\alpha}$ as follows
    • $p_{1}$ is spherical cap of area $4\pi/M$,
    • decompose remaining semi-sphere into $\sqrt{M}/2$ collars,
    • leave corridors of width $2R\coloneqq 2\ \text{supp}\hat{V}$ between adjacent patches,
    • define patches of southern semi-sphere by reflection $k\mapsto-k$.
    Define $\omega_{\alpha}\in S^{2}$ as centers of the patch $p_{\alpha}$. A useful graphical sketch of this patch construction is given in Figure 1 of [BNP+19].
  3. (iii).
    For given $k\in\Gamma$ define the set of north and south patch indices
    $\mathcal{I}_{k}^{+}$$\coloneqq\{\alpha\in\{1,\ldots,M\}\mid k\cdot\hat{\omega}_{\alpha}\geq N^{-\delta}\},$ (2.16)
    $\mathcal{I}_{k}^{-}$$\coloneqq\{\alpha\in\{1,\ldots,M\}\mid k\cdot\hat{\omega}_{\alpha}\leq-N^{-\delta}\}$ (2.17)
    and $\mathcal{I}_{k}\coloneqq\mathcal{I}_{k}^{+}\cup\mathcal{I}_{k}^{-}$. This has the effect of excluding a strip around the equator of the Fermi ball, where the number of momenta per patch may become too small. Note that $\delta>0$ coincides with the parameter in step 1.
  4. (iv).
    Define the collective almost-bosonic creation operator and its normalization factor as
    $c_{\alpha}^{\ast}(k)\coloneqq\begin{cases}b_{\alpha}^{\ast}(+k)&\text{if }\alpha\in\mathcal{I}_{k}^{+},\\ b_{\alpha}^{\ast}(-k)&\text{if }\alpha\in\mathcal{I}_{k}^{-}.\end{cases},\ \ n_{\alpha}(k)\coloneqq\begin{cases}m_{\alpha}(k)&\text{if }\alpha\in\mathcal{I}_{k}^{+},\\ m_{\alpha}(-k)&\text{if }\alpha\in\mathcal{I}_{k}^{-}.\end{cases}$ (2.18)
    with
    $b_{\alpha}^{\ast}(k)\coloneqq\frac{1}{m_{\alpha}(k)}\sum_{\begin{subarray}{c}p\in B_{F}^{c}\cap B_{\alpha},\\ p-k\in B_{F}\cap B_{\alpha}\end{subarray}}a_{p}^{\ast}a_{p-k}^{\ast},\quad m_{\alpha}(k)^{2}\coloneqq\sum_{\begin{subarray}{c}p\in B_{F}^{c}\cap B_{\alpha},\\ p-k\in B_{F}\cap B_{\alpha}\end{subarray}}1$ (2.19)
    being sensitive to being on the north or south hemisphere. The creation operator can be seen as collective in the sense that it involves a superposition of all possible fermion pairs with relative momentum $k$.
Similarly to (2.11) introduced in the previous subsection, we define as in [BNP+21a]
$c^{\ast}(\eta)\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}c_{\alpha}^{\ast}(k)$ (2.20)
with inner product $\langle\eta,\varphi\rangle\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}\varphi_{\alpha}(k)$ for all $\eta,\varphi\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$.
The following statements hold as a consequence of the above construction.
  • The surface area of a patch satisfies $\sigma(p_{\alpha})\in\mathcal{O}(1/M)$.
  • The Canonical Commutation Relations (CCR) are satisfied up to an error term (see [BNP+19, Lemma 4.1]): It holds for all $k^{\prime},k\in\Gamma$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$
    $[c_{\alpha}(k),c_{\beta}(k^{\prime})]$ $=[c_{\alpha}^{\ast}(k),c_{\beta}^{\ast}(k^{\prime})]=0,$ (2.21)
    ${}[c_{\alpha}(k),c_{\beta}^{\ast}(k^{\prime})]$ $=\delta_{\alpha,\beta}\big{(}\delta_{k,k^{\prime}}+\mathcal{E}_{\alpha}(k,k^{\prime})\big{)}$ (2.22)
    satisfying $\mathcal{E}_{\alpha}(k,k)\leq 0$, $\mathcal{E}_{\alpha}(k,l)=\mathcal{E}_{\alpha}(l,k)^{\ast}$ and
    $\forall\psi\in\mathcal{F}:\qquad\|\mathcal{E}_{\alpha}(k,k^{\prime})\psi\|\leq\frac{2}{n_{\alpha}(k)n_{\alpha}(k^{\prime})}\|\mathcal{N}\psi\|.$ (2.23)
  • The almost-bosonic operators change the number operator by two (see [BNP+19, Lemma 2.3]) in the following sense
    $c_{\alpha}(k)\mathcal{N}$ $=(\mathcal{N}+2)c_{\alpha}(k).$ (2.24)
  • The normalization constant satisfies (see [BNP+19, Proposition 3.1])
    $n_{\alpha}(k)^{2}=\frac{4\pi k_{\textnormal{F}}^{2}}{M}|k\cdot\hat{\omega}_{\alpha}|\big{(}1+o(1)\big{)}.$ (2.25)
Also note that the summation in the definition of the almost-bosonic operators $c_{\alpha}^{\ast}(k)$ and $c_{\alpha}(k)$ involves only finite sets. Unlike in the exactly bosonic case our almost-bosonic operators therefore inherit boundedness from the fermionic constituents which satisfy $\|a_{k}^{\ast}\|=\|a_{k}\|=1$. Subtle questions about the mutual adjointness and domain of the almost-bosonic operators remain trivial in our case.
3 Main results
3.1 Effective time evolution
We are now focusing on the effective time evolution of the initial state $\psi_{0}=R\psi\equiv\phi\otimes R\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$, i.e. a uncorrelated product state with the non-interacting Fermi gas prepared as its non-degenerate ground state with no initial excitations. This set-up corresponds to a system where the impurity does not interact with the cold Fermi gas at time $t=0$. Over time, we expect that the influence of the impurity particle creates and annihilates excitations of the Fermi ball. Therefore we will use the particle-hole transformation as defined in (2.3) to connect the microscopic description to the following effective description: Let
$\mathbb{H}^{\text{eff}}\coloneqq$ $-\beta\Delta_{y}+\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)+E_{N}^{\text{pw}}$
$+\lambda\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\hat{V}(k)e^{iky}n_{\alpha}(k)\bigg{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(-k)\bigg{)}$ (3.1)
be our effective Hamiltonian with $\epsilon_{\alpha}(k)=2k_{F}|k\cdot\omega_{\alpha}|$ and $E_{N}^{\text{pw}}$ as defined in (2.5). We introduce for all $k\in\Gamma$ and $\alpha\in\mathcal{I}_{k}$
$(h_{y})_{\alpha}(k)\coloneqq\lambda\hat{V}(k)e^{iky}n_{\alpha}(k).$ (3.2)
Note that the effective Hamiltonian acts on the components of the Hilbert space $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ in the sense that we can write
$\mathbb{H}^{\text{eff}}=(-\beta\Delta_{y})\otimes 1+1\otimes\mathbb{D}_{\text{B}}+\Phi(h_{y})+E_{N}^{\text{pw}}$ (3.3)
with $\Phi(h_{y})\coloneqq c^{\ast}(h_{y})+c(h_{y})$ and
$c^{\ast}(h_{y})$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\ c_{\alpha}^{\ast}(k),$ (3.4)
$c(h_{y})$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{(h_{y})_{\alpha}(k)}\ c_{\alpha}(k),$ (3.5)
$\mathbb{D}_{\text{B}}$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)\quad\text{with }\epsilon_{\alpha}(k)=2k_{F}|k\cdot\omega_{\alpha}|$ (3.6)
describing the kinetic energy of the almost-bosonic pair excitations with linear dispersion relation. Note that by the Kato-Rellich theorem the effective Hamiltonian (3.1) is self-adjoint in its natural domain and generates a unitary time evolution.
To state an effective description of the time evolution of those excitations we compare the particle-hole transformed microscopic dynamics $R^{\ast}e^{-i\mathbb{H}t}R\psi$ with the effective time evolution $e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi$ in Hilbert space norm:
Theorem 3.1 (Effective dynamics of the system).
Assume that $\hat{V}\geq 0$ is compactly supported and satisfies $\hat{V}(-k)=\hat{V}(k)$ for all $k\in\mathbb{Z}^{3}$. Let $\lambda\in[k_{\textnormal{F}}^{-1/6},1]$ be the interaction parameter as introduced in (1.2) and take the number of patches to be $M=N^{\frac{16}{45}}$ with $\delta=\frac{2}{15}$ as introduced in (2.15). Then it holds for the initial state $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ that there is a $C>0$ depending only on the interaction $V$ such that for all $k_{\textnormal{F}}\geq 2$ and $t\geq 0$
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi\|\leq C\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)k_{\textnormal{F}}^{-\frac{1}{5}}.$
Remark 3.2.
Note that the right hand side of the bound is indeed small as long as $t\in\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$. The error $k_{\textnormal{F}}^{-1/5}$ is a result of the optimized choice of $M$ and $\delta$. The non-optimal error bound depends on the patch parameters and is given by
$C\left((1+\lambda^{-1})\lambda^{-1}(M^{-\frac{1}{2}}+MN^{-\frac{2}{3}+\delta})+(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\right)\left(e^{C\lambda k_{\textnormal{F}}t}-1\right).\$ (3.7)
3.2 Effective coherent state
The effective time evolved state from Theorem 3.1 can be further simplified. Our second result shows how the dynamics can be approximated on the level of states. In order to state the second main result, we introduce almost-bosonic coherent states.
Note that since $B\coloneqq c^{\ast}(\eta)-c(\eta)$ defines for all $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ a bounded operator and satisfies $B=-B^{\ast}$, the exponential operator $e^{B}$ is well-defined and is unitary.
Definition 3.3.
Define for $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ the Weyl operator
$W(\eta)\coloneqq e^{B}\coloneqq e^{c^{\ast}(\eta)-c(\eta)}.$ (3.8)
If $\eta\equiv\eta^{y}$ is additionally a bounded multiplication operator for each $y\in\mathbb{R}^{3}$, we call $W(\eta)\phi\otimes\Omega$ a coupled coherent state with $\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$.
Remark 3.4.
If $c^{\ast},c$ would satisfy the CCR without error, one could use the Baker-Campbell-Hausdorff formula to formally write
$W(\eta)\phi\otimes\Omega=$ $e^{-\|\eta\|^{2}/2}e^{c^{\ast}(\eta)}\phi\otimes\bigg{\{}1,0,\ldots,0,\ldots\bigg{\}}$
$=$ $e^{-\|\eta\|^{2}/2}e^{c^{\ast}(\eta)}\bigg{\{}\phi,0,\ldots,0,\ldots\bigg{\}}$
$=$ $e^{-\|\eta_{s}\|^{2}/2}\bigg{\{}\phi,\eta\phi,\frac{\eta^{\otimes 2}\phi}{\sqrt{2!}},\ldots,\frac{\eta^{\otimes n}\phi}{\sqrt{n!}},\ldots\bigg{\}}$ (3.9)
i.e. the coupled coherent state corresponds to a superposition of different particle number.
Remark 3.5.
The coupled coherent state from 3.3 satisfies the following well-known properties of the Weyl operator (cf. for example [BPS16, Chapter 3] or [FZ17, Appendix A]) up to certain error terms:
  • (shift property) For $\eta,\xi\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ it holds
    $c(\xi)W(\eta)\phi\otimes\Omega\simeq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\xi_{\alpha}(k)}\eta_{\alpha}(k)\phi\otimes\Omega,$ (3.10)
  • (expectation of the number operator) For $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ it holds
    $\langle W(\eta)\phi\otimes\Omega,\mathcal{N}W(\eta)\phi\otimes\Omega\rangle\simeq 2\|\eta\|^{2},$ (3.11)
  • (time derivative of the Weyl operator) For $\eta_{t}\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ differentiable in $t$ with derivative $\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ it holds: for all $t\in\mathbb{R}$:
    $\partial_{t}W(\eta_{t})\simeq\left(c^{\ast}(\dot{\eta}_{t})-c(\dot{\eta}_{t})+i\textnormal{Im}\langle\dot{\eta}_{t},\eta_{t}\rangle\right)W(\eta_{t}),$ (3.12)
We give rigorous statements on the error terms and proofs of the approximate properties in Lemma 5.1, Proposition 5.3, Lemma 5.5 of Section 5.
Consider now the following state for all times $t\in\mathbb{R}$
$\psi_{t}$ $\coloneqq e^{iP(t)}W(\eta_{t})\phi\otimes\Omega$ (3.13)
with $\text{with }P(t)$ $=2\text{Im}(\nu_{t})-E_{N}^{\text{pw}}t-\text{Im}\int_{0}^{t}\textnormal{d}s\langle\dot{\eta}_{s},\eta_{s}\rangle_{\Gamma}$ (3.14)
with the choices of
$\big{(}\eta_{s}\big{)}_{\alpha}(k)\coloneqq$ $\big{(}\eta_{s}^{y}\big{)}_{\alpha}(k)\coloneqq\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}(h_{y})_{\alpha}(k)=\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}\lambda\hat{V}(k)n_{\alpha}(k)e^{iky},$ (3.15)
$\nu_{s}\coloneqq$ $\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}+is\epsilon_{\alpha}(k)-1}{\epsilon_{\alpha}(k)^{2}}|(h_{y})_{\alpha}(k)|^{2}$ (3.16)
for all $k\in\Gamma,\alpha\in\mathcal{I}_{k}$. Due to $\big{(}\eta_{s}\big{)}_{\alpha}(k)=-ie^{-is\epsilon_{\alpha}(k)/2}\frac{\text{sin}\left(\epsilon_{\alpha}(k)s/2\right)}{\epsilon_{\alpha}(k)/2}(h_{y})_{\alpha}(k)$ and Lemma A.2 the norm is bounded for all $s\in\mathbb{R}$
$\|\eta_{s}\|$ $\equiv\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\eta_{s}\big{)}_{\alpha}(k)|^{2}\bigg{)}^{1/2}\leq\min\left\{\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}\lambda k_{\textnormal{F}}s,\sqrt{2\pi}\|\hat{V}\|_{2}\lambda\log(4k_{\textnormal{F}}s+2)\right\}$ (3.17)
as shown later in Lemma 5.6 and similarly for $|\nu_{s}|$.
In particular, it holds $(\eta_{0},\nu_{0})=(0,0)$ and $\lim_{\epsilon_{\alpha}(k)\to 0}(\eta_{s},\nu_{s})=(-ish_{y},0)$ and therefore $\psi_{0}=\phi\otimes\Omega$. Thus, as mentioned before, $\eta\equiv\eta^{y}$ as defined in (3.15) corresponds to an interaction term with a bounded multiplication operator acting on $L^{2}(\Lambda,\textnormal{d}y)$. The state $\psi_{t}=e^{iP(t)}W(\eta_{t})\phi\otimes\Omega$ can therefore be seen as a time-dependent coupled coherent state.
The following theorem states that $\psi_{t}$ approximately corresponds to the effective time evolution generated by $\mathbb{H}^{\textnormal{eff}}$.
Theorem 3.6 (Effective coherent dynamics).
Consider the initial state $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\sum_{i=1}^{3}\big{(}\|\partial_{y_{i}}\phi\|+\|\partial_{y_{i}}^{2}\phi\|\big{)}\leq c<\infty$ for a constant $c>0$. Under the assumptions of Theorem 3.1, there exists a constant $C>0$ and a function $Q:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ monotonically increasing with $Q(0)=0$ such that for all $t\geq 0$
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|\leq CQ(\lambda k_{\textnormal{F}}t)\max\{k_{\textnormal{F}}^{-\frac{2}{15}},c\beta t\}$
with $P(t)$ given by (3.14), $\eta_{t}$ given by (5.30), $C$ and $Q$ depending only on the interaction $V$.
Remark 3.7.
Note that the upper bound above is indeed meaningful in the sense that the bound is small for $t\in\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$ and $\beta\in o(\lambda k_{\textnormal{F}})$. The latter condition together with the upper bound for $\|\Delta_{y}W(\eta_{t})\phi\otimes\Omega\|$ from Lemma 5.9 ensures that the contribution from the kinetic energy term $h_{0}=-\beta\Delta_{y}$ of the impurity particle remains negligible on the relevant time scale. This seems to be crucial in our approach since otherwise $h_{0}$ would generate non-trivial correlations and make the coherent state form inapplicable.
Remark 3.8.
Due to the explicit formulation of the coupled coherent state provided in (3.15), we can quantify the number of collective excitations over time. More concretely, using (3.11) the term $\|\eta_{t}\|^{2}$, which is calculated in Lemma 5.6, represents the expected number of excitations generated by the interaction with the impurity. Assuming that the impurity, together with its excitations, can be interpreted as polaron-like quasi-particle, the quantity $\|\eta_{t}\|^{2}$ offers insight into the quasi-particle formation process. As displayed in Figure 1 on page 1 the graph shows a parabolic growth followed by a logarithmic increase as qualitative feature.
Figure 1: Plot of $\|\eta_{t}\|^{2}$ using Lemma 5.6 with constant $\hat{V}$ and $\lambda=1$. As a qualitative feature one observes a parabolic growth followed by a logarithmic increase.
Remark 3.9.
The proof is based on the following observation: It holds by virtue of Duhamel’s formula
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|$
$=\|\psi-e^{i\mathbb{H}^{\text{eff}}t}e^{-iE_{N}^{\text{pw}}t}e^{i2\text{Im}(\nu_{t})}e^{-i\text{Im}\int_{0}^{t}\textnormal{d}s\langle\dot{\eta}_{s},\eta_{s}\rangle}W(\eta_{t})\psi\|$ (3.18)
$=\|\int_{0}^{t}\textnormal{d}s\ e^{i\mathbb{H}^{\text{eff}}s}e^{iP(s)}\left\{\big{(}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\right\}\|$ (3.19)
$\leq\int_{0}^{t}\textnormal{d}s\|\big{(}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\|.$ (3.20)
If the collective operators $c_{\alpha}^{\ast}(k)$ and $c_{\alpha}(k)$ were exactly bosonic, i.e. the CCR held without error, we could use the shift property (3.10) of the Weyl operator to commute $c_{\alpha}(k)$ to the vacuum $\Omega$ with the cost of some inner product terms. In this case we would observe with the short-hand notation $c^{\ast}c(\epsilon)\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)$ and by applying (3.12) that
$\big{[}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-i\left\{c^{\ast}(\dot{\eta}_{s})-c(\dot{\eta}_{s})\right\}\big{]}W(\eta_{s})\psi$
$=$ $\big{[}h_{0}+c^{\ast}c(\epsilon)+c^{\ast}(h_{y})+c(h_{y})+2\text{Im}(\dot{\nu}_{s})+c^{\ast}(-i\dot{\eta}_{s})-c(i\dot{\eta}_{s})\big{]}W(\eta_{s})\phi\otimes\Omega$
$=$ $\big{[}h_{0}+c^{\ast}(\epsilon\eta_{s})+c^{\ast}(h_{y})+\langle h_{y},\eta_{s}\rangle+2\text{Im}(\dot{\nu}_{s})+c^{\ast}(-i\dot{\eta}_{s})-\langle i\dot{\eta}_{s},\eta_{s}\rangle\big{]}W(\eta_{s})\phi\otimes\Omega$
$=$ $h_{0}W(\eta_{s})\phi\otimes\Omega$ (3.21)
is exactly vanishing up to the kinetic energy term $h_{0}$ of the impurity particle since $\eta_{t}$ as defined in (3.15) solves the ODE $\epsilon\eta_{t}+h_{y}=i\dot{\eta}_{t}$ and $\nu_{t}$ as defined in (3.16) absorbs all scalar terms.
Remark 3.10.
Note that as long as $\|\eta_{t}\|$ is of order 1, the coupled coherent state is different from a free decoupled dynamics with $\psi_{t}^{\textnormal{free}}\sim e^{i\beta\Delta_{y}t}\phi\otimes\Omega$. Vice versa, if $\|\eta_{t}\|\in o(1)$ in terms of $k_{\textnormal{F}}$ one can easily see with a Duhamel argument, analogue to the previous remark, that $\|\psi_{t}^{\textnormal{free}}-e^{iP(t)}W(\eta_{t})\psi\|\to 0$ for large $k_{\textnormal{F}}$. This is because all the terms appearing in $\partial_{t}W(\eta_{t})$ from (3.12) can be bounded in terms of $\|\eta_{t}\|$. Using (3.17) we can identify the following cases:
  • $\lambda=1,t\in o(k_{\textnormal{F}}^{-1})$: $\|\eta_{t}\|\in o(1)$ with respect to $k_{\textnormal{F}}$
    The time scale is too short for forming excitations thus $\psi_{t}^{\textnormal{free}}$ is a good approximation. Note that this is compatible with [MP21] as mentioned in Subsection 1.2.
  • $\lambda=1,t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$: $\|\eta_{t}\|\in\mathcal{O}(1)$ with respect to $k_{\textnormal{F}}$
    On this time scale exitations of the Fermi gas become important and thus the free decoupled evolution is not a good approximation anymore. We refer to the proof of Corollary 3.11 on how to show rigorously that the derived effective description is relevant on this time scale.
  • $\lambda\in o(1),0\leq t\in\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$: $\|\eta_{t}\|\in o(1)$ with respect to $k_{\textnormal{F}}$
    Also here, $\psi_{t}^{\textnormal{free}}$ is a good description. Note that the number of excitations grows only logarithmically in time. Thus for any $\lambda\in o(1)$ on gets $\|\eta_{t}\|\in o(1)$. In the case of $\lambda\in o(1)$ we expect this result of free decoupling to hold for even longer times, possibly of order 1. In order to show this expected behavior, we expect a method different from ours to be more suitable, for example a perturbative expansion in the spirit of [MP21] to higher orders to exploit the small coupling.
Now by virtue of the previous statement we are able to show that the linear coupling term of the effective Hamiltonian is essential for the effective description and cannot be neglected on an approximate level for $\lambda=1$ and $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$.
Let
$\widetilde{\mathbb{H}^{\text{eff}}}\coloneqq\mathbb{H}^{\textnormal{eff}}-c^{\ast}(h_{y})-c(h_{y})\equiv\beta(-\Delta_{y})+\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)+E_{N}^{\text{pw}}$ (3.22)
be the effective Hamiltonian without the linear coupling.
Corollary 3.11.
Under the assumptions of Theorem 3.6 with $\lambda=1$ and a $T>0$, there exists a monotonically increasing function $C:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ only depending on $V$ such that for all $t\in[0,T]$
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\widetilde{\mathbb{H}^{\textnormal{eff}}}t}\psi\|\geq C(t)-\mathcal{O}\left(\max\{k_{\textnormal{F}}^{-\frac{2}{15}},\beta k_{\textnormal{F}}^{-1}\}\right).$
In particular it holds $C(t)\in\mathcal{O}(1)$ with respect to $k_{\textnormal{F}}$ for all $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$.
4 Proof of Theorem 3.1
In all of our estimates, we need to control the number operator $\mathcal{N}$ acting on the time evolved state $\psi_{t}=e^{-i\mathbb{H}^{\textnormal{eff}}t}\phi\otimes\Omega$. The following statement shows that the effective time evolution preserves the order of magnitude of the number of excitations:
Proposition 4.1 (Effective time evolution of the number operator).
There exists a constant $C>0$ only depending on $V$ such that it holds for all $t\in\mathbb{R},n\in\mathbb{N}$ and $\psi\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
$\langle e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi,(\mathcal{N}+1)^{n}e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi\rangle\leq e^{nC\lambda k_{\textnormal{F}}t}\langle\psi,(\mathcal{N}+3)^{n}\psi\rangle.$
Proof.
We want to use Grönwall’s lemma and therefore estimate the derivative
$\bigg{|}i\partial_{t}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\bigg{|}$
$\leq$ $\bigg{|}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,\sum_{j=0}^{n-1}(\mathcal{N}+3)^{j}[\mathcal{N},\mathbb{H}^{\text{eff}}](\mathcal{N}+3)^{n-j-1}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\bigg{|}$
$=$ $\bigg{|}2\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\sum_{j=0}^{n-1}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{j}\big{(}c_{\alpha}^{\ast}(k)-c_{\alpha}(k)\big{)}(\mathcal{N}+3)^{n-j-1}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\bigg{|}$ (4.1)
We split the difference of $\big{(}c_{\alpha}^{\ast}(k)-c_{\alpha}(k)\big{)}$ and consider the term with $c_{\alpha}^{\ast}(k)$ first. Insert here $\textnormal{id}=(\mathcal{N}+1)^{\frac{n}{2}-1-j}(\mathcal{N}+1)^{j+1-\frac{n}{2}}$ between $(\mathcal{N}+3)^{j}$ and $c_{\alpha}^{\ast}(k)$ and use the commutation $\mathcal{N}c_{\alpha}^{\ast}(k)=c_{\alpha}^{\ast}(k)(\mathcal{N}+2)$ to obtain
$(\mathcal{N}+3)^{j}c_{\alpha}^{\ast}(k)(\mathcal{N}+3)^{n-j-1}=(\mathcal{N}+3)^{j}(\mathcal{N}+1)^{\frac{n}{2}-1-j}c_{\alpha}^{\ast}(k)(\mathcal{N}+3)^{\frac{n}{2}}.$ (4.2)
We introduce the notation $\xi_{j}\coloneqq(\mathcal{N}+1)^{\frac{n}{2}-1-j}(\mathcal{N}+3)^{j}e^{-i\mathbb{H}^{\text{eff}}t}\psi$ and $\tilde{\xi}\coloneqq(\mathcal{N}+3)^{\frac{n}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi$ to estimate
$\bigg{|}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\sum_{j=0}^{n-1}\langle\xi_{j},c_{\alpha}^{\ast}(k)\tilde{\xi}\rangle\bigg{|}$
$\leq$ $\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|(h_{y})_{\alpha}(k)|\sum_{j=0}^{n-1}\|c_{\alpha}(k)\xi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $C\lambda\sum_{k\in\Gamma}|\hat{V}(k)|\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}\bigg{)}^{1/2}\sum_{j=0}^{n-1}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\xi_{j}\|^{2}\bigg{)}^{1/2}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\sum_{k\in\Gamma}|\hat{V}(k)|\sum_{j=0}^{n-1}\|\mathcal{N}^{1/2}\xi_{j}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\sum_{j=0}^{n-1}\|(\mathcal{N}+1)\xi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $Cn\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle$ (4.3)
where we used Lemma A.4 and Lemma A.1 in the fourth line and the operator inequality $\mathcal{N}^{1/2}\leq(\mathcal{N}+1)$ in the fifth line. Note that $|k|<C$ for all $k\in\Gamma$ since $\Gamma\subset\textnormal{supp}\hat{V}$ and $\hat{V}$ has bounded support by assumption.
The second term with $c_{\alpha}(k)$ can be treated by inserting $\textnormal{id}=(\mathcal{N}+1)^{\frac{n}{2}-j}(\mathcal{N}+1)^{j-\frac{n}{2}}$ and using the commutation $c_{\alpha}(k)\mathcal{N}=(\mathcal{N}+2)c_{\alpha}(k)$.
$(\mathcal{N}+3)^{j}c_{\alpha}(k)(\mathcal{N}+3)^{n-j-1}=(\mathcal{N}+3)^{\frac{n}{2}}c_{\alpha}(k)(\mathcal{N}+1)^{j-\frac{n}{2}}(\mathcal{N}+3)^{n-j-1}.$ (4.4)
We introduce the notation $\chi_{j}\coloneqq(\mathcal{N}+1)^{j-\frac{n}{2}}(\mathcal{N}+3)^{n-j-1}e^{-i\mathbb{H}^{\text{eff}}t}\psi$
$\bigg{|}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\sum_{j=0}^{n-1}\langle\tilde{\xi},c_{\alpha}(k)\chi_{j}\rangle\bigg{|}$
$\leq$ $\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|(h_{y})_{\alpha}(k)|\sum_{j=0}^{n-1}\|c_{\alpha}(k)\chi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $C\lambda\sum_{k\in\Gamma}|\hat{V}(k)|\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}\bigg{)}^{1/2}\sum_{j=0}^{n-1}\bigg{(}\sum_{\alpha}\|c_{\alpha}(k)\chi_{j}\|^{2}\bigg{)}^{1/2}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\sum_{k\in\Gamma}|\hat{V}(k)|\sum_{j=0}^{n-1}\|\mathcal{N}^{1/2}\chi_{j}\|\tilde{\xi}\|$
$\leq$ $C\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\sum_{j=0}^{n-1}\|(\mathcal{N}+1)\chi_{j}\|\ \|\tilde{\xi}\|$
$\leq$ $Cn\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle.$ (4.5)
Altogether it follows with the Grönwall’s lemma that
$\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathcal{N}+3)^{n}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\leq\exp\big{(}Cn\lambda k_{\textnormal{F}}\|\hat{V}\|_{1}t\big{)}\langle\psi,(\mathcal{N}+3)^{n}\psi\rangle$ (4.6)
which is the desired result since $\|\hat{V}\|_{1}<C$. ∎
The following statement shows that the fermionic kinetic energy term (2.8) can be approximated by the almost-bosonic kinetic energy term.
Proposition 4.2 (Approximation of the kinetic energy).
There exists a constant $C>0$ only depending on $V$ such that it holds for all $t\in\mathbb{R}$ and $\psi\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
$\left|\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi\|-\|(\mathbb{H}_{0}-\mathbb{D}_{B})\psi\|\right|$
$\leq C(1+\lambda^{-1})N^{\frac{1}{3}}\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)\left(M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)$ (4.7)
using Proposition 4.1.
Remark 4.3.
Note that in the case of $\psi\equiv\phi\otimes\Omega$ it holds $(\mathbb{H}_{0}-\mathbb{D}_{B})\psi=0$ and $\|(\mathcal{N}+3)^{2}\psi\|=9\leq C$. Also note that $\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)=\lambda k_{\textnormal{F}}t+\mathcal{O}((\lambda k_{\textnormal{F}}t)^{2})$.
Proof.
It holds
$\left|\partial_{t}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\right|$
$=\left|\partial_{t}\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$=\left|\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,[(\mathbb{H}_{0}-\mathbb{D}_{B})^{2},\mathbb{H}^{\text{eff}}]e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$=\left|\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,\left((\mathbb{H}_{0}-\mathbb{D}_{B})[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]+[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}](\mathbb{H}_{0}-\mathbb{D}_{B})\right)e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$\leq 2\left|\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$ (4.8)
and make use of
$[(\mathbb{H}_{0}-\mathbb{D}_{B}),c_{\alpha}^{\ast}(k)]$ $\eqqcolon\mathfrak{E}_{\alpha}^{\text{lin}}(k)^{\ast}-\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)^{\ast}\eqqcolon\mathfrak{E}_{\alpha}(k)^{\ast},$ (4.10)
${}[(\mathbb{H}_{0}-\mathbb{D}_{B}),c_{\alpha}(k)]$ $=-\mathfrak{E}_{\alpha}(k)$ (4.11)
to arrive at
$[(\mathbb{H}_{0}-\mathbb{D}_{B}),\mathbb{H}^{\text{eff}}]$ $=[(\mathbb{H}_{0}-\mathbb{D}_{B}),c^{\ast}c(\epsilon)+c^{\ast}(h_{y})+c(h_{y})]$ (4.12)
$=\langle\mathfrak{E},c(\epsilon)\rangle_{\Gamma}-\langle c(\epsilon),\mathfrak{E}\rangle_{\Gamma}+\langle\mathfrak{E},h_{y}\rangle_{\Gamma}-\langle h_{y},\mathfrak{E}\rangle_{\Gamma}$ (4.13)
with the short notation $\langle A,B\rangle_{\Gamma}\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}A_{\alpha}^{\ast}(k)B_{\alpha}(k)$.
Bounds for the error terms are readily provided in Lemma A.5 and Lemma A.6 of the appendix:
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|^{2}$ $\leq C\left(N^{\frac{1}{3}}M^{-\frac{1}{2}}\right)^{2}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|^{2},$ (4.14)
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$ $\leq C\left(k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2}.$ (4.15)
Furthermore use the bounds $\sum_{\alpha\in\mathcal{I}_{k}}c_{\alpha}^{\ast}(k)c_{\alpha}(k)\leq\mathcal{N}$, $\epsilon_{\alpha}(k)\leq Ck_{\textnormal{F}}$ and that $\mathfrak{E}^{\textnormal{B}}_{\alpha}(k),\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k),c_{\alpha}(k)$ all annihilate exactly two fermions to estimate
$\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})\langle c(\epsilon),\mathfrak{E}^{\textnormal{lin}}-\mathfrak{E}^{\textnormal{B}}\rangle_{\Gamma}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle$
$\leq$ $\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle\epsilon_{\alpha}(k)c_{\alpha}(k)(\mathcal{N}+1)^{-1/2}(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$+\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle\epsilon_{\alpha}(k)c_{\alpha}(k)(\mathcal{N}+1)^{-1/2}(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$ (4.16)
$\leq$ $\sum_{k\in\Gamma}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\epsilon_{\alpha}(k)c_{\alpha}(k)(\mathcal{N}+1)^{-1/2}(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}$
×$\quad\times\left(\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}+\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)(\mathcal{N}+1)^{1/2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}\right)$ (4.17)
$\leq$ $Ck_{\textnormal{F}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\left(N^{\frac{1}{3}}M^{-\frac{1}{2}}\|(\mathcal{N}+1)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+1)^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right)$ (4.18)
$\leq$ $C^{2}k_{\textnormal{F}}N^{\frac{1}{3}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\left(M^{-\frac{1}{2}}\|(\mathcal{N}+1)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+1)^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right).$ (4.19)
The term $\langle c(\epsilon),\mathfrak{E}\rangle_{\Gamma}$ can be treated analogously. The other terms can be estimated similarly using $|(h_{y})_{\alpha}(k)|\leq C\lambda|\hat{V}(k)||n_{\alpha}(k)|$ and Cauchy-Schwarz inequality with (A.1)
$\langle e^{-i\mathbb{H}^{\text{eff}}t}\psi,(\mathbb{H}_{0}-\mathbb{D}_{B})\langle\mathfrak{E}^{\textnormal{lin}}-\mathfrak{E}^{\textnormal{B}},h_{y}\rangle_{\Gamma}e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle$
$\leq$ $C\lambda\sum_{k\in\Gamma}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle n_{\alpha}(k)(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$
$+C\lambda\sum_{k\in\Gamma}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\left|\langle n_{\alpha}(k)(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi,\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)e^{-i\mathbb{H}^{\text{eff}}t}\psi\rangle\right|$ (4.20)
$\leq$ $C\lambda k_{\textnormal{F}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\sum_{k\in\Gamma}|\hat{V}(k)|\Bigg{\{}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|^{2}\bigg{)}^{1/2}$ (4.21)
[…]
$\leq$ $C\lambda\|\hat{V}\|_{1}k_{\textnormal{F}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\bigg{\{}M^{-\frac{1}{2}}N^{\frac{1}{3}}\|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.23)
[…]
$\leq$ $C\lambda k_{\textnormal{F}}M^{-\frac{1}{2}}N^{\frac{1}{3}}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\bigg{\{}\|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.24)
[…]
Thus we derive
$\partial_{t}\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$
$\leq$ $C\lambda k_{\textnormal{F}}M^{-\frac{1}{2}}N^{\frac{1}{3}}\left(\|(\mathcal{N}+1)e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+1)^{2}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right)$
$+C\lambda k_{\textnormal{F}}N^{\frac{1}{3}}\left(\|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}e^{-i\mathbb{H}^{\text{eff}}t}\psi\|\right)$ (4.26)
$\leq$ $C^{2}k_{\textnormal{F}}N^{\frac{1}{3}}\left(e^{C\lambda k_{\textnormal{F}}t}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}e^{2C\lambda k_{\textnormal{F}}t}\|(\mathcal{N}+3)^{2}\psi\|\right)$
$+C\lambda k_{\textnormal{F}}N^{\frac{1}{3}}\left(e^{\frac{1}{2}C\lambda k_{\textnormal{F}}t}\|(\mathcal{N}+3)^{\frac{1}{2}}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}e^{\frac{3}{2}C\lambda k_{\textnormal{F}}t}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|\right)$ (4.27)
where we used the Grönwall bound from Proposition 4.1 in the last inequality. Now integrating over $t$ yields
$\|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}t}\psi\|-\|(\mathbb{H}_{0}-\mathbb{D}_{B})\psi\|$
$\leq$ $C^{2}\lambda^{-1}M^{-\frac{1}{2}}N^{\frac{1}{3}}\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)\left(\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)$
$+CM^{-\frac{1}{2}}N^{\frac{1}{3}}\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)\left(\|(\mathcal{N}+3)^{\frac{1}{2}}\psi\|+k_{\textnormal{F}}M^{\frac{3}{2}}N^{-1+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|\right)$ (4.28)
which corresponds to the desired result. ∎
We are now ready to give the proof of the main theorem.
Proof of of Theorem 3.1.
We employ Duhamel’s formula for $t\mapsto e^{i\mathbb{H}t}Re^{-i\mathbb{H}^{\text{eff}}t}$:
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ $=\|R\psi-e^{i\mathbb{H}t}Re^{-i\mathbb{H}^{\text{eff}}t}\psi\|$ (4.29)
$=\|\int_{0}^{t}\textnormal{d}s\ e^{i\mathbb{H}s}\big{(}\mathbb{H}R-R\mathbb{H}^{\text{eff}}\big{)}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$ (4.30)
$\leq\int_{0}^{t}\textnormal{d}s\ \|(R^{\ast}\mathbb{H}R-\mathbb{H}^{\text{eff}})e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$ (4.31)
where we used that $R$ is unitary.
From our considerations in (2.12) it follows that
$R^{\ast}\mathbb{H}R-\mathbb{H}^{\textnormal{eff}}$ $=-\beta\Delta_{y}+\mathbb{H}_{0}+b^{\ast}(\tilde{h}_{y})+b(\tilde{h}_{y})+E_{N}^{\text{pw}}+\mathcal{E}-\mathbb{H}^{\textnormal{eff}}$
$=\mathbb{H}_{0}-\mathbb{D}_{B}+b^{\ast}(\tilde{h}_{y})-c^{\ast}(h_{y})+b(\tilde{h}_{y})-c(h_{y})+\mathcal{E}.$ (4.32)
The interaction terms can be approximated in the sense of Lemma A.7:
$\|\left(b^{\ast}(\tilde{h}_{y})-c^{\ast}(h_{y})\right)\psi\|$ $\leq\lambda\sum_{k\in\Gamma}|\hat{V}(k)|\ \|\big{(}b(k)-\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)c_{\alpha}(k)\big{)}\psi\|$
$\leq C\lambda N^{\frac{1}{3}}\|\hat{V}\|_{1}(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|.$
Therefore by combining the bounds from Proposition 4.2 with $(\mathbb{H}_{0}-\mathbb{D}_{B})\psi=0$, Lemma A.8 with $\|\hat{V}\|_{1}<C$ and Proposition 4.1 we obtain
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|$
$\leq\int_{0}^{t}\textnormal{d}s\ \|(\mathbb{H}_{0}-\mathbb{D}_{B})e^{-i\mathbb{H}^{\text{eff}}s}\psi\|+\int_{0}^{t}\textnormal{d}s\ \|\mathcal{E}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\quad+\int_{0}^{t}\textnormal{d}s\ \|\left(b^{\ast}(\tilde{h}_{y})-c^{\ast}(h_{y})\right)e^{-i\mathbb{H}^{\text{eff}}s}\psi\|+\int_{0}^{t}\textnormal{d}s\ \|\left(b(\tilde{h}_{y})-c(h_{y})\right)e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\leq C(1+\lambda^{-1})N^{\frac{1}{3}}\left(M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+k_{\textnormal{F}}MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)\int_{0}^{t}\textnormal{d}s\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)$
$\quad+C\lambda\int_{0}^{t}\textnormal{d}s\ \|\mathcal{N}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\quad+C\lambda N^{\frac{1}{3}}(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\int_{0}^{t}\textnormal{d}s\ \|(\mathcal{N}+1)^{\frac{1}{2}}e^{-i\mathbb{H}^{\text{eff}}s}\psi\|$
$\leq C(1+\lambda^{-1})\lambda^{-1}N^{\frac{1}{3}}\left(k_{\textnormal{F}}^{-1}M^{-\frac{1}{2}}\|(\mathcal{N}+3)\psi\|+MN^{-1+\delta}\|(\mathcal{N}+3)^{2}\psi\|\right)\left(e^{C\lambda k_{\textnormal{F}}t}-\lambda k_{\textnormal{F}}t-1\right)$
$\quad+Ck_{\textnormal{F}}^{-1}\|(\mathcal{N}+3)\psi\|\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)$
$\quad+C(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\|(\mathcal{N}+3)^{\frac{1}{2}}\psi\|\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)$
$\leq C\tilde{C}\|(\mathcal{N}+3)^{2}\psi\|\left(e^{C\lambda k_{\textnormal{F}}t}-1\right).$ (4.33)
The prefactor is given by
$\tilde{C}=\max\left\{(1+\lambda^{-1})\lambda^{-1}k_{\textnormal{F}}^{-1}N^{\frac{1}{3}}\left(M^{-\frac{1}{2}}+k_{\textnormal{F}}MN^{-1+\delta}\right),k_{\textnormal{F}}^{-1},(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\right\}$ (4.34)
where we optimize over $M$ and $\delta$ with $\lambda\geq k_{\textnormal{F}}^{-\frac{1}{6}}$ to obtain the desired result. ∎
5 Proof of Theorem 3.6
Properties of almost-bosonic coherent states
We first show the following rigorous properties of the Weyl operator which was introduced in 3.3:
Lemma 5.1 (Approximate shift property).
Let $\eta,\xi\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ and $W_{\sigma}(\eta)\coloneqq e^{\sigma B}\coloneqq e^{\sigma c^{\ast}(\eta)-\sigma c(\eta)}$ for all $\sigma\in[0,1]$, then it holds
$W_{\sigma}(\eta)^{\ast}c(\xi)W_{\sigma}(\eta)$ $=c(\xi)+\sigma\langle\xi,\eta\rangle+\langle\xi,\mathcal{R}^{\sigma}\rangle_{\Gamma},$
$W_{\sigma}(\eta)^{\ast}c^{\ast}(\xi)W_{\sigma}(\eta)$ $=c^{\ast}(\xi)+\sigma\langle\eta,\xi\rangle+\langle\mathcal{R}^{\sigma},\xi\rangle_{\Gamma},$
with $\langle\cdot,\cdot\rangle_{\Gamma}:\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})\times\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})\to\mathbb{C}$ given by
$\langle\xi,\mathcal{R}^{\sigma}\rangle_{\Gamma}\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\xi_{\alpha}(k)^{\ast}\mathcal{R}^{\sigma}_{\alpha}(k)$ (5.1)
and a $\sigma$-dependent error term $\mathcal{R}^{\sigma}_{\alpha}(k)\coloneqq\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}$.
Remark 5.2.
We will give an estimate for $\mathcal{R}$ to show that this term corresponds indeed to a small error. Note that it holds $\mathcal{E}_{\alpha}(k,l)=\mathcal{E}_{\alpha}(l,k)^{\ast}$ for all $l,k\in\Gamma$ and $\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}$ from Lemma A.3 and therefore
$\mathcal{R}^{\sigma}_{\gamma}(l)^{\ast}\xi_{\gamma}(l)$ $=\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma}\mathcal{E}_{\gamma}(l,k)\overline{\eta_{\gamma}(k)}\xi_{\gamma}(l)\right)e^{\tau B},$ (5.2)
$\overline{\xi_{\gamma}(l)}\mathcal{R}^{\sigma}_{\gamma}(l)$ $=\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma}\overline{\xi_{\gamma}(l)}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}.$ (5.3)
For $\xi=\eta$ the above equations coincide, i.e.
$\langle\eta,\mathcal{R}^{\sigma}\rangle_{\Gamma}=\langle\mathcal{R}^{\sigma},\eta\rangle_{\Gamma}$ (5.4)
from which it follows immediately that $\big{(}c^{\ast}(\eta)-c(\eta)\big{)}W_{\sigma}(\eta)=W_{\sigma}(\eta)\big{(}c^{\ast}(\eta)-c(\eta)\big{)}$, i.e. $[B,W_{\sigma}(\eta)]=0$.
Proof.
We observe that $W_{\sigma}(\eta)=e^{\sigma B}$ defines a strongly continuous one-parameter semigroup. Thus we can define a derivative and make use of Duhamel’s formula of the form
$e^{-\sigma B}c_{\gamma}(l)e^{\sigma B}=c_{\gamma}(l)+\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}[c_{\gamma}(l),B]e^{\tau B}.\$ (5.5)
The interested reader is referred to [Paz83, EN06] where definitions and properties of operator derivatives are discussed. The desired statement follows with the CCR as stated in (2.22)
$[c_{\gamma}(l),e^{\sigma B}]$ $=e^{\sigma B}\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}[c_{\gamma}(l),B]e^{\tau B}$
$=e^{\sigma B}\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\eta_{\gamma}(l)+\sum_{k\in\Gamma}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}$
$=\sigma\eta_{\gamma}(l)e^{\sigma B}+e^{\sigma B}\mathcal{R}^{\sigma}_{\gamma}(l)$ (5.6)
Since $c(\xi)\equiv\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\xi_{\alpha}(k)}c_{\alpha}(k)$ is linear the result follows from the above identity. ∎
The following statement shows that the number of particles in the state $W(\eta)\phi\otimes\Omega$ corresponds to a random variable with expectation approximately being $2\|\eta\|^{2}$.
Proposition 5.3 (Expectation of the number operator).
Let $\zeta=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$, then it holds for all $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=2\|\eta\|^{2}+4\int_{0}^{1}\textnormal{d}\sigma\langle\zeta,\langle\eta,\mathcal{R}^{\sigma}\rangle_{\Gamma}\zeta\rangle.$
Proof.
Using $W^{\ast}W=\textnormal{id}$ yields for all $\zeta\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\|\zeta\|=1$
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=\langle W(\eta)\zeta,[\mathcal{N},W(\eta)]\zeta\rangle+\langle\zeta,\mathcal{N}\zeta\rangle.$ (5.7)
We use Duhamel’s formula to calculate
$[\mathcal{N},e^{B}]$ $=e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}[\mathcal{N},B]e^{\tau B}$
$=2e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(c^{\ast}(\eta)+c(\eta)\right)e^{\tau B}$
$=2e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(B+2c(\eta)\right)e^{\tau B}$
$=2e^{B}B+4e^{B}\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}c(\eta)e^{\tau B}$ (5.8)
where we used (2.24). Therefore
$\langle W(\eta)\zeta,[\mathcal{N},W(\eta)]\zeta\rangle$ $=2\langle\zeta,B\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\ \langle e^{\tau B}\zeta,c(\eta)e^{\tau B}\zeta\rangle$
$=2\langle\zeta,B\zeta\rangle+2\|\eta\|^{2}+4\langle\zeta,c(\eta)\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{\Gamma}\zeta\rangle$ (5.9)
where we used the shift property Lemma 5.1 and that $e^{\tau B}$ is unitary
$\langle e^{\tau B}\zeta,c(\eta)e^{\tau B}\zeta\rangle$ $=\langle e^{\tau B}\zeta,[c(\eta),e^{\tau B}]\zeta\rangle+\langle\zeta,c(\eta)\zeta\rangle$
$=\tau\|\eta\|^{2}+\langle\zeta,\langle\eta,\mathcal{R}^{1}\rangle_{\Gamma}\zeta\rangle+\langle\zeta,c(\eta)\zeta\rangle.$ (5.10)
Inserting (5.9) and (5.8) into (5.7) we obtain
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=2\|\eta\|^{2}+2\langle\zeta,\left(c^{\ast}(\eta)+c(\eta)\right)\zeta\rangle+\langle\zeta,\mathcal{N}\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{\Gamma}\zeta\rangle.$ (5.11)
The desired result holds for $\zeta=\phi\otimes\Omega$ since $c(\eta)\phi\otimes\Omega=0$.
For later purposes, we can bound the expectation of the number operator in the following way:
Proposition 5.4 (Stability of the number operator).
Let $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$. There exists a constant $C>0$ such that it holds for all $\tau\in[-1,1],n\in\mathbb{N}$ and $\zeta\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
$\langle e^{\tau B}\zeta,(\mathcal{N}+1)^{n}e^{\tau B}\zeta\rangle\leq e^{C\|\eta\|n|\tau|}\langle\zeta,(\mathcal{N}+3)^{n}\zeta\rangle.$
Proof.
The proof works analogously to the proof of Proposition 4.1 with a Grönwall argument and $B$ instead of $\mathbb{H}^{\textnormal{eff}}$ which is given later. Note that
$[\mathcal{N},B]=[\mathcal{N},c^{\ast}(\eta)-c(\eta)]=2\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\eta_{\alpha}(k)\big{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(k)\big{)}$ (5.12)
where we again used (2.24). The result is than obtained by using the same estimates with $\|\eta\|$ taking the role of $\|h_{y}\|$. ∎
Lemma 5.5.
Let $\eta_{t}\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ be differentiable in $t$ with derivative $\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ for all $t\in\mathbb{R}$. Then it holds for all $t\in\mathbb{R}$
$\partial_{t}W(\eta_{t})=\left(c^{\ast}(\dot{\eta}_{t})-c(\dot{\eta}_{t})+i\textnormal{Im}\langle\dot{\eta}_{t},\eta_{t}\rangle\right)W(\eta_{t})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{(1-\tau)}(\eta_{t})\textnormal{Im}\langle\dot{\eta}_{t},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{t})$
with the shorthand notation $\textnormal{Im}\langle A,B\rangle_{\Gamma}\coloneqq-\frac{i}{2}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(A_{\alpha}^{\ast}(k)B_{\alpha}(k)-B_{\alpha}^{\ast}(k)A_{\alpha}(k))$
Proof.
For arbitrary $s\in\mathbb{R}$ it holds
$W(\eta_{s})^{\ast}\partial_{s}W(\eta_{s})$ $=\left.e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}\right|_{\tau=0}^{\tau=1}=\int_{0}^{1}\textnormal{d}\tau\ \partial_{\tau}\left(e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}\right)$
$=\int_{0}^{1}\textnormal{d}\tau\left(-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\partial_{\tau}e^{\tau B_{s}}\right)$ (5.13)
$=\int_{0}^{1}\textnormal{d}\tau\left(-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\left(B_{s}e^{\tau B_{s}}\right)\right)$ (5.14)
$=\int_{0}^{1}\textnormal{d}\tau\left(-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}+e^{-\tau B_{s}}B_{s}\partial_{s}e^{\tau B_{s}}\right)$ (5.15)
$=\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}.$ (5.16)
Thus
$\partial_{s}W(\eta_{s})$ $=\int_{0}^{1}\textnormal{d}\tau\ e^{(1-\tau)B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}$
$=\int_{0}^{1}\textnormal{d}\tau\ \left(\partial_{s}B_{s}\right)e^{(1-\tau)B_{s}}e^{\tau B_{s}}+\int_{0}^{1}\textnormal{d}\tau\ \left[e^{(1-\tau)B_{s}},\partial_{s}B_{s}\right]e^{\tau B_{s}}$ (5.17)
$=\left(\partial_{s}B_{s}\right)e^{B_{s}}+\int_{0}^{1}\textnormal{d}\tau\ \left[e^{(1-\tau)B_{s}},\partial_{s}B_{s}\right]e^{\tau B_{s}}.$ (5.18)
With
$\partial_{s}B_{s}=\partial_{s}\left\{c^{\ast}(\eta_{s})-c(\eta_{s})\right\}=c^{\ast}(\dot{\eta}_{s})-c(\dot{\eta}_{s})$ (5.19)
from the linearity of $c(\eta_{s})$ it follows
$\left[e^{(1-\tau)B_{s}},\partial_{s}B_{s}\right]e^{\tau B_{s}}$
$=\left([W_{(1-\tau)}(\eta_{s}),c^{\ast}(\dot{\eta}_{s})]-[W_{(1-\tau)}(\eta_{s}),c(\dot{\eta}_{s})]\right)W_{\tau}(\eta_{s})$ (5.20)
$=W_{(1-\tau)}(\eta_{s})\bigg{(}(1-\tau)\langle\dot{\eta}_{s},\eta_{s}\rangle-(1-\tau)\langle\eta_{s},\dot{\eta}_{s}\rangle$ (5.21)
[…]
$=W(\eta_{s})(1-\tau)2i\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle+2iW_{(1-\tau)}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}).$ (5.23)
Inserting the above identity yields
$\partial_{s}W(\eta_{s})=\left(c^{\ast}(\dot{\eta}_{s})-c(\dot{\eta}_{s})+i\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{(1-\tau)}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}).$ (5.24)
Proof of the main theorem
First, we collect some useful observations on the function $\eta_{s}$ in the form of the following two lemmata. We postpone the proofs to the end of the section in order to concentrate on presenting the proof of the main result.
Lemma 5.6.
Let $\eta_{s}$ be defined as in (3.15) for $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$, then it holds for all $s\in\mathbb{R}$
$\|\eta_{s}\|^{2}$ $=\pi\lambda^{2}\sum_{k\in\Gamma}\frac{\hat{V}(k)^{2}}{|k|}\left(\log(2k_{\textnormal{F}}|k|s)-\textnormal{Ci}(2k_{\textnormal{F}}|k|s)+\gamma\right)$
×$\times\left\{1+g(k_{\textnormal{F}}|k|s)\ \mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$
where $\gamma$ is the Euler-Mascheroni constant and $g:\mathbb{R}\to\mathbb{R}_{\geq 0}$ is a function independent of $k_{\textnormal{F}}$ and monotonically increasing.
Furthermore, define $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ by
$(y,x)\mapsto f_{y}(x)\coloneqq\min\{e^{\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}yx},e^{\sqrt{2\pi}\|\hat{V}\|_{2}(\log(18)+\frac{1}{9})y}e^{\frac{\sqrt{8\pi}}{9}\|\hat{V}\|_{2}yx}\}.$
Then there exists a $C>0$ independent of $k_{\textnormal{F}}$ such that for $c_{0}>0$ and $k_{\textnormal{F}}$ sufficiently large
$\|\eta_{s}\|$ $\leq\log\left(f_{1}(\lambda k_{\textnormal{F}}s)\right),$ (5.25)
$e^{c_{0}\|\eta_{s}\|}$ $\leq f_{c_{0}}(\lambda k_{\textnormal{F}}s).$ (5.26)
Remark 5.7.
Note that $f_{y}$ is for all $y\geq 0$ monotonically increasing with $f_{y}(0)=1$.
The previous statement is useful when combined with the following estimate:
Lemma 5.8.
There exists a constant $C>0$ only depending on $V$ such that it holds for all $s\in\mathbb{R}$, $n\in\mathbb{N}$, $\psi\in\mathcal{F}$
  1. (i).
    $\sum_{k\in\Gamma}\||k|^{n}\eta_{s}(k)\|_{l^{2}}\leq C\|\eta_{s}\|$,
  2. (ii).
    $\langle\eta_{s},|k|^{n}\eta_{s}\rangle\leq C\|\eta_{s}\|^{2}$,
  3. (iii).
    $\|c^{\ast}(|k|^{n}\eta_{s})\psi\|\leq C\|\eta_{s}\|\ \|(\mathcal{N}+1)^{1/2}\psi\|$.
We will now give the proof of the second main theorem.
Proof of Theorem 3.6.
We use the approach as sketched in Remark 3.9. Since the bosonic property holds only with an error, the equality (3.21) holds only approximately:
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|$ $=\|\psi-e^{i\mathbb{H}^{\text{eff}}t}e^{-iE_{N}^{\text{pw}}t}e^{i2\text{Im}(\nu_{t})}e^{-i\text{Im}\int_{0}^{t}\textnormal{d}s\langle\dot{\eta}_{s},\eta_{s}\rangle}W(\eta_{t})\psi\|$
$\leq\int_{0}^{t}\textnormal{d}s\|\big{(}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\|$
$\leq\int_{0}^{t}\textnormal{d}s\|h_{0}W(\eta_{s})\phi\otimes\Omega\|+\|\text{Error}_{1}\|+\|\text{Error}_{2}\|.$ (5.27)
We will first estimate the error terms and then subsequently treat the $h_{0}$ term in a separate lemma. We give an explicit expression for the first error term using Lemma 5.1 on the approximate shift property applied to $c_{\alpha}(k)$ in the $c^{\ast}c(\epsilon),c(h_{y})$ and $c(i\dot{\eta}_{s})$ terms. Thus, the error of (3.21) is given by
$\text{Error}_{1}\coloneqq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\left(\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right)W(\eta_{s})\mathcal{R}^{1}_{\alpha}(k)\psi.$ (5.28)
The second error term is given by Lemma 5.5 on the time derivative of the almost-bosonic Weyl operator and therefore
$\text{Error}_{2}\coloneqq-$ $2i\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\psi$
$\equiv$ $2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau e^{(1-\tau)B}\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}e^{is\epsilon_{\alpha}(k)}\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)$
$-\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{-is\epsilon_{\alpha}(k)}(h_{y})_{\alpha}(k)\bigg{)}e^{\tau B}\psi.$ (5.29)
Firstly, we show that the term $\mathcal{R}^{\lambda}_{\alpha}(k)\psi=\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi$ as defined in Lemma 5.1 constitutes indeed a small error. We estimate
$\sum_{l\in\Gamma}\sum_{\gamma\in\mathcal{I}_{l}}\|\mathcal{R}^{1}_{\gamma}(l)\psi\|^{2}$
$\leq\sum_{l\in\Gamma}\sum_{\gamma\in\mathcal{I}_{l}\cap\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma}|\eta_{\gamma}(k)|\int_{0}^{1}\textnormal{d}\tau\|\mathcal{E}_{\gamma}(k,l)e^{\tau B}\psi\|\bigg{)}^{2}$
$\leq C\left(MN^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)\psi\|\right)^{2}$ (5.30)
where we used $\sum_{\alpha\in\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma}|\eta_{\alpha}(k)|\bigg{)}^{2}\leq\sum_{k^{\prime},k\in\Gamma}\|\eta(k)\|_{l^{2}}\|\eta(k^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.6 and
$\int_{0}^{1}\textnormal{d}\tau\|\mathcal{E}_{\gamma}(k,l)e^{\tau B}\psi\|$ (5.31)
$\leq\int_{0}^{1}\textnormal{d}\tau\langle e^{\tau B}\psi,|\mathcal{E}_{\gamma}(k,l)|^{2}e^{\tau B}\psi\rangle^{1/2}\leq CMN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\langle e^{\tau B}\psi,\mathcal{N}^{2}e^{\tau B}\psi\rangle^{1/2}$ (5.32)
$\leq CMN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau e^{C\|\eta_{s}\|\tau}\|(\mathcal{N}+3)\psi\|\leq C\|\eta_{s}\|^{-1}(e^{C\|\eta_{s}\|}-1)MN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)\psi\|$ (5.33)
which follows from $e^{\tau B}$ is unitary in the first inequality, Lemma A.3 in the second inequality and Proposition 5.4 in the third inequality.
Secondly, we estimate
$\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\leq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\|c_{\alpha}^{\ast}(k)e^{(1-\tau)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|c_{\alpha}^{\ast}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
$\quad+\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|[c_{\alpha}^{\ast}(k),e^{(1-\tau)B}]\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\bigg{(}\|(\mathcal{N}+1)^{\frac{3}{2}}e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}+C\|\eta_{s}\|^{2}MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ (1-\tau)\|\mathcal{N}e^{\tau B}\psi\|$
$\quad+\sum_{k\in\Gamma}\|\eta_{s}\|\int_{0}^{1}\textnormal{d}\tau\bigg{(}\sum_{l\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|\mathcal{R}^{1}_{\alpha}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}$
$\leq CMN^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{\frac{3}{2}}\psi\|+CMN^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-C\|\eta_{s}\|-1)\|(\mathcal{N}+3)\psi\|$
$\quad+CM^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)^{2}\|(\mathcal{N}+3)^{2}\psi\|$
$\leq CM^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{2}\psi\|.$ (5.34)
where we used $[c_{\alpha}^{\ast}(k),e^{\sigma B}]=-\lambda\eta_{\alpha}(k)e^{\sigma B}-e^{\sigma B}\mathcal{R}^{1}_{\alpha}(k)$ from Lemma 5.1, the Cauchy-Schwarz inequality for the $\alpha$-summation, Lemma A.3 in the third inequality and in the fourth inequality we used Proposition 5.4 and (5.30).
In total by combining (5.30) and (5.34) we end up with the following estimate
$\|\text{Error}_{1}\|$ $\leq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\left\{\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right\}e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\leq Ck_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\quad+C\lambda\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\|n_{\alpha}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\leq Ck_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\quad+C\lambda k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\|\hat{V}\|_{2}\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1}_{\alpha}(k)\psi\|^{2}\bigg{)}^{1/2}$
$\leq Ck_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\ (e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{2}\psi\|$
$\quad+C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\ (e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)\psi\|$
$\leq C\ (f_{C}(\lambda k_{\textnormal{F}}t)-1)(\lambda+1)k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)^{2}\psi\|.$ (5.35)
where we used (A.2) and $e^{B}$ unitary in the third inequality and Lemma 5.6 in the last line. Using $\psi=\phi\otimes\Omega$ we obtain the desired bound.
Similarly, we obtain an estimate for the second error term using Cauchy-Schwarz, (5.30) and Proposition 5.4
$\|\text{Error}_{2}\|$
$\leq 2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\bigg{(}\|\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)e^{\tau B}\psi\|\bigg{)}$
$\leq 2\lambda\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|\hat{V}(k)n_{\alpha}(k)|\bigg{(}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|\bigg{)}$
$\leq C\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ \lambda k_{\textnormal{F}}\|\hat{V}\|_{2}\bigg{(}\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}+\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}\bigg{)}$
$\leq C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|(1-\tau)}-1)\|(\mathcal{N}+3)e^{\tau B}\psi\|$
$\leq C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+5)\psi\|$
$\leq C(f_{C}(\lambda k_{\textnormal{F}}t)-1)\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+5)\psi\|.$ (5.36)
Again, by using $\psi=\phi\otimes\Omega$ we obtain the desired bound.
With the subsequent Lemma 5.9, we can conclude with a bound on $h_{0}=-\beta\Delta_{y}$ of the form
$\int_{0}^{t}\textnormal{d}s\|h_{0}W(\eta_{s})\phi\otimes\Omega\|\leq C\beta\int_{0}^{t}\textnormal{d}s\left\{(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)\right\}$
$\leq C\beta t\left\{\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]^{4}\right\}\left\{f_{C}(\lambda k_{\textnormal{F}}t)+1\right\}$ (5.37)
[…]
Lemma 5.9.
Under the assumptions of Theorem 3.6, it holds that for all $t\geq 0$
$\|\Delta_{y}W(\eta_{t})\phi\otimes\Omega\|\leq C(\|\eta_{t}\|+\|\eta_{t}\|^{4})(e^{C\|\eta_{t}\|}+1)$
for $C>0$ independent of $k_{\textnormal{F}}$.
Proof of Lemma 5.9.
We explicitly calculate the action of the Laplacian on the coupled coherent state $W(\eta_{s})\psi=e^{B}\psi$ i.e.
$-\Delta_{y}W(\eta_{s})\psi=-\big{(}\Delta_{y}W(\eta_{s})\big{)}\psi-2\beta\nabla_{y}W(\eta_{s})\cdot\nabla_{y}\psi-W(\eta_{s})\Delta_{y}\psi.$ (5.38)
In total we expect, that all terms can be bounded by assumption on the initial condition on $\phi$. We will first focus on the term $\Delta_{y}W(\eta_{s})$. Recall that it holds
$W(\eta_{s})^{\ast}\partial_{y_{i}}W(\eta_{s})=\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(\partial_{y_{i}}B_{s}\right)e^{\tau B}$ (5.39)
due to the same calculation as in the proof of Lemma 5.5. With
$\partial_{y_{i}}B_{s}$ $=\partial_{y_{i}}\left\{c^{\ast}(\eta_{s})-c(\eta_{s})\right\}=c^{\ast}(\partial_{y_{i}}\eta_{s})-c(\partial_{y_{i}}\eta_{s}),$ (5.40)
$\partial_{y_{i}}\eta_{s}$ $=\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}\lambda\hat{V}(k)n_{\alpha}(k)ik_{i}e^{iky}=ik_{i}\eta_{s}$ (5.41)
it follows analogously to Lemma 5.5 that
$\partial_{y_{i}}W(\eta_{s})$ $=\int_{0}^{1}\textnormal{d}\tau\ e^{(1-\tau)B_{s}}\left(\partial_{y_{i}}B_{s}\right)e^{\tau B_{s}}$
$=\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})$
$\hphantom{=}+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}).$ (5.42)
And repeating the differentiation with
$\partial_{y_{i}}W_{\tau}(\eta_{s})$ $=\tau\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\tau\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W_{\tau}(\eta_{s})$
$\hphantom{=}+2i\int_{0}^{\tau}\textnormal{d}\sigma\ W_{1-\sigma}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\tau-\sigma}\rangle_{\Gamma}W_{\sigma}(\eta_{s})$ (5.43)
yields
$\Delta W(\eta_{s})=\sum_{i=1}^{3}\partial_{y_{i}}^{2}W(\eta_{s})$
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})$
$+\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\partial_{y_{i}}W(\eta_{s})$
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}$ (5.44)
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})$
$+\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\times$
×$\times\left\{\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\right\}$
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}$ (5.45)
$\eqqcolon$ $I_{1}+I_{2}+I_{3}+I_{4}$
with
$I_{1}$ $\coloneqq\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})$
$\hphantom{\coloneqq}-\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s}),$ (5.46)
$I_{2}$ $\coloneqq 2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}),$ (5.47)
$I_{3}$ $\coloneqq 2\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\int_{0}^{1}\textnormal{d}\tau\ iW_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}),$ (5.48)
$I_{4}$ $\coloneqq 2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}.$ (5.49)
We show that each term can be bounded here by a constant at most of order $1$.
For $I_{1}$, we can treat all $c^{\ast}(\cdots)$ and $c(\cdots)$ terms with Lemma 5.6 and Lemma 5.8. Furthermore we use that $c_{\alpha}(k)\mathcal{N}=(\mathcal{N}+2)c_{\alpha}(k)$ to estimate
$\|I_{1}\psi\|$ $\leq C\|\eta_{s}\|\ \|(\mathcal{N}+1)^{1/2}W(\eta_{s})\psi\|+C(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})\|W(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|(\mathcal{N}+3)W(\eta_{s})\psi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}\|(\mathcal{N}+5)\psi\|+C(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})$ (5.50)
where we used Proposition 5.4.
For $I_{2}$, we use a similar approach to (5.30) to obtain
$\|I_{2}\psi\|$ $\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|(1-\tau)}-1)\|(\mathcal{N}+3)W_{\tau}(\eta_{s})\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ (e^{C\|\eta_{s}\|}e^{(\tilde{C}-C)\|\eta_{s}\|\tau}-e^{\tilde{C}\|\eta_{s}\|\tau})\|(\mathcal{N}+5)\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)\psi\|.$ (5.51)
For $I_{3}$, we first observe that for $n\in\mathbb{N}$ and $\mu\in[0,1]$
$\|\mathcal{N}^{n}\mathcal{R}^{\mu}_{\beta}(k)\psi\|$ $\leq\int_{0}^{\mu}\textnormal{d}\tau\ \|\mathcal{N}^{n}e^{-\tau B}\left(\sum_{l\in\Gamma}\eta_{\beta}(l)\mathcal{E}_{\beta}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma}|\eta_{\beta}(l)|\ \|\mathcal{E}_{\beta}(l,k)(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma}|\eta_{\beta}(l)|\ MN^{-\frac{2}{3}+\delta}\|\mathcal{N}(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\|\eta_{s}\|^{-1}(e^{C\|\eta_{s}\|\mu}-1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\beta}(l)|\ \|(\mathcal{N}+5)^{n+1}\psi\|.$ (5.52)
We use a similar approach to (5.34) and insert the above inequality (5.52) to obtain
$\|I_{3}\psi\|$ $\leq C\|\eta_{s}\|\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\left(\|\eta_{s}\|(\mathcal{N}+1)^{1/2}+C\|\eta_{s}\|^{2}\right)W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\mathcal{N}W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ e^{C\|\eta_{s}\|(1-\tau)}\|\mathcal{N}\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|^{2}MN^{-\frac{2}{3}+\delta}\int_{0}^{1}\textnormal{d}\tau\ e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|(1-\tau)}-1)\|(\mathcal{N}+5)^{2}\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)^{2}\psi\|$
$\quad+CMN^{-\frac{2}{3}+\delta}\|\eta_{s}\|^{2}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}(\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+5)^{2}\psi\|.$ (5.53)
For $I_{4}$, we first calculate
$\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}$
$=\int_{\tau}^{1}\textnormal{d}\sigma\ \partial_{y_{i}}e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}$
$\hphantom{=}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\partial_{y_{i}}e^{\sigma B}$ (5.54)
$\eqqcolon I_{4,1}+I_{4,2}+I_{4,3}+I_{4,4}+I_{4,5}$
with
$I_{4,1}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i(1-\sigma)\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)$
[…]
$I_{4,2}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\sigma-a}\rangle_{\Gamma}e^{aB}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.56)
$I_{4,3}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.57)
$I_{4,4}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\sigma\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B},$ (5.58)
$I_{4,5}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{\Gamma}e^{aB}.$ (5.59)
We approach each term similarly to (5.30).
For $I_{4,1}$, it holds
$\|I_{4,1}\psi\|$
$\leq C\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\left(\|\eta_{s}\|\ \|(\mathcal{N}+1)e^{(1-\sigma)B}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|^{2}\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\left(e^{C\|\eta_{s}\|(1-\sigma)}\|(\mathcal{N}+3)\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|\,\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)\|(\mathcal{N}+3)^{2}e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+5)^{2}\psi\|$ (5.60)
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\left((1-\tau)e^{C\|\eta_{s}\|(1-\tau)}+C\|\eta_{s}\|e^{C\|\eta_{s}\|}\right)\|(\mathcal{N}+5)^{2}\psi\|$ (5.61)
where we used Lemma A.4, Lemma 5.6 and Lemma 5.8 in the first inequality, Lemma A.3 in the third inequality and Proposition 5.4 in the second and forth inequality. Therefore using $\int_{0}^{1}(1-\tau)e^{y(1-\tau)}\textnormal{d}\tau=y^{-2}\left((y-1)e^{y}+1\right)$ yields
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,1}\psi\|^{2}}$ (5.62)
$\leq CMN^{-\frac{2}{3}+\delta}\left((\|\eta_{s}\|^{3}+\|\eta_{s}\|-1)e^{C\|\eta_{s}\|}+1\right)\|(\mathcal{N}+5)^{2}\psi\|$ (5.63)
where we used $\sum_{\alpha\in\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\bigg{)}^{2}\leq\sum_{l^{\prime},l\in\Gamma}\|\eta(l)\|_{l^{2}}\|\eta(l^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.6 and Lemma 5.8.
For $I_{4,2}$, it holds
$\|I_{4,2}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ \|(\mathcal{N}+3)e^{aB}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|a}\|(\mathcal{N}+5)^{2}e^{\sigma B}\psi\|$
$\leq C\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma(e^{C\|\eta_{s}\|(1-\sigma)}-1)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+8)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}\left(MN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|e^{C\|\eta_{s}\|(1-\tau)}\|(\mathcal{N}+8)^{2}\psi\|$ (5.64)
where we used Cauchy-Schwarz with (5.30), Lemma 5.6 and Lemma 5.8 in the first inequality, Proposition 5.4 and Lemma A.3 in the second inequality. Therefore it follows
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,2}\psi\|^{2}}\leq CM^{2}N^{-\frac{4}{3}+2\delta}(e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+8)^{2}\psi\|.$
The term $I_{4,3}$ is estimated similarly to (5.30) by
$\|I_{4,3}\psi\|\leq C\|\eta_{s}\|^{-1}MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|(e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+3)\psi\|$ (5.65)
and therefore
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,3}\psi\|^{2}}\leq CMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)\psi\|.$ (5.66)
Similarly for $I_{4,4}$, we estimate
$\|I_{4,4}\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|\mathcal{N}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{2})MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma\|(\mathcal{N}+1)^{2}e^{\sigma B}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\ \left(e^{C\|\eta_{s}\|}+(1-C\|\eta_{s}\|\tau)e^{C\|\eta_{s}\|\tau}\right)\|(\mathcal{N}+3)^{2}\psi\|$
$\leq C(\|\eta_{s}\|^{-1}+1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\ e^{C\|\eta_{s}\|}\|(\mathcal{N}+3)^{2}\psi\|$ (5.67)
using $\int_{\tau}^{1}\textnormal{d}\sigma\ \sigma e^{y\sigma}=y^{-2}\left((y-1)e^{y}-(\tau y-1)e^{y\tau}\right)$ and therefore
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,4}\psi\|^{2}}$
$\leq CMN^{-\frac{2}{3}+\delta}(\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}\|(\mathcal{N}+1)^{2}\psi\|.$ (5.68)
For $I_{4,5}$, we estimate
$\|I_{4,5}\psi\|$ $=\|2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{\Gamma}e^{aB}\psi\|$
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ \|\mathcal{N}e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{\Gamma}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}\bigg{(}\sum_{\alpha\in\mathcal{I}_{k}}\|(\mathcal{N}+3)\mathcal{R}^{\sigma-a}_{\alpha}(k)e^{aB}\psi\|^{2}\bigg{)}^{1/2}$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|(1-a)}(e^{C\|\eta_{s}\|(\sigma-a)}-1)\|(\mathcal{N}+8)^{2}e^{aB}\psi\|$
$\leq C\|\eta_{s}\|M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{\sigma}\textnormal{d}a\ e^{C\|\eta_{s}\|}e^{C^{\prime}\|\eta_{s}\|\sigma}e^{-C^{\prime\prime}\|\eta_{s}\|a}\|(\mathcal{N}+10)^{2}\psi\|$
$\leq C\|\eta_{s}\|^{-1}M^{2}N^{-\frac{4}{3}+2\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\ (e^{C\|\eta_{s}\|}-e^{C\|\eta_{s}\|\tau})\|(\mathcal{N}+10)^{2}\psi\|$ (5.69)
where we used (5.52) in the third inequality. Therefore we obtain
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,5}\psi\|^{2}}\leq CM^{2}N^{-\frac{4}{3}+2\delta}\|(\mathcal{N}+10)^{2}\psi\|.$ (5.70)
Taking the five bounds together we finally obtain
$\|I_{4}\psi\|$ $=\|2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}\psi\|$
$\leq C\int_{0}^{1}\textnormal{d}\tau\sum_{n=1}^{5}\|\eta_{s}\|\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,n}\psi\|^{2}\bigg{)}^{1/2}$
$\leq CMN^{-\frac{2}{3}+\delta}\left((\|\eta_{s}\|^{3}+\|\eta_{s}\|-1)e^{C\|\eta_{s}\|}+1\right)\|(\mathcal{N}+8)^{2}\psi\|.$ (5.71)
Combining (5.50), (5.51), (5.53) and (5.71) with $\psi\equiv\phi\otimes\Omega$ yields the desired final result of
$\|\big{(}\Delta_{y}W(\eta_{s})\big{)}\psi\|\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1).$ (5.72)
Similarly to (5.50) and (5.51), we further estimate the second term of (5.38)
$\|\nabla_{y}W(\eta_{s})\cdot\nabla_{y}\psi\|$
$=\|\sum_{i=1}^{3}\Big{(}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})$
$+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\Big{)}\partial_{y_{i}}\psi\|$
$\leq C\left((\|\eta_{s}\|+\|\eta_{s}\|^{2})e^{C\|\eta_{s}\|}+(\|\eta_{s}\|^{2}+\|\eta_{s}\|^{4})\right)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|$
$+CMN^{-\frac{2}{3}+\delta}e^{C\|\eta_{s}\|}(e^{C\|\eta_{s}\|}-1)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|$ (5.73)
$\leq C(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)\sum_{i=1}^{3}\|\partial_{y_{i}}\phi\|.$ (5.74)
Therefore in total for all $t\in\mathbb{R}$
$\|\Delta_{y}W(\eta_{t})\phi\otimes\Omega\|$ (5.75)
$\leq C\left\{(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)+\sum_{i=1}^{3}\big{(}(\|\eta_{s}\|+\|\eta_{s}\|^{4})(e^{C\|\eta_{s}\|}+1)\|\partial_{y_{i}}\phi\|+\|\partial_{y_{i}}^{2}\phi\|\big{)}\right\}.$ (5.76)
Proofs of properties of $\eta$
Proof of Lemma 5.6.
Recall that by definition it holds
$\|\eta_{s}\|^{2}=\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\eta_{s}\big{)}_{\alpha}(k)|^{2}=\lambda^{2}\sum_{k\in\Gamma}|\hat{V}(k)|^{2}\underbrace{\sum_{\alpha\in\mathcal{I}_{k}}\bigg{|}n_{\alpha}(k)\frac{\sin\big{(}\epsilon_{\alpha}(k)s/2\big{)}}{\epsilon_{\alpha}(k)/2}\bigg{|}^{2}}_{\eqqcolon S_{k}}$ (5.77)
with $\epsilon_{\alpha}(k)=2k_{F}|k\cdot\omega_{\alpha}|$. We will first approximate $S_{k}$ and then give a upper bound.
Firstly, we approximate the $\alpha$-sum $S_{k}$ by an integral by identifying
$n_{\alpha}(k)^{2}=k_{\textnormal{F}}^{2}|k|\sigma(p_{\alpha})u_{\alpha}(k)^{2}\left(1+\mathcal{O}(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta})\right)$ (5.78)
with $\cos\theta_{\alpha}\coloneqq|\hat{k}\cdot\hat{\omega}_{\alpha}|\equiv u_{\alpha}(k)^{2}$ analogously to Lemma A.1. Thus, we calculate the half-sphere integral as approximation for $S_{k}$
$S_{k}$ $=\frac{1}{|k|}\int_{0}^{2\pi}\textnormal{d}\varphi\int_{0}^{\pi/2}\textnormal{d}\theta\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\sin\theta+\mathcal{E}$
$=\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(k_{\textnormal{F}}|k|su)^{2}}{u}+\mathcal{E}$
$=\frac{\pi}{|k|}\left\{\log(2k_{\textnormal{F}}|k|s)-\text{Ci}(2k_{\textnormal{F}}|k|s)+\gamma\right\}+\mathcal{E}$ (5.79)
with total error $\mathcal{E}=\mathcal{E}_{1}+\mathcal{E}_{2}+\mathcal{E}_{3}$ consisting three terms: the error $\mathcal{E}_{1}$ from (5.78), the error $\mathcal{E}_{2}$ from approximating the discrete variables $\theta_{\alpha}$ by continuous $\theta$ given by
$\mathcal{E}_{2}$ $=\left|\int_{p_{\alpha}}\textnormal{d}\sigma\ \frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}-\sigma(p_{\alpha})\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta_{\alpha})^{2}}{\cos\theta_{\alpha}}\right|$
$\leq\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\textnormal{d}}{\textnormal{d}\theta}\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\sup_{(\theta,\varphi)\in p_{\alpha}}\left|\theta-\theta_{\alpha}\right|\sigma(p_{\alpha})$
$\leq C(k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}$ (5.80)
and the error $\mathcal{E}_{3}$ from the patch construction which is given by
$\mathcal{E}_{3}=\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\sin(k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\left|\sigma\Big{(}\bigcup_{\alpha\in\mathcal{I}_{k}}p_{\alpha}\Big{)}-\sigma(S)\right|\leq Ck_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right).$ (5.81)
Thus with the triangle inequality
$\left|S_{k}-\frac{\pi}{|k|}\left\{\log(2k_{\textnormal{F}}|k|s)-\text{Ci}(2k_{\textnormal{F}}|k|s)+\gamma\right\}\right|$
$\leq\frac{C}{|k|}\left\{M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+(k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}+k_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)\right\}.$ (5.82)
We will now give an estimate for $S_{k}$ by approximating for any $m\in\mathbb{R}$
$\frac{\sin(mx)^{2}}{x}=\frac{1}{2}\frac{1-\cos(2mx)}{x}\leq\begin{cases}m^{2}x&\text{if }2mx\in[0,\pi/2],\\ \frac{1}{x}&\text{if }2mx>\pi/2\end{cases}$
where we used $1-x^{2}/2\leq\cos(x)$ for all $x\in[0,\pi/2]$. Therefore
$\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(k_{\textnormal{F}}|k|su)^{2}}{u}$
$\leq\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(k_{\textnormal{F}}|k|s)^{2}u\ \chi(2k_{\textnormal{F}}|k|su\leq\pi/2)+\frac{1}{u}\chi(2k_{\textnormal{F}}|k|su>\pi/2)\right\}$
$\leq\frac{2\pi}{|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(k_{\textnormal{F}}|k|s)^{2}u\ \chi\Big{(}u\leq\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}+\frac{1}{u}\ \chi\Big{(}u>\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}\right\}$
$=\frac{2\pi}{|k|}\left\{\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2}\Big{(}\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}^{2}-\log\Big{(}\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}\right\}\ \chi\Big{(}1>\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}$
$\qquad+\frac{2\pi}{|k|}\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}1\leq\frac{\pi}{4k_{\textnormal{F}}|k|s}\Big{)}$
$=\frac{2\pi}{|k|}\left\{\frac{\pi^{2}}{32}-\log(\pi)+\log(4k_{\textnormal{F}}|k|s)\right\}\ \chi\Big{(}k_{\textnormal{F}}|k|s>\frac{\pi}{4}\Big{)}$
$\qquad+\frac{2\pi}{|k|}\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}k_{\textnormal{F}}|k|s\leq\frac{\pi}{4}\Big{)}$
$\leq\frac{2\pi}{|k|}\min\left\{\frac{1}{2}(k_{\textnormal{F}}|k|s)^{2},\log(4k_{\textnormal{F}}|k|s+1)\right\}.$ (5.83)
Note that $\log(4k_{\textnormal{F}}|k|s+1)\leq\log(4k_{\textnormal{F}}s+1)+\log|k|$ using $\frac{1}{2}x^{2}\leq\log(4x+1)$ for all $x\leq 2$ and $|k|\geq 1$ for all $k\in\mathbb{Z}_{\ast}^{3}$. Thus, it holds
$\|\eta_{s}\|^{2}$ $\leq 2\pi\lambda^{2}\sum_{k\in\mathbb{Z}_{\ast}^{3}}|\hat{V}(k)|^{2}\min\left\{\frac{(k_{\textnormal{F}}s)^{2}}{2}|k|,\frac{\log(4k_{\textnormal{F}}|k|s+1)}{|k|}\right\}$
$\leq 2\pi\lambda^{2}\min\left\{\|\hat{V}(\cdot)^{1/2}\|_{2}^{2}(k_{\textnormal{F}}s)^{2}/2,\|\hat{V}\|_{2}^{2}\left(\log(4k_{\textnormal{F}}s+1)+1\right)\right\}.$ (5.84)
Thus by using $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ and $\sqrt{\log(x+1)}\leq\log(x+2)$ we obtain the desired
$\|\eta_{s}\|$ $\leq\lambda\min\left\{bk_{\textnormal{F}}s,\log(4k_{\textnormal{F}}s+2)^{a}+a\right\},$ (5.85)
$e^{c_{0}\|\eta_{s}\|}$ $\leq\min\left\{e^{bc_{0}\lambda k_{\textnormal{F}}s},e^{ac_{0}\lambda}(4k_{\textnormal{F}}s+2)^{ac_{0}\lambda}\right\}$ (5.86)
with $a=\sqrt{2\pi}\|\hat{V}\|_{2},b=\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}.$
Proof of Lemma 5.8.
For the inequalities, we observe that for $n\in\mathbb{N}_{0}$ it holds
$\sum_{k\in\Gamma}\||k|^{n}\eta_{s}(k)\|_{l^{2}}$ $\equiv\sum_{k\in\Gamma}|k|^{n}\Big{(}\sum_{\alpha\in\mathcal{I}_{k}}|(\eta_{s})_{\alpha}(k)|^{2}\Big{)}^{1/2}=\lambda\sum_{k\in\Gamma}|k|^{n}|\hat{V}(k)|\left(S_{k}\right)^{1/2}$ (5.87)
with $S_{k}$ from (5.77) Since by assumption $\|(\cdot)^{n}\hat{V}\|_{1}$ is bounded for each $n\in\mathbb{N}$, the first statement follows. The second statement simply follows from the same calculation and recalling that $\langle\eta_{s},|k|^{n}\eta_{s}\rangle\equiv\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}|\eta_{\alpha}(k)|^{2}$. The third statement follows from Lemma A.4 and
$\|c^{\ast}(|k|^{n}\eta_{s})\psi\|\leq\sum_{k\in\Gamma}\|\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}\eta_{\alpha}(k)c_{\alpha}^{\ast}(k)\psi\|\leq\sum_{k\in\Gamma}\||k|^{n}\eta_{s}(k)\|_{l^{2}}\|(\mathcal{N}+1)^{1/2}\psi\|.$ (5.88)
6 Proof of Corollary 3.11
Proof of Corollary 3.11.
We observe that by the inverse triangle inequality it holds
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|$
$\geq\|e^{iP(t)}W(\eta_{t})\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|$ (6.1a)
$\hphantom{\geq}-\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|-\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\text{eff}}t}\psi\|.$ (6.1b)
The second line (6.1b) is to be bounded from below by bounds of order $o(1)$ from Theorem 3.1 and Theorem 3.6. Therefore it is sufficient to show that the first line (6.1a) is large.
The first line (6.1a) can be explicitly estimated by the same approach as in the proof of Theorem 3.6. That is use Duhamel’s formula, commute all $c_{\alpha}(k)$-operators with $W(\eta_{s})$ and collect the error terms via Lemma 5.1:
$\bigg{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})-1\bigg{)}\psi$
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\left(\big{(}\widetilde{\mathbb{H}^{\text{eff}}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\eta_{s},\dot{\eta}_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\right)$ (6.2)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{(}\big{(}h_{0}-c^{\ast}(h_{y})-c(h_{y})\big{)}W(\eta_{s})\psi+\text{Error}\bigg{)}$ (6.3)
$=i\int_{0}^{t}\textnormal{d}s\ e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}\bigg{(}-W(\eta_{s})\left(c^{\ast}(h_{y})+\langle\eta_{s},h_{y}\rangle+\langle h_{y},\eta_{s}\rangle\right)\psi+h_{0}W(\eta_{s})\psi+\widetilde{\text{Error}}\bigg{)}$ (6.4)
with $\int_{0}^{t}\textnormal{d}s\ \text{Error}=\text{Error}_{1}+\text{Error}_{2}$ from (5.27) and where we also commuted $c^{\ast}(h_{y})$ and $c(h_{y})$ with $W(\eta_{s})$ such that the new error term is of the form
$\widetilde{\text{Error}}$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\left(\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)-e^{is\epsilon_{\alpha}(k)}\right)\overline{(h_{y})_{\alpha}(k)}W(\eta_{s})\mathcal{R}^{1}_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}W(\eta_{s})\mathcal{R}^{1}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\psi.$ (6.5)
Note that $\widetilde{\text{Error}}$ is $s$-dependent even though we did not include the dependence explicitly in the notation.
We employ the Cauchy-Schwarz inequality, insert our previous finding with the triangle inequality and use that $e^{-i\widetilde{\mathbb{H}^{\text{eff}}}s}$ is unitary to estimate
$\|e^{iP(t)}W(\eta_{t})\psi-e^{-i\widetilde{\mathbb{H}^{\text{eff}}}t}\psi\|$
$=\|\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})-1\big{)}\psi\|\geq\left|\langle\psi,\big{(}1-e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})\big{)}\psi\rangle\right|$
$=\left|\int_{0}^{t}\textnormal{d}s\langle e^{-iP(s)}e^{-i\widetilde{\mathbb{H}^{\text{eff}}}s}\psi,\big{(}\widetilde{\mathbb{H}^{\text{eff}}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\eta_{s},\dot{\eta}_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\rangle\right|$
$\geq\left|\int_{0}^{t}\textnormal{d}s\langle\psi,e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})\left(c^{\ast}(h_{y})+2\text{Re}\langle h_{y},\eta_{s}\rangle\right)\psi\rangle\right|$
$\hskip 113.81102pt-\int_{0}^{t}\textnormal{d}s\left|\langle e^{-iP(s)}e^{-i\widetilde{\mathbb{H}^{\text{eff}}}s}\psi,\widetilde{\text{Error}}+h_{0}W(\eta_{s})\psi\rangle\right|$ (6.6)
$\geq\left|\int_{0}^{t}\textnormal{d}s\langle\psi,\big{(}1+e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})-1\big{)}\left(c^{\ast}(h_{y})+2\text{Re}\langle h_{y},\eta_{s}\rangle\right)\psi\rangle\right|$
$\hskip 113.81102pt-\int_{0}^{t}\textnormal{d}s\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$ (6.7)
$\geq\left|\int_{0}^{t}\textnormal{d}s\left\{\langle\psi,c^{\ast}(h_{y})\psi\rangle+2\text{Re}\langle h_{y},\eta_{s}\rangle\right\}\right|$
$\hphantom{\geq}\qquad-\left|\int_{0}^{t}\textnormal{d}s\langle\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})-1\big{)}^{\ast}\psi,\left(c^{\ast}(h_{y})+2\text{Re}\langle h_{y},\eta_{s}\rangle\right)\psi\rangle\right|$
$\hphantom{\geq}\qquad\qquad-\int_{0}^{t}\textnormal{d}s\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$ (6.8)
$\geq 2\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\langle h_{y},\eta_{s}\rangle\right|-\big{(}\|c^{\ast}(h_{y})\psi\|+2\sup_{s\in[0,t]}|\text{Re}\langle h_{y},\eta_{s}\rangle|\big{)}\int_{0}^{t}\textnormal{d}s\|\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}s}e^{iP(s)}W(\eta_{s})-1\big{)}\psi\|$
$\hphantom{\geq}\qquad\qquad-t\sup_{s\in[0,t]}\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$ (6.9)
where we used $\langle\psi,c^{\ast}(h_{y})\psi\rangle=0$ since $c(h_{y})\phi\otimes\Omega=0$.
The three parts of the above inequality (6.9) can be bounded in the following:
Firstly, the error term can be bounded with (5.37) and analogously to (5.35) and (5.36):
$t\sup_{s\in[0,t]}\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$
$\leq C\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}(f_{C}(\lambda k_{\textnormal{F}}t)-1)+C\beta\left(\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]^{4}\right)\left(f_{C}(\lambda k_{\textnormal{F}}t)+1\right)$
$\leq CQ_{V}(k_{\textnormal{F}}t)\max\{\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta},\beta t\}$
$\leq C\max\{MN^{-\frac{2}{3}+\delta},\beta k_{\textnormal{F}}^{-1}\lambda^{-1}\}\eqqcolon d$ (6.10)
where we used $0\leq t\lesssim\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$ and defined the error variable $d$.
Secondly, the prefactor of the integral term is bounded by Lemma A.4
$\|c^{\ast}(h_{y})\psi\|+2\sup_{s\in[0,t]}|\text{Re}\langle h_{y},\eta_{s}\rangle|$
$\leq\|h_{y}\|\|(\mathcal{N}+1)^{1/2}\psi\|+2\sup_{s\in[0,t]}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\frac{1-\cos(s\epsilon_{\alpha}(k))}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}$ (6.11)
$\leq\|h_{y}\|+\frac{2}{k_{\textnormal{F}}}\|h_{y}\|^{2}\leq C\lambda k_{\textnormal{F}}\big{(}\|\hat{V}\|+\lambda\|\hat{V}\|^{2}\big{)}\eqqcolon\theta$ (6.12)
where we introduced the variable $\theta$.
Thirdly, we want to show that the remaining term $|\int_{0}^{t}\textnormal{d}s\,\text{Re}\langle h_{y},\eta_{s}\rangle|$ is large:
$\left|\int_{0}^{t}\textnormal{d}s\,\text{Re}\langle h_{y},\eta_{s}\rangle\right|$ $=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}\right|$
$=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\lambda^{2}\hat{V}(k)^{2}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}\right|$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\hat{V}(k)^{2}\frac{e^{-i2k_{\textnormal{F}}|k|s}-1}{2k_{\textnormal{F}}|k|s}\right|\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$ (6.13)
where we used that the sum over $\alpha$ is approximated similarly to (A.2) by integrating over the half sphere
$\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}$ $=2\pi\int_{0}^{\pi/2}\textnormal{d}\theta\ \frac{e^{-is2k_{\textnormal{F}}|k|\cos\theta}-1}{2k_{\textnormal{F}}|k|\cos\theta}\cos\theta\sin\theta\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$=i\frac{2\pi k_{\textnormal{F}}^{2}}{2k_{\textnormal{F}}}\frac{e^{-i2k_{\textnormal{F}}|k|s}-1}{2k_{\textnormal{F}}|k|s}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$ (6.14)
Therefore we obtain
$\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\hat{V}(k)^{2}\frac{e^{-i2k_{\textnormal{F}}|k|s}-1}{2k_{\textnormal{F}}|k|s}\right|=\left|\sum_{k\in\Gamma}\hat{V}(k)^{2}\int_{0}^{t}\textnormal{d}s\frac{1-\cos(2k_{\textnormal{F}}|k|s)}{2k_{\textnormal{F}}|k|s}\right|.$ (6.15)
It is well-known that for all $x>0$ it holds $\cos(x)<1-4x^{2}/\pi^{2}$ and thus $1-\cos(x)\geq 4x^{2}/\pi^{2}$ for all $x\in(0,\pi/2)$ Also it holds for all $x>\pi/2$ that
$\int_{\frac{\pi}{2}}^{x}\frac{1-\cos y}{y}\textnormal{d}y=\ln(x)-\ln(\frac{\pi}{2})+\text{Ci}(\frac{\pi}{2})-\text{Ci}(x)\geq\ln(x)-\ln(\frac{\pi}{2})$
since $\text{Ci}(\pi/2)>0$ and $\text{Ci}(\pi/2)\geq\text{Ci}(x)$ for all $x\geq\pi/2$. Thus we obtain
$\int_{0}^{x}\frac{1-\cos y}{y}\textnormal{d}y$ $\geq\chi(x>\frac{\pi}{2})\left\{\int_{0}^{\frac{\pi}{2}}\frac{4y}{\pi^{2}}\textnormal{d}y+\int_{\frac{\pi}{2}}^{x}\frac{1-\cos y}{y}\textnormal{d}y\right\}+\chi(x\leq\frac{\pi}{2})\int_{0}^{x}\frac{4y}{\pi^{2}}\textnormal{d}y$ (6.16)
$=\chi(x>\frac{\pi}{2})\left\{\frac{1}{2}+\ln(x)-\ln(\frac{\pi}{2})\right\}+\chi(x\leq\frac{\pi}{2})\frac{2x^{2}}{\pi^{2}}$ (6.17)
which is a differentiable lower bound.
Therefore we find
$2\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\langle h_{y},\eta_{s}\rangle\right|$
$\geq\frac{2\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\sum_{k\in\Gamma}\frac{\hat{V}(k)^{2}}{2k_{\textnormal{F}}|k|}\bigg{(}\chi(2k_{\textnormal{F}}|k|t>\frac{\pi}{2})\left\{\frac{1}{2}+\log(2k_{\textnormal{F}}|k|t)-\ln(\frac{\pi}{2})\right\}$
$+\chi(2k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{2(2k_{\textnormal{F}}|k|t)^{2}}{\pi^{2}}\bigg{)}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$\eqqcolon b_{t}+d$ (6.18)
with $d\in\max\{MN^{-\frac{2}{3}+\delta},\beta k_{\textnormal{F}}^{-1}\lambda^{-1}\}$ defined in (6.10) and
$\dot{b}_{t}$ $=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\sum_{k\in\Gamma}\hat{V}(k)^{2}\left(\chi(2k_{\textnormal{F}}|k|t>\frac{\pi}{2})\frac{1}{2k_{\textnormal{F}}|k|t}+\chi(2k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{8k_{\textnormal{F}}|k|t}{\pi^{2}}\right).$ (6.19)
In total, we can re-write the inequality (6.9) for $g_{t}\coloneqq\|\big{(}e^{i\widetilde{\mathbb{H}^{\text{eff}}}t}e^{iP(t)}W(\eta_{t})-1\big{)}\psi\|$ as the following integral inequality
$g_{t}\geq b_{t}-\theta\int_{0}^{t}\textnormal{d}sg_{s}$ (6.20)
with $b_{t}$ and $\theta$ defined in (6.18) and (6.12), respectively.
Claim.
We can bound $g_{t}$ from below for all $t\geq 0$ by a differentiable map $h_{t}$ which obeys the initial value problem
$h_{t}$ $=b_{t}-\theta\int_{0}^{t}\textnormal{d}sh_{s}\qquad\text{with }h_{0}=-d<0=g_{0}.$ (6.21)
Proof of claim.
Assume for the proof by contradiction that there exists a $t_{0}>0:h_{t_{0}}>g_{t_{0}}$ and set $I\subseteq\mathbb{R}_{>0}$ as the largest open interval satisfying $t_{0}\in I$ and $h_{t}>g_{t}$ for all $t\in I$. Note that since $h_{t}$ is differentiable such an open interval exists. It holds $t_{1}\coloneqq\inf I>0$ since $h_{0}<g_{0}$ and for all $t\in I$
$\dot{g}_{t}\geq\dot{b}_{t}-\theta g_{t}>\dot{b}_{t}-\theta h_{t}=\dot{h}_{t}$
and therefore
$\int_{t_{1}}^{t_{0}}\dot{g}_{s}\textnormal{d}s>\int_{t_{1}}^{t_{0}}\dot{h}_{s}\textnormal{d}s\implies h_{t_{1}}-g_{t_{1}}>h_{t_{0}}-g_{t_{0}}>0.$
Thus we obtain the desired contradiction to $t_{1}$ being the infimum of the largest set satisfying $h_{t}>g_{t}$. ∎
The solution of the initial value problem (6.21) is uniquely given by
$h_{t}=e^{-\theta t}\int_{0}^{t}\dot{b}_{s}e^{\theta s}\textnormal{d}s-d.$ (6.22)
Consequently it holds by inserting (6.19)
$h_{t}$ $=e^{-\theta t}\int_{0}^{t}\dot{b}_{s}e^{\theta s}\textnormal{d}s$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\sum_{k\in\Gamma}\hat{V}(k)^{2}e^{-\theta t}\int_{0}^{t}\bigg{(}\chi(2k_{\textnormal{F}}|k|s>\frac{\pi}{2})e^{\theta s}\frac{1}{2k_{\textnormal{F}}|k|s}$
$+\chi(2k_{\textnormal{F}}|k|s\leq\frac{\pi}{2})e^{\theta s}\frac{8k_{\textnormal{F}}|k|s}{\pi^{2}}\bigg{)}\textnormal{d}s-d$ (6.23)
$\geq\pi\lambda^{2}k_{\textnormal{F}}\sum_{k\in\Gamma}\hat{V}(k)^{2}\bigg{(}\chi(|k|t\leq\frac{\pi}{4k_{\textnormal{F}}})\frac{8k_{\textnormal{F}}|k|}{\theta^{2}\pi^{2}}\left(e^{-\theta t}+\theta t-1\right)+$
$+\chi(|k|t>\frac{\pi}{4k_{\textnormal{F}}})\frac{8k_{\textnormal{F}}|k|}{\theta^{2}\pi^{2}}\big{(}e^{-\theta\pi/(4k_{\textnormal{F}}|k|)}+\frac{\theta\pi}{4k_{\textnormal{F}}|k|}-1\big{)}\bigg{)}-d$ (6.24)
$\geq\frac{\lambda^{2}k_{\textnormal{F}}^{2}}{\theta^{2}\pi}\sum_{k\in\Gamma}\hat{V}(k)^{2}|k|\min\bigg{(}f(\theta t),f\big{(}\frac{\pi\theta}{4k_{\textnormal{F}}|k|}\big{)}\bigg{)}-d$ (6.25)
with $f(t)\coloneqq\left(e^{-t}+t-1\right)$ defining a non-negative monotonically increasing function. Recall that $d\in\max\{MN^{-\frac{2}{3}+\delta},\beta k_{\textnormal{F}}^{-1}\lambda^{-1}\}$ from (6.10) for all $0\leq t\lesssim k_{\textnormal{F}}^{-1}\lambda^{-1}$ and $\theta\leq C\lambda k_{\textnormal{F}}$ with a constant $C>0$ depending only on $V$ from (6.12). Thus, we obtain the desired result that (5.34) has a lower bound of order 1. ∎
Appendix A Estimates within the bosonization framework
In this section we collect all relevant estimates of the bosonization framework which was first developed in [BNP+19, BNP+21a, BNP+21b, BPSS23] in the semiclassical regime. In addition to the references we present brief proof sketches where we think they are helpful for the interested reader.
Lemma A.1 (Approximation of $n_{\alpha}(k)$, [BNP+21b, Lemma 5.1]).
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma,\alpha\in\mathcal{I}_{k}$ it holds
$n_{\alpha}(k)^{2}=\frac{4\pi k_{\textnormal{F}}^{2}}{M}|k\cdot\hat{\omega}_{\alpha}|\big{(}1+o(1)\big{)}.$
Note that $|k\cdot\hat{\omega}_{\alpha}|>N^{-\delta}$ by construction of $\mathcal{I}_{k}$.
Lemma A.2 (Approximation of $\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}$).
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma$ it holds
$\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}=k_{\textnormal{F}}^{2}|k|\pi\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$
Proof.
It holds from [BNP+19, Proposition 3.1]
$n_{\alpha}(k)^{2}=k_{\textnormal{F}}^{2}|k|\sigma(p_{\alpha})u_{\alpha}(k)^{2}\left(1+\mathcal{O}(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta})\right)$
with $\cos\theta_{\alpha}\coloneqq|\hat{k}\cdot\hat{\omega}_{\alpha}|\equiv u_{\alpha}(k)^{2}$. Choose $\varphi_{\alpha}$ for the azimuth angle of $\omega_{\alpha}$. We estimate the $\alpha$-sum by an appropriate surface integral over the patch $p_{\alpha}$
$\left|\int_{p_{\alpha}}\textnormal{d}\sigma\ \cos\theta-\sigma(p_{\alpha})\cos\theta_{\alpha}\right|$ $\leq\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\textnormal{d}}{\textnormal{d}\theta}\cos\theta\right|\sup_{(\theta,\varphi)\in p_{\alpha}}\left|\theta-\theta_{\alpha}\right|\sigma(p_{\alpha})$
$\leq CM^{-\frac{3}{2}}$
where we used $\left|\theta-\theta_{\alpha}\right|\leq CM^{-\frac{1}{2}}$ and $\sigma(p_{\alpha})\leq CM^{-1}$ by the patch construction. Note that the integral over $\tilde{S}\coloneqq\bigcup_{\alpha\in\mathcal{I}_{k}}p_{\alpha}$ which excludes a collar of width $N^{-\delta}$ can be approximated by an integral over the half-sphere $S$
$\left|\int_{\tilde{S}}\textnormal{d}\sigma\ \cos\theta-\int_{S}\textnormal{d}\sigma\ \cos\theta\right|<C\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)$
which can be calculated explicitly
$\int_{S}\textnormal{d}\sigma\ \cos\theta=2\pi\int_{0}^{\pi/2}\textnormal{d}\theta\ \cos\theta\sin\theta=\pi.$
Thus in total we obtain with the triangle inequality and $|k|<C$
$\left|\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}-k_{\textnormal{F}}^{2}|k|\pi\right|$ $=\left|\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)^{2}-k_{\textnormal{F}}^{2}|k|\int_{S}\textnormal{d}\sigma\ \cos\theta\right|$
$\leq Ck_{\textnormal{F}}^{2}|k|\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+M^{-\frac{1}{2}}+N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)$
$\leq Ck_{\textnormal{F}}^{2}|k|\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right).$
Lemma A.3 (Estimates on the CCR error term, [BNP+21b, Lemma 5.2]).
For $k^{\prime},k\in\Gamma$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$ the error term $\mathcal{E}_{\alpha}(k,k^{\prime})$ as defined in (2.22) satisfies $\mathcal{E}_{\alpha}(k,k^{\prime})=\mathcal{E}_{\alpha}(k^{\prime},k)^{\ast}$, commutes with $\mathcal{N}$ and for all $\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}$ it holds for all $\zeta\in\mathcal{F}$
$|\mathcal{E}_{\gamma}(k,k^{\prime})|^{2}\leq\sum_{\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}}|\mathcal{E}_{\gamma}(k,k^{\prime})|^{2}$ $\leq C\left(MN^{-\frac{2}{3}+\delta}\mathcal{N}\right)^{2},$
$\sum_{\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}}\|\mathcal{E}_{\gamma}(k,k^{\prime})\zeta\|$ $\leq CM^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}\|\mathcal{N}\zeta\|.$
Lemma A.4 (Pair operator bounds, [BNP+21b, Lemma 5.3]).
It holds for all $k\in\Gamma$ and $\psi\in\mathcal{F},f\in l^{2}(\mathcal{I}_{k})$:
  1. (i).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|^{2}\leq\|\mathcal{N}^{\frac{1}{2}}\psi\|^{2}$,
  2. (ii).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|\leq M^{\frac{1}{2}}\|\mathcal{N}^{\frac{1}{2}}\psi\|$,
  3. (iii).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|\leq M^{\frac{1}{2}}\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|$,
  4. (iv).
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|^{2}\leq\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|^{2}$,
  5. (v).
    $\|\sum_{\alpha\in\mathcal{I}_{k}}f_{\alpha}c_{\alpha}^{\ast}(k)\psi\|\leq\|f\|_{l^{2}}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|$,
  6. (vi).
    $\sum_{\alpha\in\mathcal{I}_{k}}c_{\alpha}^{\ast}(k)c_{\alpha}(k)\leq\mathcal{N}$.
Lemma A.5 (Error of linearized kinetic energy, [BNP+21b, Lemma 8.2]).
It holds for $k\in\Gamma$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{H}_{0},c_{\alpha}^{\ast}(k)]=2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
with
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|^{2}$ $\leq C\left(N^{\frac{1}{3}}M^{-\frac{1}{2}}\right)^{2}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|^{2}$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|$ $\leq CN^{\frac{1}{3}}\|\mathcal{N}^{\frac{1}{2}}\psi\|.$
Proof.
Observe that
$[\mathbb{H}_{0},c_{\alpha}^{\ast}(k)]$ $=\frac{1}{n_{\alpha}(k)}\sum_{p\in B_{\textnormal{F}}^{c}\cap B_{\alpha},p-k\in B_{\textnormal{F}}\cap B_{\alpha}}\sum_{l\in\mathbb{Z}^{3}}\Big{[}e(l)a_{l}^{\ast}a_{l},a_{p}^{\ast}a_{p-k}^{\ast}\Big{]}$
$=\frac{1}{n_{\alpha}(k)}\sum_{p\in B_{\textnormal{F}}^{c}\cap B_{\alpha},p-k\in B_{\textnormal{F}}\cap B_{\alpha}}\left(e(p)+e(p-k)\right)a_{p}^{\ast}a_{p-k}^{\ast}$
$=2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
One makes the identification $\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\equiv c_{\alpha}^{g}(k)$ representing a weighted operator (see [BNP+21b, eq. (5.11)]) with
$g(p,k)$ $=e(p)+e(p-k)-2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|=|p|^{2}-|p-k|^{2}-2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|$
$=\left(2k\cdot(p-k_{\textnormal{F}}\hat{\omega}_{\alpha})-|k|^{2}\right)$
where we used $e(p)$ as defined in (2.8). The bound follows from [BNP+21b, Lemma 5.4] which depends on $\|g\|_{l^{\infty}}$. This can be estimated by using $\text{diam}(B_{\alpha})\leq CN^{\frac{1}{3}}M^{-\frac{1}{2}}$ such that $|g(p,k)|\leq CN^{\frac{1}{3}}M^{-\frac{1}{2}}$. ∎
Lemma A.6 (Error of bosonized kinetic energy, [BNP+21b, eq. (8.6)]).
It holds for $k\in\Gamma$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{D}_{\textnormal{B}},c_{\alpha}^{\ast}(k)]=2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)^{\ast}$
with
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$ $\leq C\left(k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2},$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|$ $\leq Ck_{\textnormal{F}}M^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|.$
Proof.
It holds $\mathfrak{E}_{\alpha}^{\text{B}}(k)=2k_{\textnormal{F}}\sum_{l}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\chi(\alpha\in\mathcal{I}_{l})$ and therefore
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$
$\leq\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}||2k_{\textnormal{F}}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}\leq Ck_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}||\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq Ck_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}CMN^{-\frac{2}{3}+\delta}||\mathcal{N}c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq Ck_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\sum_{l\in\Gamma}\sum_{l^{\prime}\in\Gamma}\|c_{\alpha}(l)(\mathcal{N}-2)\psi\|\ \|c_{\alpha}(l^{\prime})(\mathcal{N}-2)\psi\|$
$\leq Ck_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2}$
where we used in the second line $l\in\Gamma$ bounded, Lemma A.3 in the third line, $\mathcal{N}c_{\alpha}(l)=c_{\alpha}(l)(\mathcal{N}-2)$ in the fourth line and Cauchy-Schwarz and Lemma A.4 in the last line.
The second statement simply follows from Cauchy-Schwarz. ∎
Lemma A.7 (Approximation of patch decomposed operators).
It holds for all $k\in\Gamma$
$\|\big{(}b(k)-\sum_{\alpha\in\mathcal{I}_{k}}n_{\alpha}(k)c_{\alpha}(k)+\text{h.c.}\big{)}\psi\|\leq C(N^{\frac{1}{3}-\frac{\delta}{2}}+N^{\frac{1}{6}}M^{\frac{1}{4}})\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|.$
Lemma A.8 (Estimate of non-bosonizable terms, [BNP+21a, eq. (4.6)]).
It holds for $\mathcal{E}$ as defined in (2.9b) the following estimate for all $\psi\in\mathcal{F}$
$\|\mathcal{E}\psi\|\leq C\lambda\|\hat{V}\|_{1}\|\mathcal{N}\psi\|.$
Acknowledgment
We are grateful for helpful discussions with David Mitrouskas, Benjamin Schlein and Karla Schön. This work has been partially supported by SFB/Transregio TRR 352 – Project-ID 470903074.
Declaration
Conflict of interest
The authors declare that there is no conflict of interest.